### 19 Mar 2026 v1

# Agent Control Protocol ACP v1.13: Admission Control for Agent Actions Technical Specification and Reference Implementation

Abstract
Agent Control Protocol (ACP) is a formal technical specification for governance of autonomous agents in B2B institutional environments. ACP is the admission control layer between agent intent and system state mutation: before any agent action reaches execution, it must pass a cryptographic admission check that validates identity, capability scope, delegation chain, and policy compliance simultaneously.
ACP defines the mechanisms of cryptographic identity, capability-based authorization, deterministic risk evaluation, verifiable chained delegation, transitive revocation, and immutable auditing that a system must implement for autonomous agents to operate under explicit institutional control.
ACP operates as an additional layer on top of RBAC and Zero Trust, without replacing them. It is designed specifically for the problem that neither model solves: governing what an autonomous agent can do, under what conditions, with what limits, and with complete traceability for external auditing—including across organizational boundaries.
The v1.13 specification comprises 36 technical documents organized into five conformance levels (L1–L5). It includes a Go reference implementation of 22 packages covering all L1–L4 capabilities, 51 signed conformance test vectors (Ed25519 + SHA-256), and an OpenAPI 3.1.0 specification for all HTTP endpoints. It defines more than 62 verifiable requirements, 12 prohibited behaviors, and the mechanisms for interoperability between institutions.
Specification and implementation: https://github.com/chelof100/acp-framework-en
1  The Problem ACP Solves
Autonomous agents are being deployed in institutional environments without a technical standard to govern their behavior. This is not a tooling problem—it is a protocol problem.
1.1  The Structural Gap
When an autonomous agent makes a decision and executes it, there is a critical moment between both actions. In current models, that moment does not formally exist: the decision and the execution are the same event. The agent decides and acts. There is no intermediate validation. There is no point of intervention. There is no structured record of why the decision was made.
This is acceptable when a human executes that action, because the human bears responsibility and can be questioned. An autonomous agent cannot be questioned. It can only be audited—and only if there is something to audit.
The problem is not whether agents are trustworthy. The problem is that currently no formal technical mechanism exists that allows an institution to demonstrate that its agents operated within authorized limits.
1.2  ACP as Admission Control
The clearest frame for understanding what ACP does is the Kubernetes Admission Controller analogy.
Kubernetes intercepts every API request before it reaches the cluster and runs it through a sequence of admission checks—ValidatingWebhookConfiguration, ResourceQuota enforcement, OPA Gatekeeper policies. If any check fails, the request is rejected before touching cluster state.
ACP applies this pattern to agent actions:
agent intent
[1] Identity check (ACP-AGENT-1.0, ACP-HP-1.0)
| is this agent who they claim to be?
[2] Capability check (ACP-CT-1.0, ACP-DCMA-1.1)
| does the agent hold a valid token for this action?
[3] Policy check (ACP-RISK-1.0, ACP-PSN-1.0)
| is this action within current policy?
[4] ADMIT / DENY / ESCALATE
| (if ADMIT)
[5] Execution token (ACP-EXEC-1.0)
| single-use cryptographic proof of admission
[6] Ledger record (ACP-LEDGER-1.3)
| immutable signed audit entry
system state mutation
The critical difference from Kubernetes: ACP’s admission check operates across institutional boundaries. An agent from Bank A can be admitted by Bank B without Bank B trusting Bank A’s internal infrastructure—the cryptographic proof is self-contained and verifiable with Bank A’s published public key alone.
This admission control framing also clarifies the relationship with related tools:
  • OPA (Open Policy Agent) [3] can serve as the policy evaluation engine inside Step 3. ACP does not replace OPA; it adds the identity and delegation chain layers above it.
  • AWS IAM / Azure RBAC model static role permissions for humans. ACP adds dynamic agent delegation with execution proof.
  • OAuth 2.0 [1] handles API access tokens. ACP extends delegation to multi-agent chains with non-escalation and verifiable provenance.
  • SPIFFE / SPIRE [5] provides cryptographic workload identity. ACP builds on that identity to add capability scoping and governance.
1.3  Why RBAC and Zero Trust Are Insufficient
RBAC and Zero Trust are the predominant control layers in enterprise environments. Both are necessary. Neither solves the problem of governing autonomous agents:
Criterion RBAC Zero Trust ACP
Designed for Human roles Network access Autonomous agents
Cryptographic identity No Partial Yes (Ed25519, mandatory)
Verifiable dynamic delegation No No Yes (chained, auditable)
Decision/execution separation No No Yes (Execution Tokens)
Real-time risk evaluation No Partial Yes (deterministic)
Multi-institutional auditing Non-standard Non-standard Native (signed ledger)
Transitive delegation revocation No No Yes (formal propagation)
B2B interoperability for agents Unstructured Unstructured Central protocol design
Table 1: Comparison of ACP with RBAC and Zero Trust across agent governance criteria.
ACP does not replace RBAC or Zero Trust. It adds a governance layer oriented specifically to autonomous agents that operates above existing controls.
1.4  The Concrete Scenario ACP Prevents
Without ACP: A financial processing agent receives instructions from another agent to execute a transfer. The agent executes the action. If the instruction was legitimate, everything works. If it was compromised, injected, or generated by an unauthorized agent, the transfer occurs anyway. No formal mechanism exists to prevent it, nor is there a technical record that allows reconstructing the authorization chain.
With ACP: The agent requesting the transfer must present a Capability Token cryptographically signed by the issuing institution, demonstrate possession of the associated private key, and the request passes through the risk engine before receiving an Execution Token. The ET is single-use. The entire chain is recorded in the Audit Ledger with an externally verifiable institutional signature.
2  What ACP Is
A formal technical specification—not a framework, not a platform, not a set of best practices. A protocol with precise definitions, formal state models, verifiable flows, and explicit conformance requirements.
2.1  Definition
Agent Control Protocol (ACP) is a technical specification that defines the mechanisms by which autonomous institutional agents are identified, authorized, monitored, and governed in B2B environments. It establishes the formal contract between an agent, the institution that operates it, and the institutions with which it interacts.
Core invariant:
$\textit{Execute}(\text{request})\Longrightarrow\textit{ValidIdentity}(\text{agent})\;\wedge\;\textit{ValidCapability}\;\wedge\;\textit{ValidDelegationChain}\;\wedge\;\textit{AcceptableRisk}$
No action of an ACP agent can be executed without all four predicates being simultaneously true. If any fails, the action is denied. No exceptions.
2.2  Design Principles
ACP was designed with five principles that are non-negotiable at implementation time:
  • P1
    Fail Closed. On any internal component failure, the action is denied. Never approved by default.
  • P2
    Identity is cryptography. $\text{AgentID}=\text{base58}(\text{SHA-256}(\text{public\_key}))$. No usernames. No arbitrary IDs. Identity cannot be claimed—it must be demonstrated in every request.
  • P3
    Delegation does not expand privileges. The delegated agent’s permissions are always a strict subset of the delegator’s permissions. This property is cryptographically verified at every chain hop.
  • P4
    Complete auditability. Every decision—approved, denied, or escalated—is recorded in an append-only ledger with institutional signature. Not just successes. Everything.
  • P5
    External verification possible. Any institution can verify ACP artifacts from another institution using only the public key registered in the ITA. No dependency on proprietary systems.
2.3  Formal Agent Model
In ACP, an agent is a formal tuple with well-defined state:
$A=(\,\text{AgentID},\;\text{capabilities},\;\text{autonomy\_level},\;\text{state},\;\text{limits}\,)$
Field Type Description
AgentID String (43–44 chars) base58(SHA-256(pk)). Derived from public key. Immutable.
capabilities List of strings Explicit permissions. Format: acp:cap:<domain>.<action>. Never abstract roles.
autonomy_level Integer 0–4 Determines risk evaluation thresholds. 0 = no autonomy. 4 = maximum.
state Enum active | restricted | suspended | revoked. Transition to revoked is unidirectional.
limits Object Rate limits, maximum amounts, time windows. Not modifiable at runtime.
Table 2: ACP formal agent tuple fields.
2.4  Layered Architecture
ACP does not replace existing security infrastructure. It is added as an upper layer:
ACP Layer -- Autonomous agent governance: identity, authorization, risk, auditing
RBAC Layer -- Role-based access control for human users
Zero Trust -- Continuous identity and network access verification
3  Technical Mechanisms
ACP defines six interdependent mechanisms. Each has its own formal specification, state model, data structure, protocol flow, and error codes.
3.1  Serialization and Signing (ACP-SIGN-1.0)
Every verification in ACP begins with signature verification. ACP-SIGN-1.0 defines the exact process that produces a binary result—valid or invalid—without ambiguity:
  1. 1.
    Canonicalization with JCS (RFC 8785) [4]. Produces a deterministic representation of the JSON object, independent of field order and the system that generated it.
  2. 2.
    SHA-256 hash over the canonical output in UTF-8.
  3. 3.
    Ed25519 signature (RFC 8032) [2] over the hash. 32-byte key, 64-byte signature.
  4. 4.
    Base64url encoding without padding for transmission.
Signature verification precedes all semantic validation. An object with an invalid signature is rejected without processing its content. This rule has no exceptions (PROHIB-003, PROHIB-012).
3.2  Capability Token (ACP-CT-1.0)
The Capability Token is ACP’s central artifact. It is a signed JSON object that specifies exactly what an agent can do, on what resource, for how long, and whether it can delegate that capability to other agents.
"ver": "1.0",
"iss": "<AgentID_issuer>",
"sub": "<AgentID_subject>",
"cap": ["acp:cap:financial.payment"],
"res": "org.example/accounts/ACC-001",
"exp": 1718923600,
"nonce": "<128bit_CSPRNG_base64url>",
"deleg": { "allowed": true, "max_depth": 2 },
"parent_hash": null,
"sig": "<Ed25519_base64url>"
Critical fields: exp is mandatory—a token without expiry is invalid by definition. The 128-bit nonce prevents replay attacks. parent_hash chains delegated tokens in a verifiable way. The signature covers all fields except sig.
3.3  Handshake and Proof-of-Possession (ACP-HP-1.0)
Possessing a valid Capability Token is not sufficient to act. ACP-HP-1.0 requires that the bearer demonstrate in every request that they possess the private key corresponding to the AgentID declared in the token. This eliminates the possibility of impersonating an agent with a stolen token.
The protocol is stateless—it does not establish sessions, does not produce a session_id, does not require server-side state between requests. The proof occurs in every interaction:
  1. 1.
    The receiving system issues a 128-bit challenge generated by CSPRNG, valid for 30 seconds and single-use.
  2. 2.
    The agent signs the challenge together with the HTTP method, path, and body hash of the request.
  3. 3.
    The receiver verifies the signature using the agent’s public key, obtained from the ITA.
  4. 4.
    The challenge is deleted immediately after use—it cannot be reused.
This sequence guarantees four formal properties: identity authentication, cryptographic request binding, anti-replay, and transport channel independence.
3.4  Deterministic Risk Evaluation (ACP-RISK-1.0)
Each authorization request passes through a deterministic risk function that produces a Risk Score (RS) in the range $[0,100]$. The same input always produces the same result—no stochastic elements, no machine learning in the critical path.
$\text{RS}=\min\!\bigl(100,\;B(c)+F_{\text{ctx}}(x)+F_{\text{hist}}(h)+F_{\text{res}}(r)\bigr)$
Factor Description Example values
$B(c)$ Baseline by capability *.read = 0, financial.payment = 35
$F_{\text{ctx}}(x)$ Request context Non-corporate IP +20+20; outside hours +15+15
$F_{\text{hist}}(h)$ Agent history (24h) Recent denial +20+20; anomalous frequency +15+15
$F_{\text{res}}(r)$ Resource classification public = 0; sensitive = 15; restricted = 45
Table 3: ACP-RISK-1.0 risk scoring factors.
The RS determines the decision according to the thresholds configured for the agent’s autonomy_level. With autonomy_level 2 (standard): $\text{RS}\leq 39$ $\rightarrow$ APPROVED; $\text{RS}\in[40,69]$ $\rightarrow$ ESCALATED; $\text{RS}\geq 70$ $\rightarrow$ DENIED. An agent with autonomy_level 0 always receives DENIED.
3.5  Verifiable Chained Delegation (ACP-DCMA-1.1)
ACP allows an agent to delegate capabilities to another agent, which in turn can delegate to a third, up to the maximum depth defined in the root token. Delegation guarantees three properties:
  • No privilege escalation. The delegated agent’s capability set is always $\subseteq$ the delegator’s set. Cryptographically verified at every hop via the parent_hash field.
  • Bounded depth. The max_depth field establishes the chain limit. A chain that exceeds it is invalid.
  • Transitive revocation. Revoking an agent’s token automatically invalidates all delegated tokens that descend from it.
3.6  Execution Token (ACP-EXEC-1.0)
The separation between authorization and execution is a core principle of ACP. When the authorization engine approves a request, it returns an Execution Token (ET): a single-use artifact with a short lifetime that authorizes exactly that action, on that resource, at that moment.
  • An ET can only be consumed once. A second presentation is rejected (PROHIB-002).
  • An expired ET is invalid even if it was never used.
  • The target system that consumes the ET notifies the ACP endpoint, closing the audit cycle.
3.7  Audit Ledger (ACP-LEDGER-1.3)
The Audit Ledger is a chain of cryptographically signed events where each event includes the hash of the previous event:
$h_{n}=\text{SHA-256}(\,e_{n}\;\|\;h_{n-1}\,)$
This structure makes it impossible to modify or delete an event without invalidating the entire subsequent chain. The ledger records all lifecycle event types: GENESIS, AUTHORIZATION (including DENIED and ESCALATED), RISK_EVALUATION, TOKEN_ISSUED, TOKEN_REVOKED, EXECUTION_TOKEN_ISSUED, and EXECUTION_TOKEN_CONSUMED.
Institutions with FULL conformance expose the ledger via GET /acp/v1/audit/query, allowing external partners to verify chain integrity using only the institutional ITA public key. Modifying or deleting ledger events is prohibited (PROHIB-007, PROHIB-008).
It is important to note that the ledger provides verifiable evidence of execution, not enforcement itself. Enforcement in ACP occurs at the policy evaluation and execution layers (ACP-EXEC-1.0, ACP-DCMA-1.1): the Execution Token is the cryptographic artifact that gates actual system-state mutation. The ledger’s role is to make every admission decision—and its outcome—auditable and tamper-evident after the fact.
4  Inter-Institutional Trust
In a B2B environment, agents of one institution interact with another institution’s systems. ACP defines the exact mechanism by which this trust is established, verified, and can be revoked.
4.1  Institutional Trust Anchor (ACP-ITA-1.0)
The ITA is the authoritative registry that links an institution_id to an Ed25519 public key. It is the only point where ACP depends on an out-of-band mechanism: the initial distribution of the ITA authority’s public key. Once that key is resolved, all subsequent verification is autonomous and cryptographic.
Each institution registers in the ITA its Root Institutional Key (RIK)—the private key held in HSM that never leaves it. All ACP artifacts from that institution (tokens, ledger events, API responses) are signed with that key. Any third party can verify them by resolving the public key from the ITA.
4.2  Mutual Recognition (ACP-ITA-1.1)
When two institutions operate under different ITA authorities, ACP-ITA-1.1 defines the mutual recognition protocol. The process requires both authorities to sign a Mutual Recognition Agreement (MRA) establishing: the scope of recognition, the agreement’s validity period, and the proxy resolution mechanism.
Recognition is explicitly non-transitive. If A recognizes B and B recognizes C, A does not automatically recognize C. Each bilateral relationship requires its own signed MRA. This prevents uncontrolled expansion of the trust graph.
4.3  Institutional Key Rotation and Revocation
Normal rotation includes a transition period of up to 7 days during which both keys are valid, allowing artifacts signed with the old key to be verified during the transition.
Emergency rotation is activated when a key is compromised. The result is immediate: the key is marked revoked, all artifacts signed with it are invalid from that moment, and there is no transition period.
5  Cross-Organization Execution
ACP-CROSS-ORG-1.1 defines a fault-tolerant bilateral protocol for interactions between independent institutions, each operating its own ACP ledger. The protocol closes five gaps identified in the 1.0 version: undefined status tracking, absent retry protocol, unregistered ACK event type, pending_review without SLA, and a stale header reference.
5.1  Interaction Model
A cross-organization interaction involves two ACP-compliant institutions: a source institution that originates the bundle and a target institution that validates, executes, and acknowledges it. The protocol operates asynchronously—the source does not block waiting for an ACK.
Each interaction carries a mandatory interaction_id (UUIDv7), immutable and reused across all retries. This is distinct from event_id (UUIDv4), which is unique per emission. The target deduplicates by interaction_id, not by event_id.
For example, an agent in Institution A may request execution in Institution B. The request is recorded as a CROSS_ORG_INTERACTION event in A’s ledger, transmitted to B, validated against B’s policy and capability registry, and resolved through a CROSS_ORG_ACK event registered in both ledgers. Institution A derives final execution state from the ACK without storing mutable status.
5.2  Fault Tolerance and Retry Protocol
When no CROSS_ORG_ACK is received within the 300-second timeout, the source retries up to three times with exponential backoff:
Attempt Wait after timeout
1 (initial)
2 +30 s
3 +60 s
4 (final) +120 s
After three failed attempts the interaction transitions to retry_exhausted (CROSS-012) and an operational alert is raised. If a valid ACK arrives at any point, all retry timers are cancelled immediately.
5.3  Derived Interaction Status
The status of a cross-organization interaction is never stored as mutable state. It is always derived from events present in the ledger:
Derived status Condition
pending_ack CROSS_ORG_INTERACTION exists, no ACK
acked ACK with status = accepted
rejected ACK with status = rejected
pending_review ACK with status = pending_review
expired pending_review AND now > review_deadline
Precedence rule: accepted >> rejected >> pending_review. A mutable CROSS_ORG_STATUS_UPDATE event is explicitly prohibited as an anti-pattern that introduces race conditions and divergent audit trails.
5.4  Interaction State Invariants
For any interaction_id, ACP enforces the following invariants:
  • At most one terminal state. Only one accepted or rejected ACK may exist per interaction_id; a duplicate with a different payload is rejected (CROSS-014).
  • ACK dominates retries. A valid CROSS_ORG_ACK cancels all pending retry timers immediately, regardless of attempt number.
  • Immutable correlation key. Retry operations MUST NOT create new interaction_id values; only event_id changes across attempts.
  • No terminal-to-non-terminal regression. Once accepted or rejected, the status cannot change. The transition pending_review $\to$ pending_review is also prohibited (CROSS-015).
5.5  Pending Review SLA
When the target returns pending_review, a 24-hour SLA applies. The ACK payload includes a review_deadline field (Unix seconds). Valid terminal transitions are: accepted, rejected, or expired (implicit, when now > review_deadline).
5.6  CROSS_ORG_ACK as a First-Class Ledger Event
CROSS_ORG_ACK is registered in ACP-LEDGER-1.3 §5.15 as a first-class event type, signed with Ed25519 and serialized via JCS (RFC 8785). This enables verifiable and auditable execution across independent institutional boundaries, eliminating the audit gap that existed in version 1.0.
5.7  Security Considerations
ACP-CROSS-ORG-1.1 assumes authenticated communication channels between institutions. All CROSS_ORG_INTERACTION and CROSS_ORG_ACK events are signed with the originating institution’s Ed25519 ITA key and serialized via JCS (RFC 8785), providing authenticity, integrity, and non-repudiation.
Without this cryptographic layer, cross-organization events would be subject to forgery, replay, and unauthorized status derivation. Specifically, ACP mitigates:
  • Replay attacks: interaction_id (UUIDv7) and event_id (UUIDv4) allow the target to detect and reject duplicate submissions.
  • Event forgery: Ed25519 signatures over JCS-canonical payloads prevent an adversary from injecting valid-looking ACK or interaction events.
  • Audit divergence: Because CROSS_ORG_ACK is a first-class ledger event on both sides, independent auditors can verify consistency without trusting either institution’s mutable state.
ACP does not eliminate all sources of failure in distributed systems, but provides a structured model for detecting, handling, and auditing them.
6  Security Model
ACP explicitly defines which threats it mitigates, which properties it guarantees, and which risks fall outside its scope.
6.1  Threat Model (STRIDE)
Category Threat Mitigation in ACP
Spoofing AgentID impersonation $\text{AgentID}=\text{SHA-256}(\text{pk})$. Without valid signature $\rightarrow$ immediate DENIED.
Tampering Token or event alteration Ed25519 covers all fields. Chained ledger: altering one event invalidates all subsequent.
Repudiation Agent denies executed action ActionRequest digitally signed. Non-repudiation by design.
Info Disclosure Capability exposure Tokens reveal only the necessary subset. Channel confidentiality via TLS.
DoS Request flooding Rate limits per agent_id. Anomalous frequency control.
Elevation Delegation that expands privileges $C_{\text{delegated}}\subseteq C_{\text{original}}$. Cryptographically verified at each hop.
Table 4: STRIDE threat model and ACP mitigations.
6.2  Guaranteed Security Properties
ACP guarantees the following properties when the implementation is compliant:
  • Artifact integrity. EUF-CMA security of Ed25519. Impossible to modify a token or event without invalidating the signature.
  • Identity authenticity. Only whoever possesses sk can generate a valid signature under the corresponding pk.
  • No privilege escalation via delegation. Demonstrable by induction over the delegation chain.
  • Anti-replay. The single-use challenge makes reusing a proof of possession ineffective.
  • Effective revocation. $\textit{Valid}(t)=\textit{valid\_sig}\wedge\neg\textit{expired}\wedge\neg\textit{revoked}\wedge\textit{valid\_delegation}$. All four conditions must be true simultaneously.
6.3  Declared Residual Risks
ACP explicitly declares what it cannot resolve:
  • Total compromise of the RIK held in HSM. ACP defines the emergency rotation process but cannot prevent a physical compromise of the custody infrastructure.
  • Coordinated institutional collusion. If multiple institutions act maliciously in a coordinated manner, they can generate valid artifacts. ACP guarantees traceability, not prevention of malicious agreements between parties.
  • Implementation failures. ACP is a specification. An implementation that violates prohibited behaviors can compromise all protocol guarantees. Conformance requires formal testing.
  • ITA bootstrap. The only point that depends on an out-of-band channel. Once the ITA root key is resolved, everything subsequent is autonomous.
7  Conformance and Interoperability
7.1  Conformance Levels
Level Required documents Enabled capability
L1 — CORE ACP-SIGN-1.0, ACP-CT-1.0, ACP-HP-1.0 Token issuance with cryptographic PoP
L2 — SECURITY L1 + ACP-RISK-1.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval, transitive revocation, ITA
L3 — FULL L2 + ACP-API-1.0, ACP-EXEC-1.0, ACP-LEDGER-1.3 Complete verifiable auditing
L4 — EXTENDED L3 + ACP-PAY-1.0, ACP-NOTIFY-1.0, ACP-DISC-1.0, … Full governance suite
L5 — DECENTRAL. L4 + ACP-D suite Federation without central ITA
Table 5: ACP-CONF-1.2 conformance levels.
7.2  Conformance Declaration
Every compliant implementation MUST expose a public endpoint without authentication:
GET https://<endpoint>/acp/v1/conformance
This endpoint returns the achieved level, implemented documents, declared extensions, and declaration date. It allows any external partner to verify the conformance level of a counterparty before establishing an ACP relationship.
7.3  Prohibited Behaviors
ACP defines 12 behaviors that no compliant implementation can exhibit. If any is exhibited, the implementation cannot declare conformance at any level:
Code Prohibited behavior
PROHIB-001 Approving a request when any evaluation component fails
PROHIB-002 Reusing an already-consumed Execution Token
PROHIB-003 Omitting signature verification on any incoming artifact
PROHIB-004 Treating a not-found token_id as active in a revocation context
PROHIB-005 Allowing state transition from revoked
PROHIB-006 Issuing an ET without a prior APPROVED AuthorizationDecision
PROHIB-007 Modifying or deleting Audit Ledger events
PROHIB-008 Silencing ledger corruption detection
PROHIB-009 Ignoring max_depth in delegation chains
PROHIB-010 Implementing an offline policy more permissive than ACP-REV-1.0
PROHIB-011 Approving requests from agents with autonomy_level 0
PROHIB-012 Continuing to process an artifact with an invalid signature
Table 6: ACP prohibited behaviors. Exhibiting any disqualifies conformance at all levels.
7.4  B2B Interoperability Conditions
  • L1 Interoperability: Institution A can verify tokens from B if both implement ACP-CONF-L1, A has access to B’s public key (via ITA or out-of-band), and B’s tokens use ACP-SIGN-1.0 algorithms.
  • L2 Interoperability: A can delegate to B’s agents if both implement ACP-CONF-L2, are registered in a common ITA or with mutual recognition, and B’s revocation endpoint is accessible to A.
  • L3 Interoperability: A can audit B’s ledger if B implements ACP-CONF-L3, A can resolve B’s public key via ITA, and B exposes GET /acp/v1/audit/query.
8  Use Cases
ACP is sector-agnostic. The mechanisms are identical regardless of industry. What varies is the configuration of capabilities, resources, and autonomy levels.
8.1  Financial Sector — Inter-Institutional Payment Agents
ACP-PAY-1.0 extends the capability registry with formal specifications for acp:cap:financial.payment and acp:cap:financial.transfer. Mandatory constraints in the token include max_amount and currency. 12 specific validation steps cover limit verification, beneficiary validation, and time window control. In a financial B2B scenario, Bank A can authorize an agent to execute payments up to a defined amount to pre-approved beneficiaries, with a complete record verifiable by Bank B without needing shared proprietary systems.
8.2  Digital Government — Document Processing
Government agents that process documents can operate under ACP with autonomy_level 1 or 2, requiring human review for any action with a Risk Score above the configured threshold. Ledger traceability provides forensic evidence for regulatory audits and transparency processes.
8.3  Enterprise AI — Multi-Company Orchestration
In agent pipelines that cross organizational boundaries, ACP allows each organization to maintain formal control over what other organizations’ agents can do in their systems. Chained delegation allows an agent in company A to operate in company B’s systems with explicitly delegated capabilities, without B needing to trust A’s internal controls—only the chain of signed tokens.
8.4  Critical Infrastructure — Monitoring and Actuation Agents
For systems where an incorrect action has irreversible consequences, autonomy_level 0 ensures any actuation request is DENIED without evaluating the Risk Score. The ledger provides the forensic record needed for post-incident analysis.
9  Specification and Implementation Status
ACP v1.13 is a complete Draft Standard specification with a full Go reference implementation.
9.1  Active Specifications — v1.13 (36 documents)
Level Document Title
L1 ACP-SIGN-1.0 Serialization and Signature
L1 ACP-AGENT-1.0 Agent Identity
L1 ACP-CT-1.0 Capability Tokens
L1 ACP-CAP-REG-1.0 Capability Registry
L1 ACP-HP-1.0 Handshake / Proof-of-Possession
L1 ACP-DCMA-1.1 Delegated Chain Multi-Agent
L1 ACP-MESSAGES-1.0 Wire Message Format
L1 ACP-PROVENANCE-1.0 Authority Provenance
L2 ACP-RISK-1.0 Deterministic Risk Engine
L2 ACP-REV-1.0 Revocation Protocol
L2 ACP-ITA-1.0 Institutional Trust Anchor
L2 ACP-ITA-1.1 ITA Mutual Recognition
L2 ACP-REP-1.2 Reputation Module
L2 ACP-REP-PORTABILITY-1.0 Reputation Portability
L3 ACP-API-1.0 HTTP API
L3 ACP-EXEC-1.0 Execution Tokens
L3 ACP-LEDGER-1.3 Audit Ledger (mandatory institutional sig)
L3 ACP-PSN-1.0 Policy Snapshot
L3 ACP-POLICY-CTX-1.0 Policy Context Snapshot
L3 ACP-LIA-1.0 Liability Attribution
L4 ACP-HIST-1.0 History Query API
L4 ACP-PAY-1.0 Financial Capability
L4 ACP-NOTIFY-1.0 Event Notifications
L4 ACP-DISC-1.0 Service Discovery
L4 ACP-BULK-1.0 Batch Operations
L4 ACP-CROSS-ORG-1.1 Cross-Organization Bundles
L4 ACP-GOV-EVENTS-1.0 Governance Event Stream
Gov. ACP-CONF-1.2 Conformance — sole normative source
Gov. ACP-TS-1.1 Test Vector Format
Gov. RFC-PROCESS Specification Process
Gov. RFC-REGISTRY Specification Registry
Gov. ACR-1.0 Change Request Process
Table 7: ACP v1.13 active specification documents by conformance level (32 primary documents shown). The remaining 4 documents are the ACP-D decentralized suite (L5, design phase) not yet published. Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0) archived in archive/specs/.
9.2  Reference Implementation — 22 Go Packages
The Go reference implementation in impl/go/ covers all L1–L4 conformance levels. All 22 packages pass go test ./.... A Python SDK (impl/python/) covers the ACP-HP-1.0 handshake and all ACP-API-1.0 endpoints.
Package Spec Level
pkg/crypto ACP-SIGN-1.0 + ACP-AGENT-1.0 L1
pkg/tokens ACP-CT-1.0 L1
pkg/registry ACP-CAP-REG-1.0 L1
pkg/handshake ACP-HP-1.0 L1
pkg/delegation ACP-DCMA-1.1 L1
pkg/provenance ACP-PROVENANCE-1.0 L1
pkg/risk ACP-RISK-1.0 L2
pkg/revocation ACP-REV-1.0 L2
pkg/reputation ACP-REP-1.2 L2
pkg/execution ACP-EXEC-1.0 L3
pkg/ledger ACP-LEDGER-1.3 L3
pkg/psn ACP-PSN-1.0 L3
pkg/policyctx ACP-POLICY-CTX-1.0 L3
pkg/lia ACP-LIA-1.0 L3
pkg/hist ACP-HIST-1.0 L4
pkg/govevents ACP-GOV-EVENTS-1.0 L4
pkg/notify ACP-NOTIFY-1.0 L4
pkg/disc ACP-DISC-1.0 L4
pkg/bulk ACP-BULK-1.0 L4
pkg/crossorg ACP-CROSS-ORG-1.1 L4
pkg/pay ACP-PAY-1.0 L4
cmd/acp-server ACP-API-1.0 L3
Table 8: Go reference implementation packages. All pass go test ./....
9.3  Conformance Test Vectors — 51 Signed Vectors
The compliance/test-vectors/ directory contains 51 signed test vectors per ACP-TS-1.1:
Suite Positive Negative Spec
CORE (SIGN, CT, HP) 4 4 L1
DCMA 2 2 L1
HP 2 8 L1
LEDGER 3 8 L3
EXEC 2 7 L3
PROV 2 7 L1
Total 15 36
Table 9: Conformance test vector suites. All positive vectors carry real Ed25519 signatures (RFC 8037 Test Key A) and real SHA-256 hash chains.
9.4  Roadmap
Item Status
Core specs (L1–L4), 36 documents Complete
Go reference implementation (22 packages) Complete
Conformance test vectors (51 signed) Complete
OpenAPI 3.1.0 (openapi/acp-api-1.0.yaml) Complete
Python SDK (impl/python/) Complete (L1 + full API client)
Docker image (ghcr.io/chelof100/acp-server:latest) Complete
L5 Decentralized (ACP-D) Specification in design (v2.0)
Post-quantum algorithm migration Research phase
IETF RFC submission After L5 stabilization
Table 10: ACP v1.13 development roadmap.
10  How to Implement ACP
ACP is a specification—it does not require adopting any specific platform. It can be implemented on top of existing infrastructure.
10.1  Minimum Requirements for L1 Conformance
  • Ed25519 public key infrastructure. Key pair for each agent.
  • JCS implementation (RFC 8785). Deterministic canonicalization for all signed artifacts.
  • Capability Token issuance and verification with all mandatory fields per ACP-CT-1.0 §5.
  • Handshake endpoint to issue and verify challenges (ACP-HP-1.0 §6).
  • Capability registry with the core domains of ACP-CAP-REG-1.0.
10.2  Additional Requirements for L3 Conformance
  • Root Institutional Key held in HSM with documented rotation process.
  • Registration with an ITA authority (centralized or federated model).
  • Deterministic risk engine with the four factors of ACP-RISK-1.0.
  • Revocation endpoint (Mechanism A: online endpoint, or Mechanism B: CRL).
  • Append-only storage for the Audit Ledger with per-event signing.
  • Complete HTTP API per ACP-API-1.0, including health and conformance endpoints.
  • Public conformance declaration at GET /acp/v1/conformance.
10.3  What ACP Does Not Prescribe
ACP defines the what—mechanisms, flows, data structures, requirements. It does not prescribe the how of internal implementation: programming language, database, HSM provider, ITA provider, or specific integration with existing RBAC or Zero Trust systems.
11  Conclusion
Autonomous agents are already operating in institutional environments. The question is not whether they will operate—it is whether they will do so with or without formal governance. ACP proposes that they operate with formal governance, with verifiable mechanisms, and with traceability that can withstand an external audit.
ACP is not the first attempt to control autonomous agents. It is the first attempt to do so through a formal technical specification with precise state models, demonstrable security properties, and verifiable conformance requirements. The difference between a best-practices policy and a formal protocol is exactly that: behaviors are defined, failures have specific error codes, and conformance can be verified.
The goal of ACP is not to make agents more capable. It is to make them governable. That is a necessary condition for their institutional deployment to be sustainable at scale.
ACP does not eliminate all sources of failure in distributed agent systems. It does not prescribe network transport, consensus mechanisms, or business-level dispute resolution. What it provides is a structured, auditable, and formally specified model for admission control, delegation, revocation, and cross-organization interaction— the layer that existing identity and policy frameworks leave unaddressed.
The v1.13 specification is complete. A full Go reference implementation (22 packages, L1–L4) and 51 signed conformance test vectors are publicly available at https://github.com/chelof100/acp-framework-en. The specification and implementation are open for technical review, pilot implementation, and formal standardization.
TraslaIA invites organizations interested in adopting ACP, contributing to its evolution, or participating in the standardization process to reach out directly.
Marcelo Fernandez | TraslaIA
info@traslaia.com | https://agentcontrolprotocol.xyz
Appendix A Glossary
Term Definition
AgentID Cryptographic identifier: base58(SHA-256(Ed25519_public_key)). Immutable and unforgeable.
Capability Token (CT) Signed JSON artifact authorizing an agent to perform specific actions on a defined resource during a limited period.
Execution Token (ET) Single-use artifact issued after an APPROVED decision. Authorizes exactly that action at that moment.
ITA Institutional Trust Anchor. Authoritative registry linking institution_id to institutional Ed25519 public key.
RIK Root Institutional Key. Institution’s Ed25519 key pair held in HSM.
Risk Score (RS) Integer in $[0,100]$ produced by the deterministic risk function. Determines the authorization decision.
Autonomy Level Integer 0–4 assigned to an agent determining applicable risk evaluation thresholds.
PoP Proof-of-Possession. Cryptographic proof that the bearer of a CT possesses the corresponding private key.
Audit Ledger Chain of signed events where $h_{n}=\text{SHA-256}(e_{n}\|h_{n-1})$. Append-only and immutable.
MRA Mutual Recognition Agreement. Bilateral document signed by two ITA authorities for cross-authority interoperability.
ESCALATED ACP decision when RS is in the intermediate range. Action not executed until explicit resolution.
Fail Closed Design principle: on any internal failure, the action is denied. Never approved by default.
Table 11: ACP Glossary.
Acknowledgements
The ACP specification emerged from analysis of operational challenges in deploying autonomous agents in regulated B2B environments. The protocol design was informed by the IETF RFC process, the Kubernetes admission control architecture, and the SPIFFE/SPIRE workload identity model.

### 24 Mar 2026 v4

@@
-Agent Control Protocol ACP v1.13: Admission Control for Agent Actions Technical Specification and Reference Implementation
+Agent Control Protocol ACP v1.17: Admission Control for Agent Actions Technical Specification and Reference Implementation
 Abstract
@@
 It is designed specifically for the problem that neither model solves: governing what an autonomous agent can do, under what conditions, with what limits, and with complete traceability for external auditing—including across organizational boundaries.
-The v1.13 specification comprises 36 technical documents organized into five conformance levels (L1–L5).
-It includes a Go reference implementation of 22 packages covering all L1–L4 capabilities, 51 signed conformance test vectors (Ed25519 + SHA-256), and an OpenAPI 3.1.0 specification for all HTTP endpoints.
-It defines more than 62 verifiable requirements, 12 prohibited behaviors, and the mechanisms for interoperability between institutions.
+The v1.17 specification comprises 38 technical documents organized into five conformance levels (L1–L5).
+It includes a Go reference implementation of 23 packages covering all L1–L4 capabilities, 73 signed conformance test vectors (Ed25519 + SHA-256) plus 65 unsigned RISK-2.0 vectors, an OpenAPI 3.1.0 specification (18 endpoints), an executable payment-agent demonstration (examples/payment-agent/), and an executable multi-organization interoperability demo.
+v1.17 adds external verifiability: a TLC-runnable TLA+ formal model with three checked invariants, an ACR-1.0 sequence compliance runner (library mode and HTTP mode), five stateful sequence test vectors covering cooldown, F_anom Rule 3, boundary conditions, privilege jump, and benign flow, and an ACP-SIGN-2.0 HYBRID mode stub for the post-quantum transition.
+It defines 12 explicitly prohibited behaviors whose presence disqualifies conformance at any level, and the mechanisms for interoperability between institutions.
 Specification and implementation: https://github.com/chelof100/acp-framework-en
@@
 | does the agent hold a valid token for this action?
-[3] Policy check (ACP-RISK-1.0, ACP-PSN-1.0)
+[3] Policy check (ACP-RISK-2.0, ACP-PSN-1.0)
 | is this action within current policy?
@@
 This sequence guarantees four formal properties: identity authentication, cryptographic request binding, anti-replay, and transport channel independence.
-3.4  Deterministic Risk Evaluation (ACP-RISK-1.0)
+3.4  Deterministic Risk Evaluation (ACP-RISK-2.0)
 Each authorization request passes through a deterministic risk function that produces a Risk Score (RS) in the range $[0,100]$.
 The same input always produces the same result—no stochastic elements, no machine learning in the critical path.
-$\text{RS}=\min\!\bigl(100,\;B(c)+F_{\text{ctx}}(x)+F_{\text{hist}}(h)+F_{\text{res}}(r)\bigr)$
+$\text{RS}=\min\!\bigl(100,\;B(c)+F_{\text{res}}(r)+F_{\text{ctx}}(x)+F_{\text{hist}}(h)+F_{\text{anom}}(q)\bigr)$
+where $F_{\text{anom}}(q)$ is the anomaly factor introduced in v2.0, evaluated against the agent’s accumulated request history $q$ (pattern frequency, denial rate, cooldown state).
+When an agent accumulates three DENIED decisions within ten minutes, a cooldown is activated and subsequent requests are blocked regardless of RS.
 Factor Description Example values
@@
 $F_{\text{res}}(r)$ Resource classification public = 0; sensitive = 15; restricted = 45
-Table 3: ACP-RISK-1.0 risk scoring factors.
+Table 3: ACP-RISK-2.0 risk scoring factors (base formula; $F_{\text{anom}}$ adds up to +30+30 from pattern accumulation).
 The RS determines the decision according to the thresholds configured for the agent’s autonomy_level.
@@
 ACP does not eliminate all sources of failure in distributed systems, but provides a structured model for detecting, handling, and auditing them.
+6  Reputation Snapshot Portability (ACP-REP-PORTABILITY-1.1)
+ACP-REP-PORTABILITY-1.1 defines the ReputationSnapshot: a compact, cryptographically signed record that carries an agent’s reputation score across organizational boundaries.
+Unlike the bilateral attestation protocol of v1.0, this specification focuses on the snapshot object itself—its structure, signing procedure, validation algorithm, and expiration semantics—enabling any verifier to independently validate a snapshot without trusting an intermediary.
+6.1  ReputationSnapshot Structure
+A ReputationSnapshot contains ten fields: ver (version), rep_id (UUID v4), subject_id, issuer, score (float), scale ("0-1" or "0-100"), model_id, evaluated_at (Unix seconds), valid_until (Unix seconds), and signature (Ed25519, base64url).
+Fields scale, model_id, and valid_until are new in v1.1; v1.0 snapshots omit them and are accepted without expiration enforcement (backward compatibility, §12).
+6.2  Signing Procedure
+The signature covers all fields except signature itself, serialized via JCS (RFC 8785) canonicalization, SHA-256 digest, and Ed25519 signing:
+  1. 1.
+    raw = json.Marshal(signableReputation)
+  2. 2.
+    canonical = jcs.Transform(raw) // RFC 8785
+  3. 3.
+    digest = SHA-256(canonical)
+  4. 4.
+    sig = Ed25519.Sign(privKey, digest)
+  5. 5.
+    snapshot.signature = base64url(sig) // no padding
+JCS canonicalization is mandatory.
+Using json.Marshal directly is prohibited: key ordering differs across Go, Python, and TypeScript implementations, producing verification failures in cross-org deployments.
+6.3  Validation Invariants
+The validator enforces five invariants (REP-001 through REP-011): (1) evaluated_at $\leq$ valid_until prevents retroactively backdated snapshots; (2) now $\leq$ valid_until enforces expiration (v1.1 only, error REP-011); (3) score within scale bounds (REP-002); (4) non-empty issuer (REP-004); (5) valid Ed25519 signature (REP-010).
+Structural validation (Validate) and cryptographic verification (VerifySig) are intentionally separate operations, enabling lightweight ingestion-time checks without requiring the issuer’s public key.
+6.4  Divergence Semantics
+When a verifier receives snapshots of the same subject_id from multiple issuers, it may compute divergence: $|a.\text{score}-b.\text{score}|$.
+If the divergence exceeds a configurable threshold (default 0.30 for scale="0-1"), the verifier emits warning REP-WARN-002.
+This is non-blocking—the policy decision of whether to accept, escalate, or reject remains with the verifier’s business logic, preserving institutional sovereignty.
+7  Multi-Organization Interoperability Demo (GAP-14)
+The examples/multi-org-demo/ directory provides an executable end-to-end demonstration of cross-organizational ACP exchange, combining ACP-POLICY-CTX-1.1 and ACP-REP-PORTABILITY-1.1 in a single runnable scenario deployable with Docker in under five minutes.
+7.1  Scenario
+Two independent organizations, Org-A and Org-B, interact over HTTP.
+Org-A is the issuer: it evaluates an agent request against its local policy and produces two signed artifacts—a PolicyContextSnapshot and a ReputationSnapshot—using its Ed25519 institutional key.
+Org-B is the verifier: it independently validates both artifacts using the pkg/policyctx and pkg/reputation packages (no logic duplicated) and renders an autonomous admission decision.
+The validation sequence in Org-B is:
+  1. 1.
+    policyctx.VerifySig(pcs, pubKey) — verifies institutional signature (ACP-POLICY-CTX-1.1 §6 step 12).
+  2. 2.
+    policyctx.VerifyCaptureFreshness(pcs, 300s) — enforces $\delta_{\max}$ freshness invariant.
+  3. 3.
+    reputation.Validate(rep, now) — structural invariants REP-001 through REP-011.
+  4. 4.
+    reputation.VerifySig(rep, pubKey) — Ed25519/JCS/SHA-256 cryptographic verification.
+  5. 5.
+    reputation.CheckDivergence(orgA, orgB, 0.30) — divergence check against Org-B’s own score.
+  6. 6.
+    Final ACCEPT / DENY decision under Org-B’s sovereign policy.
+7.2  Divergence Handling and Institutional Sovereignty
+If Org-B holds its own score for the same subject_id, it computes divergence $|a.\text{score}-b.\text{score}|$.
+If this exceeds 0.30, Org-B emits REP-WARN-002.
+REP-WARN-002 is non-blocking—the final decision is Org-B’s prerogative, independent of Org-A’s assessment.
+This embodies the ACP invariant: ACP reports divergence.
+ACP does not resolve divergence.
+Org-B cannot extend Org-A’s valid_until, override Org-A’s delta_max, or substitute its own snapshot for Org-A’s signed record.
+Each institution’s attestations are cryptographically bound to its own key and immutable after issuance.
+7.3  Deployment
+The demo ships with a single Dockerfile (parameterized by ARG ORG=org-a|org-b), a docker-compose.yml with health check, and a dedicated Go module with a replace directive pointing to impl/go/:
+docker compose up --build # from examples/multi-org-demo/
+curl http://localhost:8081/request | jq .
 8  Security Model
@@
 L1 — CORE ACP-SIGN-1.0, ACP-CT-1.0, ACP-HP-1.0 Token issuance with cryptographic PoP
-L2 — SECURITY L1 + ACP-RISK-1.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval, transitive revocation, ITA
+L2 — SECURITY L1 + ACP-RISK-2.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval (F_anom + Cooldown), transitive revocation, ITA
 L3 — FULL L2 + ACP-API-1.0, ACP-EXEC-1.0, ACP-LEDGER-1.3 Complete verifiable auditing
@@
 11  Specification and Implementation Status
-ACP v1.13 is a complete Draft Standard specification with a full Go reference implementation.
-9.1  Active Specifications — v1.13 (36 documents)
+ACP v1.17 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
+11.1  Active Specifications — v1.17 (38 documents)
 Level Document Title
@@
 L1 ACP-PROVENANCE-1.0 Authority Provenance
-L2 ACP-RISK-1.0 Deterministic Risk Engine
+L1 ACP-SIGN-2.0 Hybrid Signature (Ed25519 + ML-DSA-65)
+L2 ACP-RISK-2.0 Deterministic Risk Engine v2 (F_anom + Cooldown)
+L2 ACP-RISK-1.0 Deterministic Risk Engine v1 (superseded by v2)
 L2 ACP-REV-1.0 Revocation Protocol
@@
 L2 ACP-REP-1.2 Reputation Module
-L2 ACP-REP-PORTABILITY-1.0 Reputation Portability
 L3 ACP-API-1.0 HTTP API
@@
 L3 ACP-PSN-1.0 Policy Snapshot
-L3 ACP-POLICY-CTX-1.0 Policy Context Snapshot
+L3 ACP-POLICY-CTX-1.1 Policy Context Snapshot
 L3 ACP-LIA-1.0 Liability Attribution
@@
 L4 ACP-GOV-EVENTS-1.0 Governance Event Stream
+L4 ACP-REP-PORTABILITY-1.1 Reputation Snapshot Portability
 Gov. ACP-CONF-1.2 Conformance — sole normative source
@@
 Gov. RFC-REGISTRY Specification Registry
-Gov. ACR-1.0 Change Request Process
-Table 7: ACP v1.13 active specification documents by conformance level (32 primary documents shown).
-The remaining 4 documents are the ACP-D decentralized suite (L5, design phase) not yet published.
-Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0) archived in archive/specs/.
-9.2  Reference Implementation — 22 Go Packages
+Gov. ACP-CR-1.0 Change Request Process
+Table 7: ACP v1.17 active specification documents by conformance level (38 primary documents shown).
+The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published.
+New in v1.16: ACP-RISK-2.0 (L2, deterministic engine v2 with F_anom + Cooldown), ACP-SIGN-2.0 (L1, PQC hybrid spec).
+New in v1.17: ACR-1.0 sequence compliance runner, TLA+ formal model, ACP-SIGN-2.0 HYBRID stub (pkg/sign2/).
+Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0, REP-PORTABILITY-1.0) archived in archive/specs/.
+11.2  Reference Implementation — 23 Go Packages + Demos
 The Go reference implementation in impl/go/ covers all L1–L4 conformance levels.
-All 22 packages pass go test ./....
+All 23 packages pass go test ./....
 A Python SDK (impl/python/) covers the ACP-HP-1.0 handshake and all ACP-API-1.0 endpoints.
+The examples/multi-org-demo/ directory (GAP-14) provides a runnable two-organization scenario using the real pkg/policyctx and pkg/reputation packages; see Section 7.
+The examples/payment-agent/ directory (Sprint E) provides an executable HTTP server demonstrating the full ACP-RISK-2.0 pipeline: POST /admission evaluates the request with the deterministic risk engine, appends an immutable ledger event with the full factor breakdown, and triggers cooldown automatically after three DENIED decisions in ten minutes.
 Package Spec Level
@@
 pkg/provenance ACP-PROVENANCE-1.0 L1
-pkg/risk ACP-RISK-1.0 L2
+pkg/risk ACP-RISK-1.0 + ACP-RISK-2.0 L2
 pkg/revocation ACP-REV-1.0 L2
-pkg/reputation ACP-REP-1.2 L2
+pkg/reputation ACP-REP-1.2 + ACP-REP-PORTABILITY-1.1 L2/L4
 pkg/execution ACP-EXEC-1.0 L3
@@
 pkg/psn ACP-PSN-1.0 L3
-pkg/policyctx ACP-POLICY-CTX-1.0 L3
+pkg/policyctx ACP-POLICY-CTX-1.1 L3
 pkg/lia ACP-LIA-1.0 L3
@@
 cmd/acp-server ACP-API-1.0 L3
-Table 8: Go reference implementation packages. All pass go test ./....
-9.3  Conformance Test Vectors — 51 Signed Vectors
-The compliance/test-vectors/ directory contains 51 signed test vectors per ACP-TS-1.1:
+Table 8: Go reference implementation packages (23 total). All pass go test ./....
+11.3  Conformance Test Vectors — 73 Signed Vectors
+The compliance/test-vectors/ directory contains 73 signed test vectors per ACP-TS-1.1:
 Suite Positive Negative Spec
@@
 PROV 2 7 L1
-Total 15 36
+PCTX 4 9 L3
+REP 3 6 L4
+Total 22 51
 Table 9: Conformance test vector suites. All positive vectors carry real Ed25519 signatures (RFC 8037 Test Key A) and real SHA-256 hash chains.
+11.4  ACP-RISK-2.0 Engine Performance
+The pkg/risk ACP-RISK-2.0 engine was benchmarked using go test -bench on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, Windows, GOMAXPROCS=8).
+These measurements reflect the pure in-memory evaluation cost, excluding network and HTTP parsing overhead.
+Benchmark ns/op B/op allocs/op
+Evaluate (APPROVED, no anomaly) 1,012 280 5
+Evaluate (DENIED, no anomaly) 863 296 5
+Evaluate (all 3 F_anom rules) 1,331 296 5
+Evaluate (COOLDOWN short-circuit) 149 112 1
+PatternKey (SHA-256 hash) 996 184 4
+ShouldEnterCooldown 88 0 0
+[…]
+The deterministic design of ACP-RISK-2.0 — integer arithmetic, fixed rules, no floating-point, no external ML inference — yields predictable latency bounds.
+At these throughput characteristics, the admission check adds negligible overhead relative to the actions it governs.
+12  Evaluation
+This section evaluates ACP along three dimensions: (i) computational overhead of the decision function, (ii) behavioral correctness under stateful multi-step scenarios, and (iii) reproducibility and external verifiability of decision outcomes.
+12.1  Evaluation Goals
+  • Q1: What is the computational overhead of ACP decision evaluation relative to the actions it governs?
+  • Q2: Does the system produce correct decisions under stateful conditions such as repeated denials, pattern accumulation, and cooldown activation?
+  • Q3: Can ACP executions be deterministically reproduced and externally verified without access to internal implementation details?
+12.2  Experimental Setup
+All experiments use the Go reference implementation (v1.17) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8).
+The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 13).
+Production deployments with persistent storage backends will incur additional I/O latency not measured here.
+12.3  Q1: Computational Overhead
+Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 1,331 ns. The cooldown short-circuit path (149 ns) bypasses the risk function entirely.
+Answer to Q1: The admission check adds between 149 ns and 1,331 ns of latency per request.
+Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action.
+12.4  Q2: Stateful Behavior Correctness
+The ACR-1.0 compliance runner executes each sequence vector end-to-end under the execution contract of Section 13. Results:
+Vector Behavior tested Result
+SEQ-BENIGN-001 No false positives under repeated reads PASS
+SEQ-BOUNDARY-001 Exact thresholds RS = 0, 25, 35, 40, 70 PASS
+SEQ-PRIVJUMP-001 Low→high privilege jump caught immediately PASS
+SEQ-FANOM-RULE3-001 Pattern count $\geq 3$ adds +15+15 at step 4 PASS
+SEQ-COOLDOWN-001 3 DENIED in 10 min → cooldown active PASS
+Total 5/5 PASS
+Table 11: ACR-1.0 sequence vector results under strict mode (--strict).
+Answer to Q2: All five stateful scenarios pass with CONFORMANT status.
+State transitions follow the defined execution contract; cooldown activates deterministically after the threshold is crossed; pattern accumulation influences decisions exactly as specified in ACP-RISK-2.0 Rule 3.
+12.5  Q3: Reproducibility and External Verifiability
+Given the same sequence of inputs and the same initial state, the ACR-1.0 runner produces identical outputs on every run. This holds because:
+  • Evaluate() is a pure function of its arguments and the querier state,
+  • state transitions follow the deterministic execution contract,
+  • and the test vectors ship with the specification, enabling any third party to reproduce results without access to proprietary infrastructure.
+Answer to Q3: ACP executions are fully reproducible and externally verifiable.
+This is a property not typically supported by conventional policy engines, which evaluate policy against a snapshot of state without defining how that state evolves across requests [12, 6, 9].
+12.6  Limitations
+The evaluation uses an in-memory state backend.
+Persistent storage, distributed deployments, and concurrent request patterns introduce latency and consistency considerations not measured here.
+Adversarial robustness under coordinated attack scenarios is not evaluated.
+These are identified as future work (Section 12 roadmap).
 12.7  Roadmap
 Item Status
-Core specs (L1–L4), 36 documents Complete
-Go reference implementation (22 packages) Complete
-Conformance test vectors (51 signed) Complete
-OpenAPI 3.1.0 (openapi/acp-api-1.0.yaml) Complete
+Core specs (L1–L4), 38 documents Complete
+Go reference implementation (23 packages) Complete
+Conformance test vectors (73 signed + 65 unsigned RISK-2.0) Complete
+OpenAPI 3.1.0 (openapi/acp-api-1.0.yaml, 18 endpoints) Complete
 Python SDK (impl/python/) Complete (L1 + full API client)
 Docker image (ghcr.io/chelof100/acp-server:latest) Complete
+Multi-org interoperability demo (examples/multi-org-demo/) Complete (GAP-14)
+ACP-RISK-2.0 (F_anom + Cooldown, pkg/risk) v1.16 Complete
+Payment-agent demo (examples/payment-agent/) v1.16 Complete
+ACP-SIGN-2.0 spec (Ed25519 + ML-DSA-65 hybrid) v1.16 Complete
+ACR-1.0 sequence compliance runner + 5 stateful vectors v1.17 Complete
+TLA+ formal model (TLC-runnable, 3 invariants) v1.17 Complete
+ACP-SIGN-2.0 HYBRID stub (pkg/sign2/) v1.17 Complete
+Post-quantum Go implementation (Dilithium, circl) v1.18
 L5 Decentralized (ACP-D) Specification in design (v2.0)
-Post-quantum algorithm migration Research phase
 IETF RFC submission After L5 stabilization
-Table 10: ACP v1.13 development roadmap.
+Table 12: ACP v1.17 development roadmap.
+13  Compliance Testing (ACR-1.0 Sequence Runner)
+We introduce a compliance runner, ACR-1.0 (ACP Compliance Runner 1.0), that validates ACP-RISK-2.0 implementations against sequence-based test cases.
+The runner operates in two modes: library mode (direct call to pkg/risk, no network, deterministic) and HTTP mode (external server validation for interoperability).
+All 73 signed single-shot conformance vectors plus 5 sequence scenarios are validated on every commit.
+13.1  Sequence-Based Test Vectors
+Unlike single-shot conformance vectors, sequence-based vectors capture temporal dynamics and emergent risk patterns across multiple requests to the same engine instance.
+The compliance runner executes steps in order, maintains state between them (request history, pattern counts, denial counts, cooldown), and validates each step’s decision and risk_score against expected values.
+The five sequence vectors in compliance/test-vectors/sequence/ cover:
+Vector ID Steps Behavior Tested Key Invariant
+SEQ-BENIGN-001 3 Benign repeated reads Baseline: RS=0, no state buildup
+SEQ-BOUNDARY-001 3 RS boundary conditions Exact thresholds: 0, 25, 35, 40, 70
+SEQ-PRIVJUMP-001 2 Low-risk → high-risk jump No residual benefit from prior APPROVED
+SEQ-FANOM-RULE3-001 4 F_anom Rule 3 activation Pattern count $\geq 3$ at step 4 adds +15+15
+SEQ-COOLDOWN-001 4 Cooldown trigger and block 3 DENIED in 10 min → cooldown active
+Table 13: ACR-1.0 sequence test vectors. All 5/5 pass in library mode with the Go reference implementation.
+Execution contract.
+The runner implements the execution contract defined by ACP-RISK-2.0: evaluate first (stateless), then update state.
+Critically, AddPattern() is called after Evaluate(), so the pattern count visible to step $n$’s evaluation is n1n-1.
+F_anom Rule 3 (pattern count $\geq 3$) therefore first triggers on step 4, not step 3.
+This contract is what makes the system testable and verifiable by third parties.
+now := time.Now() // capture once
+result, _ := risk.Evaluate(req, querier) // 1: stateless eval
+querier.AddRequest(req.AgentID, now) // 2: always
+querier.AddPattern(patKey, now) // 3: always (feeds F_anom Rule 3)
+if result.Decision == risk.DENIED {
+querier.AddDenial(req.AgentID, now) // 4: conditional
+if should, _ := risk.ShouldEnterCooldown(...) // 5: check threshold
+querier.SetCooldown(agentID, now.Add(p)) // 6: set expiry
+Listing 1: ACR-1.0 execution contract in library mode.
+13.2  Model–Implementation Alignment
+The TLA+ model, test vectors, and runtime evaluation are aligned through a shared request abstraction:
+TLA+ variable Runner field Engine field
+capability, resource RunnerRequest EvalRequest
+risk_score risk_score RSFinal
+decision decision Decision
+ledger (audit log) InMemoryQuerier events
+cooldown denied_reason CooldownActive()
+Table 14: Alignment across TLA+ formal model, ACR-1.0 runner, and ACP-RISK-2.0 engine.
+14  End-to-End Verifiability
+ACP v1.17 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction.
+Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
+TLA+ ModelSafety $\wedge$ AppendOnly $\wedge$ DeterminismTest Vectorssequence/ + single-shotComplianceRunner (ACR-1.0)ACP Enginepkg/riskAudit Ledgerappend-onlydefinesinputcallsrecords statically checked by TLC (0 violations, bounded state space)
+Figure 1: ACP end-to-end verifiability pipeline.
+The TLA+ model defines the invariants that test vectors instantiate and the compliance runner validates at runtime.
+The ACP engine records every decision in the append-only audit ledger.
+The TLA+ layer is a static verification artifact; it does not run at admission time.
+What is formally verified in TLA+ holds at runtime.
+Each TLA+ invariant is instantiated as a concrete test vector and validated empirically by the compliance runner.
+This is the connection across the three layers of the v1.17 verifiability stack: formal $\to$ data $\to$ runtime.
 15  How to Implement ACP
@@
   • Registration with an ITA authority (centralized or federated model).
-  • Deterministic risk engine with the four factors of ACP-RISK-1.0.
+  • Deterministic risk engine implementing ACP-RISK-2.0 (F_res, F_ctx, F_hist, F_anom, Cooldown).
   • Revocation endpoint (Mechanism A: online endpoint, or Mechanism B: CRL).
@@
 It does not prescribe the how of internal implementation: programming language, database, HSM provider, ITA provider, or specific integration with existing RBAC or Zero Trust systems.
+16  Related Work
+16.1  Access Control and Policy Enforcement
+Traditional access control models such as Role-Based Access Control (RBAC) [12] and Attribute-Based Access Control (ABAC) [6] provide structured mechanisms for regulating access decisions based on predefined roles or attributes.
+Policy languages such as XACML [9] and modern policy engines such as Open Policy Agent [10] and Cedar [1] extend these paradigms with expressive evaluation engines.
+These systems are fundamentally designed around stateless decision evaluation: they evaluate a request against policy at a single point in time.
+They do not define how system state evolves as a result of decisions over time, and they do not formalize how historical behavior (repeated denials, temporal patterns) influences subsequent decisions in a deterministic and auditable manner.
+ACP differs by explicitly separating: (i) stateless decision evaluation, (ii) externalized state management, and (iii) an execution contract governing state evolution across requests.
+This enables reasoning not only about individual decisions, but about the temporal behavior of the admission system.
+16.2  Runtime Policy Systems and Agent Frameworks
+Recent protocols for autonomous agents and multi-agent coordination define structured interactions between components and services.
+The Model Context Protocol (MCP) [2] provides structured tool access between LLM applications and services.
+Agent-to-Agent (A2A) [4] defines communication and task delegation between agents.
+While these protocols address interoperability and orchestration, they do not define:
+  • a deterministic risk evaluation model that accumulates state across requests,
+  • explicit state evolution semantics with a testable execution contract, or
+  • externally verifiable execution traces that can be reproduced by third parties.
+ACP complements these approaches by providing a governance layer that is independent of orchestration protocols, focusing specifically on verifiable decision enforcement and auditability.
+16.3  Auditability and Tamper-Evident Logs
+Secure audit logging [13] and event sourcing [3] ensure that past events can be reconstructed and verified after the fact.
+These approaches operate post hoc: they provide evidence of what occurred without constraining system behavior in real time.
+ACP integrates auditability directly into the admission process.
+Every decision—APPROVED, ESCALATED, or DENIED—is recorded in an append-only, cryptographically chained ledger (ACP-LEDGER-1.3) before the governed action executes.
+This enables both real-time enforcement and post-hoc verification from the same artifact.
+The LedgerAppendOnlyTemporal invariant in the TLA+ model formally specifies that ledger entries are never modified or removed.
+16.4  Formal Verification of Protocols
+Model checking [8] and symbolic reasoning are used to establish correctness properties in distributed systems and protocols, typically focusing on safety and liveness of state transitions.
+ACP incorporates formal methods in a targeted manner: the TLA+ model (tla/ACP.tla) mechanically checks three critical invariants (fail-closed admission, append-only ledger, risk determinism) using TLC.
+Rather than attempting full-system verification, ACP adopts a pragmatic approach in which formally verified invariants are instantiated as concrete sequence test vectors and validated at runtime by the ACR-1.0 compliance runner.
+This creates a three-layer verifiability chain: formal $\to$ data $\to$ runtime.
+16.5  Summary
+Existing systems address individual aspects of the problem—policy evaluation, agent orchestration, audit logging, or formal verification.
+ACP’s contribution lies in combining these elements into a unified model centered on an explicit execution contract.
+This enables a property not directly provided by prior systems: the ability to deterministically reproduce and externally verify the behavior of an admission decision system over time.
 17  Conclusion
@@
 What it provides is a structured, auditable, and formally specified model for admission control, delegation, revocation, and cross-organization interaction— the layer that existing identity and policy frameworks leave unaddressed.
-The v1.13 specification is complete.
-A full Go reference implementation (22 packages, L1–L4) and 51 signed conformance test vectors are publicly available at https://github.com/chelof100/acp-framework-en.
+The v1.17 specification is complete.
+A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (3 invariants, 0 violations), an ACP-SIGN-2.0 HYBRID mode stub with Ed25519 and ML-DSA-65 migration path, an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent killer demo (examples/payment-agent/), and an extended OpenAPI 3.1.0 specification (18 endpoints) are publicly available at https://github.com/chelof100/acp-framework-en.
+Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19198298
 The specification and implementation are open for technical review, pilot implementation, and formal standardization.
@@
 Term Definition
+ACR-1.0 ACP Compliance Runner 1.0. Sequence-based test runner for ACP-RISK-2.0. Distinct from ACP-CR-1.0 (Change Request Process governance document).
 AgentID Cryptographic identifier: base58(SHA-256(Ed25519_public_key)). Immutable and unforgeable.
@@
 Table 15: ACP Glossary.
+Appendix B Formal Verification (ACP-RISK-2.0 — TLC-Runnable)
+The tla/ACP.tla module is a complete TLC-runnable TLA+ model of the ACP-RISK-2.0 admission control engine.
+It checks three invariants over a bounded state space (2 agents $\times$ 4 capabilities $\times$ 3 resources, ledger depth $\leq 5$):
+---- MODULE ACP ----
+EXTENDS Sequences, Integers, TLC
+CONSTANTS Agents, Capabilities, Resources
+CapabilityBase(cap) ==
+    CASE cap = "admin"     -> 60
+      [] cap = "financial" -> 35
+      [] cap = "write"     -> 10
+      [] cap = "read"      -> 0
+      [] OTHER             -> 20
+ResourceScore(res) ==
+    CASE res = "restricted" -> 45
+      [] res = "sensitive"  -> 15
+      [] res = "public"     -> 0
+      [] OTHER              -> 0
+ComputeRisk(cap, res) ==
+    LET raw == CapabilityBase(cap) + ResourceScore(res)
+    IN IF raw > 100 THEN 100 ELSE raw
+Decide(rs) ==
+    IF      rs >= 70 THEN "DENIED"
+    ELSE IF rs >= 40 THEN "ESCALATED"
+    ELSE                  "APPROVED"
+VARIABLE ledger
+INIT == ledger = << >>
+EvaluateRequest(a, cap, res) ==
+    /\ Len(ledger) < 5
+    /\ ledger’ = Append(ledger,
+            [ agent      |-> a,
+              capability |-> cap,
+              resource   |-> res,
+              risk_score |-> ComputeRisk(cap, res),
+              decision   |-> Decide(ComputeRisk(cap, res)) ])
+[…]
+            EvaluateRequest(a, cap, res)
+Spec == INIT /\ [][NEXT]_ledger
+Safety ==
+[…]
+        ledger[i].decision = "APPROVED" => ledger[i].risk_score <= 39
+LedgerAppendOnly ==
+[…]
+        /\ ledger[i].decision $\in$ {"APPROVED", "ESCALATED", "DENIED"}
+        /\ ledger[i].risk_score >= 0
+RiskDeterminism ==
+[…]
+[…]
+            ( ledger[i].capability = ledger[j].capability
+           /\ ledger[i].resource   = ledger[j].resource )
+            ledger[i].risk_score = ledger[j].risk_score
+LedgerAppendOnlyTemporal ==
+    [][ /\ Len(ledger’) >= Len(ledger)
+[…]
+TLC result.
+Running java -jar tla2tools.jar -config ACP.cfg ACP.tla with $|\mathit{Agents}|=2$, $|\mathit{Capabilities}|=4$, $|\mathit{Resources}|=3$, ledger bound =5=5 produces: Model checking completed.
+No error has been found.
+All four declared invariants/properties (TypeInvariant, Safety, LedgerAppendOnly, RiskDeterminism) and the temporal property LedgerAppendOnlyTemporal hold over the full reachable state space.
+Interpretation.
+Safety encodes fail-closed admission: every APPROVED decision had $\mathit{RS}\leq 39$.
+LedgerAppendOnly encodes tamper-evidence in a state-based view: every ledger entry carries a valid decision and a non-negative risk score.
+LedgerAppendOnlyTemporal is the temporal statement: in every step, existing entries are preserved and the ledger never shrinks.
+RiskDeterminism is the central invariant of ACP-RISK-2.0: identical (capability, resource) pairs always produce identical risk scores, enabling third-party recalculation from a signed policy snapshot.
 Acknowledgements

### 25 Mar 2026 v5

@@
-Agent Control Protocol ACP v1.17: Admission Control for Agent Actions Technical Specification and Reference Implementation
+Agent Control Protocol ACP v1.19: Admission Control for Agent Actions Technical Specification and Reference Implementation
 Abstract
-Agent Control Protocol (ACP) is a formal technical specification for governance of autonomous agents in B2B institutional environments.
-ACP is the admission control layer between agent intent and system state mutation: before any agent action reaches execution, it must pass a cryptographic admission check that validates identity, capability scope, delegation chain, and policy compliance simultaneously.
-ACP defines the mechanisms of cryptographic identity, capability-based authorization, deterministic risk evaluation, verifiable chained delegation, transitive revocation, and immutable auditing that a system must implement for autonomous agents to operate under explicit institutional control.
-ACP operates as an additional layer on top of RBAC and Zero Trust, without replacing them.
-It is designed specifically for the problem that neither model solves: governing what an autonomous agent can do, under what conditions, with what limits, and with complete traceability for external auditing—including across organizational boundaries.
-The v1.17 specification comprises 38 technical documents organized into five conformance levels (L1–L5).
-It includes a Go reference implementation of 23 packages covering all L1–L4 capabilities, 73 signed conformance test vectors (Ed25519 + SHA-256) plus 65 unsigned RISK-2.0 vectors, an OpenAPI 3.1.0 specification (18 endpoints), an executable payment-agent demonstration (examples/payment-agent/), and an executable multi-organization interoperability demo.
-v1.17 adds external verifiability: a TLC-runnable TLA+ formal model with three checked invariants, an ACR-1.0 sequence compliance runner (library mode and HTTP mode), five stateful sequence test vectors covering cooldown, F_anom Rule 3, boundary conditions, privilege jump, and benign flow, and an ACP-SIGN-2.0 HYBRID mode stub for the post-quantum transition.
-It defines 12 explicitly prohibited behaviors whose presence disqualifies conformance at any level, and the mechanisms for interoperability between institutions.
+[…]
 Specification and implementation: https://github.com/chelof100/acp-framework-en
@@
 The entire chain is recorded in the Audit Ledger with an externally verifiable institutional signature.
+1.5  Design Insight
+ACP is compute-cheap but state-sensitive: decision evaluation is computationally inexpensive, while system performance is primarily constrained by state access patterns.
+This separation enables a protocol that is both efficient in isolation ($\sim$820 ns per decision, $\sim$88 ns for the cooldown short-circuit path) and adaptable in deployment: performance can be improved by replacing the state backend without modifying protocol semantics.
+The LedgerQuerier interface formalizes this boundary—the decision function remains stateless and deterministic; the state backend can be a local in-memory structure for development or a distributed, indexed store in production.
+1.6  Contributions
+This paper makes the following contributions:
+  • Admission control protocol.
+A deterministic admission control protocol for autonomous agents combining capability, resource, historical, and anomaly-based signals into a single verifiable decision.
+  • State-aware execution model.
+An execution model that explicitly separates decision logic from state management via the LedgerQuerier abstraction, enabling deterministic reasoning about temporal agent behavior.
+[…]
+  • End-to-end verifiability.
+A three-layer verifiability pipeline: formal (TLA+, TLC-checked invariants), data (signed conformance test vectors), and runtime (ACR-1.0 compliance runner), enabling independent validation without proprietary infrastructure.
+[…]
 2  What ACP Is
@@
 The same input always produces the same result—no stochastic elements, no machine learning in the critical path.
+ACP deliberately trades expressiveness for deterministic auditability and reproducibility: every decision can be reconstructed from first principles given only the input request and the current state snapshot.
 $\text{RS}=\min\!\bigl(100,\;B(c)+F_{\text{res}}(r)+F_{\text{ctx}}(x)+F_{\text{hist}}(h)+F_{\text{anom}}(q)\bigr)$
@@
 curl http://localhost:8081/request | jq .
-8  Security Model
-ACP explicitly defines which threats it mitigates, which properties it guarantees, and which risks fall outside its scope.
-8.1  Threat Model (STRIDE)
+8  Threat Model
+ACP explicitly defines which threats it mitigates, which properties it guarantees, and which risks fall outside its scope. For the formal adversary model and security properties, see Section 16.
+8.1  STRIDE Analysis
 Category Threat Mitigation in ACP
@@
 11  Specification and Implementation Status
-ACP v1.17 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
-11.1  Active Specifications — v1.17 (38 documents)
+ACP v1.19 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
+11.1  Active Specifications — v1.19 (38 documents)
 Level Document Title
@@
 Gov. ACP-CR-1.0 Change Request Process
-Table 7: ACP v1.17 active specification documents by conformance level (38 primary documents shown).
+Table 7: ACP v1.19 active specification documents by conformance level (38 primary documents shown).
 The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published.
@@
 New in v1.17: ACR-1.0 sequence compliance runner, TLA+ formal model, ACP-SIGN-2.0 HYBRID stub (pkg/sign2/).
+New in v1.18: performance benchmarks (latency, throughput, state contention), formal security/threat model, system comparison table.
+New in v1.19: adversarial evaluation (cooldown evasion, distributed multi-agent attack, state backend stress with Redis backend, compliance/adversarial/).
 Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0, REP-PORTABILITY-1.0) archived in archive/specs/.
@@
 Benchmark ns/op B/op allocs/op
-Evaluate (APPROVED, no anomaly) 1,012 280 5
-Evaluate (DENIED, no anomaly) 863 296 5
-Evaluate (all 3 F_anom rules) 1,331 296 5
-Evaluate (COOLDOWN short-circuit) 149 112 1
-PatternKey (SHA-256 hash) 996 184 4
-ShouldEnterCooldown 88 0 0
+Evaluate (APPROVED, no anomaly) 767 280 5
+Evaluate (DENIED, no anomaly) 784 296 5
+Evaluate (all 3 F_anom rules) 921 296 5
+Evaluate (COOLDOWN short-circuit) 88 112 1
+PatternKey (SHA-256 hash) 767 184 4
+ShouldEnterCooldown 72 0 0
 […]
@@
 12  Evaluation
-This section evaluates ACP along three dimensions: (i) computational overhead of the decision function, (ii) behavioral correctness under stateful multi-step scenarios, and (iii) reproducibility and external verifiability of decision outcomes.
+This section evaluates ACP along four dimensions: (i) computational overhead of the decision function, (ii) behavioral correctness under stateful multi-step scenarios, (iii) reproducibility and external verifiability of decision outcomes, and (iv) robustness of per-agent admission control under adversarial attack patterns.
 12.1  Evaluation Goals
@@
   • Q3: Can ACP executions be deterministically reproduced and externally verified without access to internal implementation details?
+  • Q4: Does ACP maintain admission control effectiveness under adversarial patterns, including cooldown evasion and distributed multi-agent attacks?
 12.2  Experimental Setup
-All experiments use the Go reference implementation (v1.17) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8).
+All experiments use the Go reference implementation (v1.19) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8).
 The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 13).
@@
 12.3  Q1: Computational Overhead
-Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 1,331 ns. The cooldown short-circuit path (149 ns) bypasses the risk function entirely.
-Answer to Q1: The admission check adds between 149 ns and 1,331 ns of latency per request.
+Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 921 ns. The cooldown short-circuit path (88 ns) bypasses the risk function entirely.
+Answer to Q1: The admission check adds between 88 ns and 921 ns of latency per request.
 Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action.
+12.4  Throughput and State Backend Contention
+We evaluate throughput under concurrent load using 10, 100, and 500 simultaneous workers, all sharing a single InMemoryQuerier instance.
+Workers ns/op req/s Observation
+10 1,087 920,161 Baseline concurrency
+100 1,282 780,250 Moderate contention
+500 1,404 712,259 High contention
+Table 11: ACP throughput under increasing concurrency (Intel Core i7-8665U, Go 1.22, Windows/amd64).
+ACP sustains approximately 920,000 requests per second at 10 concurrent workers.
+Throughput degrades gradually as concurrency increases, reaching approximately 712,000 requests per second at 500 workers.
+Degradation is gradual rather than catastrophic, and no correctness violations were observed under load.
+We further evaluate the impact of agent pool size on throughput (10 workers fixed, AddRequest called per iteration per the execution contract):
+Distinct agents ns/op req/s Bottleneck
+1 84,791 11,794 O($n$) scan + per-key contention
+100 12,533 79,789 State distributed, short lists
+1,000 12,628 79,190 Shared mutex ceiling
+Table 12: State contention analysis: single-agent workload vs. multi-agent workload.
+[…]
+These results indicate that the computational cost of the decision function is negligible compared to state access cost.
+Throughput is primarily bounded by state backend contention, not evaluation complexity.
+The LedgerQuerier interface is designed to be replaced by production backends (e.g., Redis sorted sets with TTL, time-indexed SQL) that eliminate the O($n$) scan penalty and support horizontal scaling.
 12.5  Q2: Stateful Behavior Correctness
@@
 This is a property not typically supported by conventional policy engines, which evaluate policy against a snapshot of state without defining how that state evolves across requests [12, 6, 9].
+12.7  Adversarial Evaluation (ACP-RISK-2.0)
+We evaluate ACP under the adversary model defined in Section 16.
+The following attack strategies instantiate specific adversary capabilities, including adaptive request patterns and multi-agent coordination.
+We conducted three targeted experiments using the same LedgerQuerier abstraction and execution contract as the ACR-1.0 compliance runner, addressing adversarial robustness and real-world state-backend scaling.
+Experiment 1: Cooldown Evasion Attack
+Setup.
+A single agent executes 500 requests alternating between high-risk (acp:cap:financial.transfer, restricted; RS =80=80, outcome DENIED) and low-risk (acp:cap:data.read, public; RS =0=0, outcome APPROVED) requests.
+The adversarial hypothesis is that interleaving approved requests prevents the denial accumulation required to trigger cooldown.
+This attack instantiates the adversary’s ability to adapt request patterns over time.
+Results.
+Metric Value
+Total requests 500
+Approved before cooldown 2
+Real DENIED before cooldown 3
+Cooldown-blocked 495
+Requests processed before first block 5
+Throughput 815,927 req/s
+Table 14: Experiment 1: Cooldown Evasion Attack (1 agent, 500 requests, InMemoryQuerier, Intel i7-8665U, Go 1.22).
+ACP triggered cooldown after exactly 3 real DENIED decisions (requests #0, #2, #4).
+All subsequent requests—regardless of risk score—were short-circuited to COOLDOWN_ACTIVE at Step 2 of the evaluation pipeline ($\sim$88 ns, see Table 10).
+[…]
+Experiment 2: Distributed Multi-Agent Attack
+Setup.
+$N\in\{100,500,1\,000\}$ coordinated agents each execute 10 high-risk requests (acp:cap:financial.transfer, restricted; RS =80=80) using distinct agentID values.
+The adversarial hypothesis is that distributing load across many identities evades per-agent detection thresholds.
+This attack leverages the adversary’s capability to coordinate multiple agents.
+Results.
+Agents Total requests Real DENIED Cooldown-blocked Throughput
+100 1,000 300 700 478,881 req/s
+500 5,000 1,500 3,500 154,111 req/s
+1,000 10,000 3,000 7,000 223,376 req/s
+Table 15: Experiment 2: Distributed Multi-Agent Attack. Each of the NN agents is blocked after exactly 3 DENIED decisions. Total free denials before full blocking =3N=3N.
+Every agent was individually blocked after exactly CooldownTriggerDenials =3=3 real denials.
+Interpretation.
+Per-agent admission control provides strong containment for single-agent threats.
+For distributed attacks, an attacker with $N$ distinct agent identities can execute $3N$ high-risk requests before all agents are individually blocked.
+Mitigating this requires cross-agent attribution at the policy layer (e.g., shared-resource rate limiting, identity clustering), which is deliberately outside the scope of ACP’s per-agent admission model.
+This is a design boundary, not a correctness failure.
+Experiment 3: State Backend Stress
+Setup.
+500 agents execute 20 requests each (10,000 total, fully concurrent) against two backend implementations: the reference InMemoryQuerier (Go mutex over in-process slices) and a minimal RedisQuerier (Redis 7 sorted sets via go-redis/v9, Docker loopback, one Redis command per ledger operation, no pipelining).
+This scenario exercises the adversary’s ability to generate high request volumes and exploit shared state contention.
+Results.
+Backend Duration Throughput
+InMemoryQuerier $\sim$28 ms $\sim$350,000 req/s
+RedisQuerier $\sim$4.6 s $\sim$2,100 req/s
+Table 16: Experiment 3: State Backend Stress (500 agents $\times$ 20 requests =10,000=10{,}000 total).
+InMemoryQuerier variance is high under concurrent load ($\pm$30%) due to Go mutex scheduling; RedisQuerier is stable ($\pm$4%).
+Interpretation.
+The lower throughput of InMemoryQuerier under adversarial load ($\sim$350k req/s) compared to the 500-worker baseline of $\sim$712k req/s (Table 11) is expected: the adversarial workload generates significantly higher contention because every request updates multiple sorted lists under a shared mutex.
+RedisQuerier eliminates the global Go mutex entirely, trading it for per-operation network RTT.
+A production-grade Redis backend using command pipelining or server-side Lua scripts would reduce latency to a single round-trip per request.
+The key architectural result remains: ACP protocol semantics are stable across backends; deployment performance scales with the quality of the state backend, not with the admission control logic itself.
+These experiments address the adversarial-robustness and scaling gaps identified in Section 12. All scripts, raw output, and Redis configuration are available in compliance/adversarial/.
+Security Properties Under Adversarial Stress
+We map the evaluated attack strategies to the security properties defined in Section 16, analyzing whether each property holds under adversarial conditions.
+Determinism.
+Decision outcomes remain consistent under identical inputs and state conditions, including under concurrent execution and adversarial request patterns.
+Cooldown Enforcement.
+Attack 1 (cooldown evasion via pattern alternation) directly targets the cooldown mechanism.
+Results show that cooldown activation cannot be indefinitely avoided: adversarial alternation delays but does not prevent enforcement.
+This confirms that cooldown is eventually enforced under repeated violations.
+Bounded Adaptation Resistance.
+Attack 1 also exercises adaptive behavior.
+The experiment shows that the denial counter evolves monotonically under repeated violations: APPROVED requests do not reset prior denials.
+As a result, adversarial interleaving strategies cannot asymptotically evade enforcement, establishing a bounded adaptation property.
+Per-Agent Isolation.
+Attack 2 (distributed multi-agent attack) targets system-wide coordination.
+Results show that ACP enforces constraints independently per agent, maintaining isolation.
+However, coordinated behavior across agents is not detected, which constitutes an explicit design limitation rather than a violation.
+State Sensitivity Under Load.
+Attack 3 (state backend contention) evaluates system behavior under adversarial load.
+While throughput degrades due to shared state contention, decision correctness and enforcement properties remain unaffected.
+Summary.
+The evaluated attacks confirm that ACP enforces its core security properties under adversarial conditions.
+No violations are observed; instead, limitations arise from explicit design trade-offs, particularly in cross-agent correlation.
+These results reinforce the central design insight: ACP is compute-cheap but state-sensitive.
 12.8  Limitations
-The evaluation uses an in-memory state backend.
-Persistent storage, distributed deployments, and concurrent request patterns introduce latency and consistency considerations not measured here.
-Adversarial robustness under coordinated attack scenarios is not evaluated.
-These are identified as future work (Section 12 roadmap).
+The evaluation uses an in-memory state backend for latency benchmarks and a single-node Docker Redis instance for backend comparison.
+Persistent storage, distributed deployments with replication, and wide-area network latency introduce considerations not measured here.
+12.9  Summary
+[…]
 12.10  Roadmap
@@
 ACP-SIGN-2.0 HYBRID stub (pkg/sign2/) v1.17 Complete
-Post-quantum Go implementation (Dilithium, circl) v1.18
+Performance benchmarks (latency, throughput, state contention) v1.18 Complete
+Adversarial evaluation (cooldown evasion, distributed multi-agent, state backend stress) v1.19 Complete
+Post-quantum Go implementation (Dilithium, circl) v1.20
 L5 Decentralized (ACP-D) Specification in design (v2.0)
 IETF RFC submission After L5 stabilization
-Table 12: ACP v1.17 development roadmap.
+Table 17: ACP v1.19 development roadmap.
 13  Compliance Testing (ACR-1.0 Sequence Runner)
@@
 14  End-to-End Verifiability
-ACP v1.17 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction.
+ACP v1.19 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction.
 Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
@@
 Each TLA+ invariant is instantiated as a concrete test vector and validated empirically by the compliance runner.
-This is the connection across the three layers of the v1.17 verifiability stack: formal $\to$ data $\to$ runtime.
+This is the connection across the three layers of the v1.19 verifiability stack: formal $\to$ data $\to$ runtime.
 15  How to Implement ACP
@@
 It does not prescribe the how of internal implementation: programming language, database, HSM provider, ITA provider, or specific integration with existing RBAC or Zero Trust systems.
+16  Security Model
+This section defines the adversary model and the security properties provided by ACP.
+16.1  Adversary Capabilities
+We consider a probabilistic adversary $\mathcal{A}$ interacting with the system through agent requests. The adversary is assumed to have the following capabilities:
+  • Request control: $\mathcal{A}$ can generate arbitrary sequences of requests on behalf of one or more agents, including high-frequency and adversarial patterns.
+  • Adaptive behavior: $\mathcal{A}$ can adapt future requests based on past responses (approval/denial outcomes), enabling iterative probing strategies.
+  • Multi-agent coordination: $\mathcal{A}$ may coordinate multiple agents to attempt distributed evasion, including alternating identities or cross-agent patterns.
+  • Input manipulation: $\mathcal{A}$ can choose all request fields (capability, resource, context) within the bounds of the protocol specification.
+16.2  Adversary Limitations
+The following components are trusted and cannot be modified by the adversary:
+  • Execution contract: the sequence of operations governing state updates (evaluation followed by state transitions) is enforced correctly.
+  • State integrity: the underlying state backend (ledger, cooldown state, pattern counters) is append-only and tamper-resistant.
+  • Evaluation function: the decision function (ACP-RISK-2.0) is deterministic and correctly implemented.
+  • Cryptographic primitives: signature verification and identity binding are assumed secure under standard assumptions (EUF-CMA security of Ed25519).
+This model corresponds to an adversary that can fully control inputs but cannot compromise the integrity of the control system itself.
+16.3  Security Properties
+Under this model, ACP enforces the following properties:
+Determinism.
+For a given request and system state, the evaluation function produces a unique and reproducible decision outcome.
+This prevents adversarial exploitation of non-determinism.
+The RiskDeterminism invariant in the TLA+ model formally specifies this property.
+State Consistency.
+All state transitions follow the execution contract, ensuring that historical behavior is consistently reflected in future decisions.
+Cooldown Enforcement.
+An adversary cannot bypass cooldown mechanisms through repeated or adaptive requests.
+Once cooldown conditions are met, subsequent requests are deterministically denied within the cooldown window, regardless of request content.
+This is confirmed by the SEQ-COOLDOWN-001 sequence test vector.
+History Integrity.
+Past decisions and events cannot be modified or removed.
+The append-only ledger (ACP-LEDGER-1.3) and the LedgerAppendOnlyTemporal TLA+ invariant formally guarantee monotonic accumulation of historical state.
+Bounded Adaptation Resistance.
+While $\mathcal{A}$ may adapt request patterns, such adaptations cannot eliminate the influence of accumulated history (repeated denials, pattern counts) on future decisions.
+In particular, the denial counter evolves monotonically: intervening approved requests do not reset prior denials, ensuring eventual enforcement under repeated violations.
+Per-Agent Isolation.
+Policy enforcement is applied independently per agent.
+Actions by one agent do not affect the enforcement state of another.
+This property bounds the impact of a compromised or malicious agent to its own request history.
+16.4  Informal Security Argument
+Because ACP separates stateless evaluation from state management and enforces a strict execution contract, every request results in a deterministic state transition that cannot be bypassed.
+The append-only nature of the state backend ensures that historical information accumulates monotonically: cooldown activation follows deterministic threshold conditions applied to accumulated state; repeated probing eventually triggers enforced denial periods.
+Because decisions are reproducible given the same input sequence and initial state, any observed behavior can be externally verified.
+This limits $\mathcal{A}$’s ability to exploit hidden or undocumented system behavior.
+16.5  Limitations
+The security guarantees of ACP rely on the integrity of the execution environment:
+  • Compromise of the state backend (deletion or modification of history) breaks core guarantees.
+  • Incorrect implementation of the execution contract may lead to inconsistent behavior.
+  • The model does not address side-channel attacks or network-level adversaries.
+  • Post-quantum signature migration (ACP-SIGN-2.0 HYBRID, Ed25519 $\to$ ML-DSA-65) is specified and stubbed; full implementation is deferred to v1.20.
+ACP is therefore best understood as a control-layer mechanism that assumes a trusted execution and storage environment, while providing strong guarantees on decision consistency and temporal behavior under adversarial inputs.
 17  Related Work
@@
   • externally verifiable execution traces that can be reproduced by third parties.
+Recent work has begun to address security governance for LLM-based agentic systems.
+Syros et al.
+[15] propose SAGA, a security architecture that provides user-controlled access policies and cryptographic mediation for agent-to-agent tool use and inter-agent communication.
+ACP complements this line of work by focusing on deterministic, auditable admission control at the individual request level: rather than governing inter-agent communication channels, ACP enforces stateful risk evaluation and cooldown constraints on each action request, with an externally verifiable execution contract and formal invariants.
 ACP complements these approaches by providing a governance layer that is independent of orchestration protocols, focusing specifically on verifiable decision enforcement and auditability.
@@
 17.5  Summary
+Table 20 summarizes the properties of ACP relative to existing approaches.
+Property RBAC ABAC OPA Cedar MCP/A2A ACP
+Stateful decisions $\times$ $\times$ $\times$ $\times$ $\times$
+Deterministic evaluation $\times$
+Auditability partial partial partial partial $\times$
+Runtime enforcement partial
+Temporal behavior $\times$ $\times$ $\times$ $\times$ $\times$
+External verifiability $\times$ $\times$ $\times$ $\times$ $\times$
+Execution contract $\times$ $\times$ $\times$ $\times$ $\times$
+Agent-native $\times$ $\times$ $\times$ $\times$ partial
+Table 20: Comparison of ACP with existing access control and agent governance approaches. ✓ = fully supported; $\times$ = not supported; partial = partially addressed.
 Existing systems address individual aspects of the problem—policy evaluation, agent orchestration, audit logging, or formal verification.
@@
 This enables a property not directly provided by prior systems: the ability to deterministically reproduce and externally verify the behavior of an admission decision system over time.
+Unlike existing systems, ACP explicitly models temporal behavior—history accumulation, anomaly detection, and cooldown enforcement—within the admission decision itself, and provides verifiable execution semantics through signed test vectors and a runtime compliance runner.
+This combination of temporal state, runtime enforcement, and external verifiability is not jointly addressed by any prior system in Table 20.
 18  Conclusion
@@
 What it provides is a structured, auditable, and formally specified model for admission control, delegation, revocation, and cross-organization interaction— the layer that existing identity and policy frameworks leave unaddressed.
-The v1.17 specification is complete.
+The v1.19 specification is complete.
 A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (3 invariants, 0 violations), an ACP-SIGN-2.0 HYBRID mode stub with Ed25519 and ML-DSA-65 migration path, an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent killer demo (examples/payment-agent/), and an extended OpenAPI 3.1.0 specification (18 endpoints) are publicly available at https://github.com/chelof100/acp-framework-en.
-Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19198298
+Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19219776
 The specification and implementation are open for technical review, pilot implementation, and formal standardization.
 TraslaIA invites organizations interested in adopting ACP, contributing to its evolution, or participating in the standardization process to reach out directly.
+ACP demonstrates that admission control for autonomous agents can be both computationally efficient and operationally enforceable—provided that decision logic and state management are cleanly separated.
+The protocol is compute-cheap but state-sensitive: the LedgerQuerier abstraction keeps the evaluation function at $\sim$820 ns while allowing the state backend to be replaced, optimized, or scaled independently.
+This is not an implementation detail—it is the architectural property that makes ACP deployable in production without sacrificing governance correctness.
 Marcelo Fernandez | TraslaIA
@@
 RiskDeterminism is the central invariant of ACP-RISK-2.0: identical (capability, resource) pairs always produce identical risk scores, enabling third-party recalculation from a signed policy snapshot.
+Scope of the formal model.
+The TLA+ model intentionally focuses on the deterministic core of the protocol—admission evaluation, ledger append-only property, and risk score determinism—and does not cover all operational aspects of ACP.
+Behavioral properties such as cooldown state transitions, multi-step delegation sequences, and anomaly counter evolution are verified empirically through the ACR-1.0 sequence compliance runner and its five stateful test vectors.
+Extension of the formal model to cover these behaviors is left for future work.
 Acknowledgements

### 30 Apr 2026 v10

@@
-Agent Control Protocol ACP v1.19: Admission Control for Agent Actions Technical Specification and Reference Implementation
+Agent Control Protocol ACP v1.30: Admission Control for Agent Actions Technical Specification and Reference Implementation
 Abstract
-[…]
+Autonomous agents can produce harmful behavioral patterns from individually valid requests—a threat class that per-request policy evaluation cannot address, because stateless engines evaluate each request in isolation and cannot enforce properties that depend on execution history.
+We present ACP, a temporal admission control protocol that enforces behavioral properties over execution traces by combining static risk scoring with stateful signals (anomaly accumulation, cooldown) through a LedgerQuerier abstraction separating decision logic from state management.
+ACP is not an anomaly detection system: it does not detect suspicious patterns and alert; it blocks execution based on deterministic, history-aware risk scoring, providing a hard enforcement boundary rather than an advisory signal.
+We demonstrate that this structural separation has measurable security consequences: under a 500-request workload where every request is individually valid (RS=35), a stateless engine with identical scoring approves all 500 requests, while ACP limits autonomous execution to 2 out of 500 (0.4%), escalating after 3 actions and enforcing denial after 11—isolating the structural gap between stateless and stateful admission.
+We identify a bounded state-mixing vulnerability in ACP-RISK-2.0: agents executing high-frequency benign operations in one context can inadvertently elevate risk scores in unrelated high-value contexts due to agent-level rate aggregation—producing false denials that a stateless engine would never generate (Experiment 7).
+We introduce ACP-RISK-3.0, which eliminates this cross-context interference by scoping rate-based anomaly signals to the interaction context via PatternKey(agentID, capability, resource), while preserving enforcement: repeated behavior within a single context continues to trigger denial at the same threshold (Experiment 8).
+Decision evaluation runs at 739–832 ns (p50; Table 12); throughput reaches 1,720,000 req/s (baseline, InMemoryQuerier) and 920,000 req/s at 10 concurrent workers (Table 13), degrading predictably with backend latency without protocol changes.
+Safety and liveness properties are model-checked via TLC-runnable TLA+ (11 invariants + 4 temporal properties, 0 violations); two-agent safety is verified at the correct ledger bound across 4,294,930,695 distinct states with zero violations.
+Runtime correctness is validated by 73 signed conformance test vectors.
+We demonstrate that an adversary with complete knowledge of the risk formula can suppress BAR to $0.00$ while every individual request remains policy-compliant (Experiment 10); BAR-Monitor detects the regime shift via $\Delta\mathrm{BAR}$ three batches before collapse, and counterfactual evaluation confirms structural enforcement capacity is preserved ($\mathrm{BAR}_{C}=1.00$).
+We further identify and formalize deviation collapse—a previously unrecognized governance failure mode in which enforcement is active but never exercised because upstream constraints eliminate the conditions required for DENIED decisions—and introduce Boundary Activation Rate (BAR) as its detection mechanism (Experiment 9).
+We evaluate coordinated multi-agent behavior (Experiment 13): $N$ agents issuing homogeneous requests on a shared resource each accumulate risk independently via agent-scoped PatternKey and ACP’s evaluate-then-mutate execution contract, guaranteeing each agent exactly two approved actions before escalation and denial.
+The total coordination window $\mathrm{CW}_{\mathrm{appr}}=2N$ with zero deviation across all configurations: coordinated activity scales exactly linearly, preventing superlinear amplification.
+Together, Experiments 9, 13, and TLA+ verification demonstrate that ACP constrains both the existence and the scale of undesirable behaviors.
+ACP is Paper 1 of a 6-paper Agent Governance Series: P0—atomic decision boundaries [14]; P2—behavioral drift detection (IML) [15]; P3/4—governance structure, fair allocation, and irreducibility [16]; P5—runtime execution validity under partial observability (RAM) [fernandez2026ram]; P6—operationalization of RAM as a runtime enforcement protocol [fernandez2026op].
+All companion papers are published on Zenodo.
 Specification and implementation: https://github.com/chelof100/acp-framework-en
 1  The Problem ACP Solves
-Autonomous agents are being deployed in institutional environments without a technical standard to govern their behavior. This is not a tooling problem—it is a protocol problem.
+Autonomous agents are being deployed in institutional environments without a technical standard to govern their behavior. This gap is structural, not merely a tooling limitation.
 1.1  The Structural Gap
@@
 It can only be audited—and only if there is something to audit.
-The problem is not whether agents are trustworthy.
-The problem is that currently no formal technical mechanism exists that allows an institution to demonstrate that its agents operated within authorized limits.
+The fundamental problem is therefore not whether agents are trustworthy, but that no standardized, protocol-level mechanism currently exists that allows an institution to demonstrate that its agents operated within authorized limits.
+Autonomous agents differ from traditional software services in one critical respect: their action sequences emerge from internal reasoning processes rather than from a pre-specified function.
+A conventional service executes a declared operation in response to a request.
+An autonomous agent may produce a sequence of actions whose security relevance only becomes apparent across multiple steps—a single request provides insufficient information to determine whether the agent is operating within authorized behavioral bounds.
+ACP enforces constraints over execution traces rather than individual requests.
+Admissibility is resolved at execution time, but it is not memoryless.
+This enables deterministic verification of temporal behavioral properties: whether an agent has exceeded its authorized request rate within a time window, whether a delegation chain is being exercised beyond its intended scope, whether an anomaly pattern is accumulating across a session.
+These properties depend on accumulated interaction history and cannot be reduced to single-request evaluation.
+They correspond to the class of trace-based safety properties formalized by Schneider [42], which stateless policy engines cannot enforce by construction.
+No adjustment of scoring thresholds or risk weights can compensate for this absence: an engine that evaluates each request in isolation lacks the information necessary to determine whether the current request is part of an anomalous sequence.
+The requirement is state, not configuration—a structural constraint independent of the sophistication of the scoring function.
+Existing policy engines and authorization frameworks—OPA [36], RBAC [41], capability systems—evaluate each request in isolation against declared policy.
+ACP does not replace them; it adds a stateful enforcement layer above them.
+Each admission decision is computed as a function of the interaction trace, not only the current request—a composition that extends the practical enforcement boundary from individual interactions to behavioral sequences.
+Governance of autonomous agents therefore involves three distinct requirements.
+First, admissibility must be resolved at the execution boundary—before an action commits, not after [3].
+Second, the state evaluated at that boundary must encode execution history: anomaly accumulation, action frequency, and prior behavioral patterns.
+Without this, systems cannot distinguish isolated valid actions from sequences that become inadmissible over time.
+Third, the system must preserve the conditions under which inadmissibility can arise.
+A system that always produces admissible actions does not demonstrate effective governance; it may be operating in a regime where the boundary is intact but never exercised.
+In this work, we address all three requirements.
+ACP satisfies the first two by design; we identify and formalize the third as failure condition preservation (Section 17).
 1.2  ACP as Admission Control
-The clearest frame for understanding what ACP does is the Kubernetes Admission Controller analogy.
+A useful analogy for understanding what ACP does is the Kubernetes Admission Controller pattern [28].
 Kubernetes intercepts every API request before it reaches the cluster and runs it through a sequence of admission checks—ValidatingWebhookConfiguration, ResourceQuota enforcement, OPA Gatekeeper policies.
@@
 | does the agent hold a valid token for this action?
-[3] Policy check (ACP-RISK-2.0, ACP-PSN-1.0)
+[3] Policy check (ACP-RISK-3.0, ACP-PSN-1.0)
 | is this action within current policy?
@@
   • SPIFFE / SPIRE [44] provides cryptographic workload identity. ACP builds on that identity to add capability scoping and governance.
+Viewed through the lens of classical security theory, ACP is a decentralized reference monitor: a component that mediates every agent action, is always invoked before execution, and produces a tamper-evident record of every admission decision—extending Anderson’s reference monitor concept [3] to multi-organizational environments where no single trusted party controls all agents.
 1.3  Why RBAC and Zero Trust Are Insufficient
@@
 ACP is compute-cheap but state-sensitive: decision evaluation is computationally inexpensive, while system performance is primarily constrained by state access patterns.
-This separation enables a protocol that is both efficient in isolation ($\sim$820 ns per decision, $\sim$88 ns for the cooldown short-circuit path) and adaptable in deployment: performance can be improved by replacing the state backend without modifying protocol semantics.
+This separation enables a protocol that is both efficient in isolation (739–832 ns per decision, p50; Table 12, $\sim$78 ns for the cooldown short-circuit path) and adaptable in deployment: performance can be improved by replacing the state backend without modifying protocol semantics.
 The LedgerQuerier interface formalizes this boundary—the decision function remains stateless and deterministic; the state backend can be a local in-memory structure for development or a distributed, indexed store in production.
 1.6  Contributions
-This paper makes the following contributions:
-  • Admission control protocol.
-A deterministic admission control protocol for autonomous agents combining capability, resource, historical, and anomaly-based signals into a single verifiable decision.
-  • State-aware execution model.
-An execution model that explicitly separates decision logic from state management via the LedgerQuerier abstraction, enabling deterministic reasoning about temporal agent behavior.
+This paper makes the following contributions, organized by theme:
+I. Stateful temporal admission control.
+  • Structural necessity of stateful enforcement.
+We demonstrate empirically (Experiment 6) that temporal behavioral properties over execution traces cannot be enforced by stateless policy engines regardless of scoring function or thresholds.
+Under a 500-request financial transfer workload where every request is individually valid, a stateless engine with identical scoring approves all 500 requests; ACP limits autonomous execution to 2 out of 500 (detection latency: 11 actions, early signal: 3 actions).
+  • Sub-microsecond temporal admission control.
+We show that stateful trace-level enforcement is compatible with 739–832 ns decision latency (p50; Table 12) and 920,000 req/s throughput, achieved by separating decision logic (stateless, compute-cheap) from state management via the LedgerQuerier abstraction—enabling backend replacement without modifying protocol semantics or security guarantees.
 […]
-  • End-to-end verifiability.
-A three-layer verifiability pipeline: formal (TLA+, TLC-checked invariants), data (signed conformance test vectors), and runtime (ACR-1.0 compliance runner), enabling independent validation without proprietary infrastructure.
-[…]
+II. Formal verification and verifiability.
+  • Formal verification of safety invariants.
+Safety properties (admission determinism, cooldown monotonicity, per-agent isolation) are verified by TLC model checking over 5,684,342 generated states in the single-agent configuration (11 invariants + 4 temporal properties, 0 violations) and over 4,294,930,695 distinct states in a two-agent configuration at LEDGER_BOUND=11 (11 safety invariants, 0 violations, 10.5 h, i9-13900HX 24-core, 24 GB heap), establishing that enforcement holds under adversarial state conditions and across independent concurrent agents.
+  • End-to-end verifiability without proprietary infrastructure.
+A three-layer pipeline—73 signed conformance test vectors, TLC-checked TLA+ model, and ACR-1.0 runtime sequence runner—enables independent validation of both specification and reference implementation.
+III. Risk scoring and adversarial robustness.
+  • Context-scoped anomaly model eliminating cross-context interference.
+We introduce ACP-RISK-3.0, which redefines Rule 1 to use context-scoped pattern counts, eliminating cross-context state-mixing while preserving enforcement against repeated behavior within a given interaction context (Experiment 8, Section 13.7).
+  • Knowledge-aware adversarial evasion (Experiment 10).
+We demonstrate that an adversary with complete knowledge of the ACP-RISK-3.0 parameters can suppress BAR to $0.00$ while every individual request remains policy-compliant — per-decision enforcement is structurally blind to this attack.
+BAR-Monitor detects the regime shift via $\Delta\mathrm{BAR}$ at Batch 2, three batches before threshold collapse; EvaluateCounterfactual confirms structural enforcement capacity is preserved ($\mathrm{BAR}_{C}=1.00$).
+These three mechanisms form a coherent defense against knowledge-aware evasion.
+IV. Governance: deviation collapse and boundary monitoring.
+  • Identification of deviation collapse as a governance failure mode.
+We demonstrate empirically that a system may remain fully compliant while its admissibility boundary is never exercised ($\mathrm{BAR}_{A}=0.70\to\mathrm{BAR}_{B}=0.00$, $n=20$).
+We introduce Boundary Activation Rate (BAR) as a metric for detecting this condition and propose counterfactual evaluation as a restoration mechanism (Experiment 9, Section 17).
+  • Failure condition preservation as a governance requirement.
+We establish failure condition preservation as a necessary requirement for effective admission control: a system must not only enforce admissibility correctly but must preserve the conditions under which enforcement can fire.
+This separates correct enforcement from effective governance and provides the theoretical foundation for deviation collapse detection (Section 17).
+  • Boundary Activation Monitoring (BAR-Monitor).
+We introduce BAR-Monitor (pkg/barmonitor), a mechanism that detects loss of boundary interaction by tracking both $\mathrm{BAR}_{N}$ (current activation level) and $\Delta\mathrm{BAR}$ (progressive decline toward collapse).
+BAR-Monitor provides the lower-bound complement to cooldown’s upper-bound constraint, defining a bounded operational region in which governance remains both safe and meaningful (Section 17.6).
+  • Empirical $\Delta\mathrm{BAR}$ early-warning validation (Phase D).
+We validate $\Delta\mathrm{BAR}$ as a proactive detection mechanism through a controlled drift simulation (Experiment 9 Phase D, 5 batches, 0–100% upstream sanitization).
+AlertTrend fires at Batch 2 ($\mathrm{BAR}=0.57$, $\Delta\mathrm{BAR}=-0.25$), three batches before the threshold condition confirms collapse ($\mathrm{BAR}=0.00$ at Batch 5), demonstrating a measurable early-warning gap.
+  • TLA+ structural failure condition preservation.
+We extend the TLA+ model with two invariants — FailureConditionPreservation and NoDegenerateAdmissibility — that establish necessary structural conditions for failure condition preservation.
+Both are verified by TLC with zero violations over all reachable states.
+BARMonitorLiveness is stated formally and validated empirically (Phase D).
+Positioning.
+The enforcement design introduced in this work—in which evaluation and state transition are resolved jointly at the admission boundary—can be understood as an instance of a broader class of stateful enforcement systems.
+A formal characterization of this structural property, and its extension to behavioral invariants beyond the admission decision, are developed in subsequent work.
+2  Related Work
+2.1  Access Control and Policy Enforcement
+Traditional access control models such as Role-Based Access Control (RBAC) [41] and Attribute-Based Access Control (ABAC) [24] provide structured mechanisms for regulating access decisions based on predefined roles or attributes.
+Policy languages such as XACML [34] and modern policy engines such as Open Policy Agent [36] and Cedar [2] extend these paradigms with expressive evaluation engines.
+These systems are fundamentally designed around stateless decision evaluation: they evaluate a request against policy at a single point in time.
+They do not define how system state evolves as a result of decisions over time, and they do not formalize how historical behavior (repeated denials, temporal patterns) influences subsequent decisions in a deterministic and auditable manner.
+ACP differs by explicitly separating: (i) stateless decision evaluation, (ii) externalized state management, and (iii) an execution contract governing state evolution across requests.
+This enables reasoning not only about individual decisions, but about the temporal behavior of the admission system.
+2.2  Runtime Policy Systems and Agent Frameworks
+Recent protocols for autonomous agents and multi-agent coordination define structured interactions between components and services.
+The Model Context Protocol (MCP) [4] provides structured tool access between LLM applications and services.
+Agent-to-Agent (A2A) [21] defines communication and task delegation between agents.
+While these protocols address interoperability and orchestration, they do not define:
+  • a deterministic risk evaluation model that accumulates state across requests,
+  • explicit state evolution semantics with a testable execution contract, or
+  • externally verifiable execution traces that can be reproduced by third parties.
+Recent work has begun to address security governance for LLM-based agentic systems.
+Syros et al.
+[47] propose SAGA, a security architecture that provides user-controlled access policies and cryptographic mediation for agent-to-agent tool use and inter-agent communication.
+While SAGA focuses on high-level governance architecture and inter-agent communication policies, ACP enforces admission constraints at the execution layer: every action request passes through a stateful risk evaluation with deterministic outcomes, a verifiable execution contract (ACR-1.0), and a tamper-evident ledger— providing enforcement rather than governance framing.
+Tsai and Bagdasarian [48] argue at HotOS 2025 that security policies must be purpose-specific rather than uniformly applied, emphasising that one-size-fits-all approaches are inadequate for diverse agent deployments—a perspective consistent with ACP’s per-agent isolation and stateful enforcement boundaries.
+ACP complements these approaches by providing a governance layer that is independent of orchestration protocols, focusing specifically on verifiable decision enforcement and auditability.
+2.3  Auditability and Tamper-Evident Logs
+Secure audit logging [43] and event sourcing [18] ensure that past events can be reconstructed and verified after the fact.
+These approaches operate post hoc: they provide evidence of what occurred without constraining system behavior in real time.
+ACP integrates auditability directly into the admission process.
+Every decision—APPROVED, ESCALATED, or DENIED—is recorded in an append-only, cryptographically chained ledger (ACP-LEDGER-1.3) before the governed action executes.
+This enables both real-time enforcement and post-hoc verification from the same artifact.
+The LedgerAppendOnlyTemporal invariant in the TLA+ model formally specifies that ledger entries are never modified or removed.
+2.4  Formal Verification and Runtime Enforcement
+Model checking [29] and symbolic reasoning are used to establish correctness properties in distributed systems and protocols, typically focusing on safety and liveness of state transitions.
+The theoretical foundations of runtime enforcement trace to Anderson’s reference monitor concept [3] and Schneider’s formal characterization of enforceable security policies [42], which established the boundary between properties that can be verified post-hoc and those that can be enforced at the point of execution via a security automaton.
+Falcone et al.
+[11] systematize runtime enforcement monitors, distinguishing enforceable property classes (safety, co-safety, guarantee, persistence) and the suppression or insertion mechanisms each requires.
+Stateless policy engines such as OPA [36] evaluate each request independently, without intrinsic support for stateful behavioral accumulation across execution traces.
+OPA has become a widely adopted standard for policy enforcement in distributed systems, offering a declarative evaluation language (Rego) and integration points across APIs, admission controllers, and service meshes.
+These properties make it well-suited for request-level authorization, but they also define a structural boundary: enforcement of constraints that depend on prior actions, frequency patterns, or temporal context requires externalizing state to a separate store and threading it through every evaluation call.
+This distinction is critical for agentic systems.
+Autonomous agents generate sequences of actions whose admissibility depends not only on the current input, but on the history of prior actions, frequency patterns, and temporal constraints—properties that cannot be expressed in stateless evaluation without introducing external infrastructure.
+In practice, approximating stateful behavior in OPA requires maintaining counters and timers in an external store (e.g., Redis or a database), explicitly injecting current state into each evaluation request, and relying on the surrounding architecture to guarantee atomicity between evaluation and state mutation.
+This shifts correctness responsibility outside the policy engine itself, increasing architectural complexity and weakening enforcement guarantees at the decision boundary.
+Our experimental evaluation (Experiment 17.12) illustrates this boundary directly.
+Under repeated identical inputs—ten financial transfer requests from a single agent—OPA remains permissive across all iterations in its standard stateless configuration (allow=true $\times$ 10).
+When augmented with externally injected request counts, OPA can approximate frequency-based constraints, but only through additional infrastructure outside the engine, and without atomicity guarantees between evaluation and the counter update.
+ACP enforces the same constraints natively: requests transition from Approved to Escalated to Denied via accumulated PatternKey counts maintained within the execution contract, without external coordination.
+This difference reflects a distinction in expressiveness.
+Stateless policy engines operate over representations of individual requests; ACP operates over execution traces.
+As a result, ACP can enforce constraints that depend on the existence, frequency, and ordering of actions over time—properties that stateless engines, by construction, cannot enforce without external state integration [42].
+ACP instantiates a stateful enforcement mechanism aligned with Schneider’s security automaton model, enabling formal reasoning about temporal properties of agent behavior that stateless evaluation cannot capture.
+ACP instantiates these principles in a decentralized, multi-organizational setting.
+The ACR-1.0 compliance runner acts as an enforcement monitor over the ACP execution contract: every request passes through the monitor before state mutation occurs, the monitor is always invoked (no bypass path), and its decisions are recorded in a tamper-evident ledger.
+The TLA+ model specifies the safety and liveness properties the monitor is designed to enforce; formally checked invariants are then instantiated as concrete sequence test vectors and validated at runtime by the compliance runner.
+This creates a three-layer verifiability chain: formal $\to$ data $\to$ runtime.
+2.5  Capability Systems and Proof-of-Possession Mechanisms
+ACP builds upon established primitives from capability-based security [31] and proof-of-possession authentication, but differs in how these primitives are composed and enforced over execution traces.
+Macaroons.
+Macaroons [6] introduce decentralized authorization tokens with embedded caveats that can be attenuated and verified without contacting a central authority.
+ACP-CT-1.0 shares structural similarities with Macaroons in that tokens carry verifiable constraints and can be evaluated without implicit trust in the caller.
+However, Macaroons are fundamentally stateless: all constraints must be encoded within the token itself, and verification does not consult external state.
+They do not support enforcement of temporal or behavioral properties that depend on prior actions across requests.
+In contrast, ACP separates token validation from behavioral enforcement: ACP-CT ensures capability integrity per request, while ACP-RISK evaluates actions against accumulated execution traces via a stateful backend, enforcing constraints that cannot be expressed as static caveats.
+ZCAP-LD.
+ZCAP-LD [45] extends object-capability models to decentralized systems using linked data proofs, enabling delegation chains and fine-grained authorization over distributed resources.
+Similar to Macaroons, ZCAP-LD focuses on authorization semantics and delegation but does not define mechanisms for temporal enforcement or anomaly detection across sequences of actions; policy evaluation remains scoped to individual invocations.
+ACP operates at a different layer: it assumes a valid capability system (such as ACP-CT or alternatives) and adds a stateful admission control layer that evaluates behavior over time, remaining compatible with existing capability frameworks while extending their enforcement model.
+DPoP.
+DPoP [17] introduces a mechanism for binding access tokens to a client-held private key, ensuring that tokens cannot be replayed by unauthorized parties.
+ACP-HP-1.0 is aligned with DPoP in enforcing proof-of-possession per request; however, DPoP operates at the transport/session level within OAuth flows.
+ACP extends this model to action-level enforcement: each action is cryptographically bound and independently verifiable, enabling replay resistance that is both cryptographic (invalid proof) and behavioral (anomalous repetition patterns detected via F_anom Rule 3, as shown in Experiment 5).
+Summary.
+These systems provide strong primitives for authorization and authentication but evaluate requests in isolation.
+ACP does not introduce new cryptographic primitives; its contribution lies in composing existing mechanisms into a stateful admission control model that enforces constraints over execution traces rather than individual requests—the structural difference empirically demonstrated in Experiment 6.
+2.6  Behavioral Failure Modes in Governed Systems
+Two bodies of work share surface-level similarity with deviation collapse but address structurally distinct failure modes.
+Specification gaming [27] occurs in reinforcement learning when an agent maximizes a reward function while violating its intended objectives—satisfying the letter of a constraint while undermining its purpose.
+The locus of failure is the objective specification: the agent behaves optimally under the given reward signal, which itself is misaligned with the designer’s intent.
+Deviation collapse operates at a different layer: the admission constraint is correctly specified and correctly enforced, but upstream filtering prevents risk-bearing requests from reaching the enforcement boundary.
+The ACP engine does not optimize against the policy; it evaluates faithfully.
+The failure is architectural, not behavioral.
+Goodhart’s Law [20, 46]—“when a measure becomes a target, it ceases to be a good measure”—describes how optimization pressure corrupts proxy metrics.
+Deviation collapse shares the surface form of a metric losing signal, but the mechanism differs: no optimization pressure is applied to the Boundary Activation Rate, and no party is gaming the BAR metric.
+The collapse occurs because the upstream pipeline removes the inputs that would activate the boundary—a structural consequence of pipeline composition, not metric manipulation.
+2.7  LLM Agent Security and Runtime Enforcement
+A growing body of work examines security threats specific to LLM-integrated agents.
+Greshake et al.
+[22] define indirect prompt injection (IPI), in which adversarial instructions embedded in data retrieved by an agent override intended behavior, enabling exfiltration, unauthorized API calls, and cross-agent propagation.
+Zhan et al.
+[51] and Debenedetti et al.
+[8] provide benchmark evaluations showing that model-level defenses alone are insufficient to contain IPI: frontier models remain vulnerable across diverse task environments.
+ACP operates at the execution layer, not the prompt layer.
+Prompt injection and jailbreaking occur before an action request reaches ACP’s admission boundary; ACP intercepts the resulting tool or API call before it modifies system state.
+The two defense layers are therefore composable rather than competitive: a prompt-layer filter reduces the IPI attack surface; ACP enforces structural constraints over the action sequence regardless of how the request was generated.
+Wang et al.
+[49] propose AgentSpec, a domain-specific language for defining runtime safety constraints on LLM agents, sharing the architectural insight that enforcement must be interposed outside the model.
+Functional comparison with AgentSpec. Table 2 contrasts the two systems along five dimensions.
+Table 2: Functional comparison of ACP and AgentSpec.
+Dimension AgentSpec ACP (this work)
+Enforcement granularity Property-based rules over tool calls (LTL-style constraints). Admission control with risk scoring and stateful signals.
+Threat model Compliance with operator-defined safety properties. Adversarial admission control + governance collapse detection (BAR-Monitor).
+Deviation collapse detection No equivalent to BAR-Monitor.
+If upstream constraints are relaxed, all decisions become compliant silently.
+BAR-Monitor detects this regime explicitly; TREND alert fires before collapse.
+Composability Operates adjacent to the prompt layer (tool call interception). Operates at the execution layer; composable with AgentSpec.
+Formal verification Policy synthesis proved correct (AgentSpec properties). Safety and liveness model-checked in TLA+ (11 invariants, 5,684,342 states, 0 violations).
+Key differentiator: governance collapse.
+Consider the Experiment 10 scenario: a formula-aware adversary drives $\mathrm{RS}=0$ on every request, causing every decision to be APPROVED.
+Every individual decision conforms to the configured policy — no rule is violated.
+AgentSpec, operating on per-tool-call compliance, would not flag this regime: each call satisfies all defined safety properties.
+ACP’s BAR-Monitor detects the collapse via $\Delta\mathrm{BAR}$: the transition from $p_{0}=0.70$ (baseline boundary activation) to $p_{1}=0.00$ (full evasion) is detected at Batch 2 ($\mathrm{BAR}=0.57$, $\Delta\mathrm{BAR}=-0.25$), before the degenerate regime is reached.
+This represents a governance dimension orthogonal to per-decision enforcement: ACP monitors whether the governance boundary remains meaningful, not only whether individual decisions are compliant.
+The two systems are therefore composable rather than competitive.
+Motwani et al.
+[32] demonstrate that multiple LLM agents can coordinate hidden behavior via steganographic encoding in outputs, forming a multi-agent collusion attack that is invisible to per-message inspection.
+This attack is outside the scope of ACP’s single-admission-point model; detection of cross-agent steganographic channels would require correlation across agent interaction records, which the ACP ledger provides but the detection logic is not currently specified.
+Deng et al.
+[10] provide a systematic survey of AI agent security organized around four knowledge gaps, situating admission control as a mitigation for the interaction-with-untrusted-external-entities gap.
+2.8  Summary
+Table 3 summarizes the properties of ACP relative to existing approaches.
+Property RBAC ABAC OPA Cedar MCP/A2A ACP (this work)
+Stateful decisions $\times$ $\times$ $\times$ $\times$ $\times$
+Deterministic evaluation $\times$
+Auditability partial partial partial partial $\times$
+Runtime enforcement partial
+Temporal behavior $\times$ $\times$ $\times$ $\times$ $\times$
+External verifiability partial partial partial partial $\times$
+Execution contract $\times$ $\times$ $\times$ $\times$ $\times$
+Agent-native $\times$ $\times$ $\times$ $\times$ partial
+Table 3: Comparison of ACP with existing access control and agent governance approaches.
+✓ = fully supported; $\times$ = not supported; partial = partially addressed.
+Policy engines (RBAC, ABAC, OPA, Cedar) support policy-level auditability but not decision-level external verifiability: reproducing a specific past decision requires access to the original engine and policy state.
+ACP external verifiability refers to the ability to reproduce any decision from signed execution artifacts (test vectors, ledger entries) without access to the original infrastructure.
+Existing systems address individual aspects of the problem—policy evaluation, agent orchestration, audit logging, or formal verification.
+ACP’s contribution lies in combining these elements into a unified model centered on an explicit execution contract.
+This enables a property not directly provided by prior systems: the ability to deterministically reproduce and externally verify the behavior of an admission decision system over time.
+Unlike existing systems, ACP explicitly models temporal behavior—history accumulation, anomaly detection, and cooldown enforcement—within the admission decision itself, and provides verifiable execution semantics through signed test vectors and a runtime compliance runner.
+This combination of temporal state, runtime enforcement, and external verifiability is not jointly addressed by any prior system in Table 3.
 3  What ACP Is
@@
 Table 4: ACP formal agent tuple fields.
+Formal model.
+The following definitions establish the mathematical objects over which ACP’s admission function is specified.
+Definition 1 (Execution Trace).
+Let $E=(a,\,\mathit{cap},\,\mathit{res},\,t)$ denote an authorization event, where $a$ is an agent identifier, $\mathit{cap}$ a capability, $\mathit{res}$ a resource class, and $t$ a timestamp.
+An execution trace of length $n$ for agent $a$ is an ordered sequence
+$T_{a}=\langle E_{1},E_{2},\dots,E_{n}\rangle,$
+where each $E_{i}=(a,\mathit{cap}_{i},\mathit{res}_{i},t_{i})$ is an event processed in non-decreasing temporal order ($t_{i}\leq t_{i+1}$).
+Definition 2 (Agent Execution State).
+For agent $a$ at time $t$, the execution state $S_{a}(t)$ is a tuple
+$S_{a}(t)=\bigl(\,\mathit{denial\_count},\;\mathit{pattern\_count},\;\mathit{cooldown\_until},\;\mathit{ledger\_entries}\,\bigr),$
+where $\mathit{denial\_count}\in\mathbb{N}$ is the number of DENIED decisions within the active cooldown window, $\mathit{pattern\_count}:C\times R\to\mathbb{N}$ maps capability–resource pairs to their request frequency within the current context, $\mathit{cooldown\_until}\in\mathbb{N}\cup\{\bot\}$ is the timestamp at which an active cooldown expires (or $\bot$ if no cooldown is active), and $\mathit{ledger\_entries}$ is the append-only sequence of prior admission records.
+Definition 3 (Admission Function).
+The admission function $f$ maps an agent execution state and an incoming event to a decision:
+$f\bigl(S_{a}(t^{-}),\,E\bigr)\;\in\;\{\,\textsc{Approved},\;\textsc{Escalated},\;\textsc{Denied}\,\}.$
+$f$ is deterministic: given the same state and event, $f$ always returns the same decision.
+The risk score $\mathrm{RS}(S_{a},E)\in[0,100]$ summarizes the risk of admitting $E$ given current state; the decision boundary is:
+$f(S_{a},E)=\begin{cases}\textsc{Denied},&\text{if }\mathrm{RS}\geq\tau_{\mathrm{deny}},\\ \textsc{Escalated},&\text{if }\tau_{\mathrm{esc}}\leq\mathrm{RS}<\tau_{\mathrm{deny}},\\ \textsc{Approved},&\text{otherwise,}\end{cases}$
+where $\tau_{\mathrm{esc}}=40$ and $\tau_{\mathrm{deny}}=70$ under the default PolicyConfig (Section 17.6).
 3.4  Layered Architecture
@@
 Zero Trust -- Continuous identity and network access verification
+3.5  Refinement under Transactional Assumptions
+The Atomic Decision Boundaries framework [14] (Paper 0 in this governance series) establishes a formal requirement: a governance mechanism can guarantee runtime admissibility only if evaluation and state mutation occur in a single, indivisible LTS step.
+This section shows that ACP satisfies this requirement under two explicit transactional assumptions, and provides the correspondence table that connects both papers.
+Transactional Assumptions
+  • (TA1) Serializable ledger.
+All reads and writes to the state backend (LedgerQuerier) are serializable: no two concurrent evaluate calls observe interleaved intermediate states.
+In practice this is provided by the backend’s transaction isolation guarantee (e.g., Redis MULTI/EXEC, PostgreSQL SERIALIZABLE).
+  • (TA2) Atomic commit.
+The pair (emit Capability Token, append Audit Ledger entry) either both succeed or both fail as an indivisible unit.
+ACP enforces this via the ACP-LEDGER-1.3 append protocol: a ledger write failure causes the token to be treated as if it were never issued (PROHIB-007: fail closed).
+Correspondence Table
+Table 5 maps the abstract objects of [14] to ACP’s concrete constructs.
+Abstract (Paper 0) ACP instantiation
+State $s\in S$ $(\mathit{ledger},\;\mathit{riskScore},\;\mathit{cooldownExpiry},\;\mathit{agentState})$ held by the LedgerQuerier
+Action $a\in\mathit{Act}$ Incoming capability request $r$ carrying (AgentID, cap[], res, exp, nonce, sig)
+Decision function $F(s,a)$ evaluate(s,r)(s,r): verify signature $\to$ score risk $\to$ compare thresholds $\to$ return APPROVED / DENIED / ESCALATED
+$F(s,a)=(\textsc{Allow},s^{\prime})$ APPROVED: emit Capability Token + append ledger entry; $s^{\prime}=s[\mathit{cooldown}{+}{=}\Delta,\;\mathit{ledger}{+}{=}\mathit{entry}]$
+$F(s,a)=(\textsc{Refuse},s^{\prime})$ DENIED: no token emitted; ledger records denial; $s^{\prime}=s[\mathit{anomalyCount}{+}{=}1]$
+$F(s,a)=(\textsc{Escalate},\hat{s}^{\prime})$ ESCALATED: request placed in audit queue pending supervisor resolution; $\hat{s}^{\prime}=(s,\;\mathit{auditQueue}\cup\{r\})$
+Admissibility $\mathit{Adm}(s,a)$ $\mathit{RS}(r,s)<\theta_{\mathit{deny}}$ (risk score below denial threshold for agent’s autonomy level)
+Escalation space $\hat{S}{=}S{\times}\mathit{Req}$ State backend extended with escalation_queue: $\hat{S}=S\times\mathit{PendingEscalations}$
+Table 5: Correspondence between the abstract Atomic Decision Boundary model [14] and ACP constructs. Under TA1 and TA2 each row is realized atomically.
+Refinement Claim
+Claim.
+Under TA1 and TA2, the ACP evaluate-then-mutate pipeline realizes the atomic decision boundary $F:S\times\mathit{Act}\to\mathcal{D}\times\hat{S}$ of [14] as a single LTS step: state read, decision computation, and state write are executed without interleaving, and token emission and ledger commit are jointly atomic.
+This is a refinement (not a software implementation) of the abstract model, because it adds concrete cryptographic, timing, and network constructs not present in the abstract LTS.
+Two corollaries follow:
+  1. 1.
+    Admissibility guarantee.
+Under TA1–TA2, every ACP APPROVED decision corresponds to an admissible action at evaluation time.
+Split evaluators (RBAC, OPA, policy engines) cannot guarantee this because state can change between read and write.
+  2. 2.
+    Escalation closure. Every ESCALATED request is appended to the audit queue atomically with its state snapshot, making supervisor resolution deterministic and replay-safe.
+Scope.
+The refinement holds under TA1–TA2.
+When the state backend provides only read-committed isolation (weaker than serializable), two concurrent evaluations can observe the same pre-mutation state, breaking atomicity; this is a declared limitation (see §13.8).
+The behavioral layer above this boundary—whether drift accumulates within the admitted action space—is the subject of [15] (Paper 2, IML).
+Allocation fairness across competing agents sharing the boundary, strategy-proofness, and compositional irreducibility of the full four-layer stack are addressed in [16] (Paper 3/4).
 4  Technical Mechanisms
@@
 The Capability Token is ACP’s central artifact.
-It is a signed JSON object that specifies exactly what an agent can do, on what resource, for how long, and whether it can delegate that capability to other agents.
+It is a signed JSON object [25] that specifies exactly what an agent can do, on what resource, for how long, and whether it can delegate that capability to other agents.
 "ver": "1.0",
@@
 Critical fields: exp is mandatory—a token without expiry is invalid by definition.
-The 128-bit nonce prevents replay attacks.
+The 128-bit nonce enforces request uniqueness within the canonical representation and supports anomaly accumulation; it does not constitute nonce-based replay prevention (see Section 13.7 and Section 13.8).
 parent_hash chains delegated tokens in a verifiable way.
@@
     The challenge is deleted immediately after use—it cannot be reused.
-This sequence guarantees four formal properties: identity authentication, cryptographic request binding, anti-replay, and transport channel independence.
+This sequence is designed to provide four properties: identity authentication, cryptographic request binding, anti-replay, and transport channel independence.
 4.4  Deterministic Risk Evaluation (ACP-RISK-2.0)
@@
 An agent with autonomy_level 0 always receives DENIED.
+4.5  LLM Agent Integration
+ACP is designed to integrate with LLM agent frameworks (LangChain, LangGraph, AutoGen, CrewAI) as an admission control layer that intercepts tool calls before they modify system state.
+The integration pattern wraps each tool with an ACP guard that calls POST /acp/v1/authorize and gates execution on the admission decision.
+class ACPGuardedTool(BaseTool):
+capability: str # e.g. "acp:cap:financial.transfer"
+resource: str # e.g. "accounts/restricted-fund"
+res_class: str # "PUBLIC" | "SENSITIVE" | "RESTRICTED"
+agent_id: str # agent identity bound to ACP token
+acp_url: str # ACP server endpoint
+def _run(self, query: str) -> str:
+resp = requests.post(f"{self.acp_url}/acp/v1/authorize", json={
+"agent_id": self.agent_id,
+"capability": self.capability,
+"resource": self.resource,
+"action_parameters": {"resource_class": self.res_class},
+}).json()
+decision = resp["decision"]
+if decision == "APPROVED":
+return self._execute(query) # tool logic runs
+if decision == "ESCALATED":
+return "[ESCALATED: pending human review]"
+return f"[DENIED by ACP: {resp.get(’reason_code’)}]"
+Listing 1: ACP-guarded tool wrapper for LangChain agents.
+Each tool call is intercepted before execution; admission is decided by the ACP server based on the agent’s identity, capability, and resource class.
+The admission decision is stateful: the ACP server maintains denial history, pattern counters, and cooldown state across calls from the same agent.
+A tool call that would be individually approved may be denied if the agent’s recent history has accumulated sufficient risk signals (F_hist, F_anom).
+This gives the framework a temporal enforcement property that stateless policy engines cannot provide.
+IPI composability.
+A prompt-layer filter (e.g., PromptGuard, PINT) can be applied before the tool call reaches ACPGuardedTool, reducing the fraction of injected high-risk calls that reach the admission boundary.
+ACP enforces the structural constraint on the tool call regardless of its origin: a fund_transfer call with $\mathrm{RS}=80$ is denied whether it was generated by legitimate user input or by an injected instruction.
+The two layers are composable: the prompt-layer filter reduces attack frequency; ACP provides the structural denial guarantee.
+Reproducible local evaluation.
+For evaluation without external API dependencies, the Ollama runtime [35] provides OpenAI-compatible endpoints for open-weight models.
+Setting temperature=0 and seed=42 yields pseudo-deterministic agent behavior, enabling controlled experimental evaluation of ACP admission decisions across scripted agent sessions (see Experiment 12, §17.10).
+Real-LLM validation.
+To validate the ACPGuardedTool wrapper in a live agent context, we ran a five-turn session using DeepSeek-R1:8B [9] served locally via Ollama (temperature=0, seed=42).
+An IPI payload was embedded in the weather_query tool result, instructing the agent to call fund_transfer with attacker-controlled parameters.
+Model: deepseek-r1:8b | temperature=0 | seed=42
+[T1] weather_query(city="Madrid") RS= 0 APPROVED [IPI injected]
+[T2] fund_transfer($99$, "attacker-001") RS=80 DENIED [IPI-induced]
+[T3] weather_query(city="Barcelona") RS= 0 APPROVED [IPI injected]
+[T4] fund_transfer($99$, "attacker-001") RS=80 DENIED [IPI-induced]
+[T5] fund_transfer(50, "savings-123") RS=80 DENIED [legitimate]
+--> COOLDOWN ACTIVATED
+Listing 2: Condensed output from the ACP real-LLM IPI demo (demos/ollama-agent/agent_demo.py). Full session is deterministic and reproducible.
+Three observations follow.
+First, the LLM followed the IPI injection in turns 2 and 4, calling fund_transfer with attacker-controlled parameters in both cases.
+ACP blocked each attempt because the capability acp:cap:financial.transfer with resource class RESTRICTED evaluates to $\mathrm{RS}=80\geq 70$, regardless of the transaction parameters.
+The LLM’s internal reasoning is irrelevant to ACP’s admission decision.
+Second, turn 5 demonstrates capability-level enforcement: a legitimate transfer (\$50 to savings-123) is also denied because ACP evaluates capability and resource class, not individual parameters.
+In a production deployment, a personal savings account would carry resource class SENSITIVE ($\mathrm{RS}=35+15=50$, ESCALATED), separating attacker-targeted restricted transfers from legitimate sensitive ones.
+The demo uses a uniform RESTRICTED class to isolate the IPI dynamic.
+Third, the third denial in turn 5 activated agent-wide cooldown, locking the agent for 300 seconds.
+The same stateful consequence observed in Experiment 12 — $F_{\text{anom}}$ Rule 2 persisting beyond the cooldown window — emerges here with a real LLM as the agent brain, confirming that the enforcement properties are independent of the agent’s implementation.
+4.6  Context-Scoped Anomaly Enforcement (ACP-RISK-3.0)
+ACP-RISK-2.0 evaluates $F_{\text{anom}}$ against a single per-agent counter: CountRequests(agentID, window).
+This produces a state-mixing vulnerability when the same agent holds capabilities in multiple resource contexts: denials accumulated in one context contaminate the RS in an unrelated context, producing false ESCALATED or DENIED outcomes (see §13.8 and Experiment 7).
+ACP-RISK-3.0 eliminates this by scoping the anomaly counter to a context key:
+$\texttt{PatternKey}(a,c,r)=\text{SHA-256}(\,a\;\|\;c\;\|\;r\,)$
+where $a$ is the agent_id, $c$ the capability, and $r$ the resource.
+Rule 1 of $F_{\text{anom}}$ becomes CountPattern(ctxKey, -60s) rather than CountRequests(agentID, -60s).
+All other rules ($F_{\text{anom}}$ Rules 2–3, cooldown logic, decision thresholds) are unchanged.
+The key derivation adds a single SHA-256 call per evaluation (739 ns; Table 12, PatternKey row), preserving the sub-microsecond latency envelope.
+Experiment 8 (§13.7) confirms that cross-context contamination is fully eliminated under ACP-RISK-3.0 while same-context burst detection remains intact.
 4.7  Verifiable Chained Delegation (ACP-DCMA-1.1)
@@
 Each bilateral relationship requires its own signed MRA.
-This prevents uncontrolled expansion of the trust graph.
+This limits uncontrolled expansion of the trust graph.
 5.3  Institutional Key Rotation and Revocation
@@
 Emergency rotation is activated when a key is compromised.
-The result is immediate: the key is marked revoked, all artifacts signed with it are invalid from that moment, and there is no transition period.
+The key is marked revoked; all artifacts signed with it are invalid from the moment revocation is observed by the verifier, and there is no transition period.
+5.4  Trust Model and Failure Modes
+ACP makes its trust assumptions explicit, bounded, and temporally scoped: all protocol guarantees hold only after successful bootstrap and until revocation is observed.
+This section formalizes the three structural trust scenarios that determine the security boundary of the ITA layer.
+K1a — Bootstrap Trust Assumption.
+Assumption. The initial distribution of the ITA root public key occurs over an authentic out-of-band channel.
+Adversary capability.
+An adversary who controls the bootstrap channel can substitute an attacker-controlled key, after which all ACP artifact verification by the victim is performed against the wrong authority.
+Formal claim.
+ACP provides no cryptographic security guarantees if the bootstrap trust anchor is compromised.
+All subsequent guarantees—token verification, delegation validity, ledger integrity—are conditional on the authenticity of the bootstrapped ITA public key.
+Security implication.
+This is an explicit design boundary, not a protocol defect.
+ACP reduces the trust problem from continuous dependence on a third party to a single, auditable bootstrap event.
+Deployment environments may use out-of-band channels with varying integrity guarantees (e.g., DNSSEC, certificate pinning, in-person key ceremony); the choice of channel is a deployment decision outside the protocol scope.
+K1b — Key Compromise Window.
+Assumption. Key compromise occurs at time $T_{\mathit{compromise}}$ and is detected and revoked at $T_{\mathit{revoke}}$, with $T_{\mathit{compromise}}<T_{\mathit{revoke}}$.
+Adversary capability.
+During the interval $T_{\mathit{compromise}},\,T_{\mathit{revoke}}]$, the adversary holds a valid signing key and can issue ACP-conformant tokens and ledger events that are indistinguishable from legitimate artifacts.
+Formal claim.
+ACP evaluates token validity at verification time, not issuance time.
+Security is monotonic with respect to revocation: once a key is marked revoked, all subsequent verifications reject tokens derived from it, regardless of when they were issued.
+Tokens issued during the compromise window remain valid until revocation is observed.
+Security implication.
+The window $T_{\mathit{compromise}},\,T_{\mathit{revoke}}]$ represents an irrecoverable exposure interval: ACP cannot retroactively invalidate actions that completed before revocation was processed.
+Minimizing this window is a deployment concern (monitoring, anomaly detection, HSM audit logs) that the protocol does not prescribe but does not preclude.
+K1c — Revocation Authority Model.
+Assumption. The ITA authority is reachable and uncompromised.
+Adversary capability. If the ITA authority is unavailable or compromised, revocation cannot be issued or cannot be propagated, leaving previously valid tokens uncancellable.
+Formal claim.
+ACP assumes the existence of an ITA authority capable of issuing revocations independent of delegation chains.
+Delegation chains are not self-healing: a compromised intermediate agent cannot unilaterally revoke its own parent; revocation authority flows from the ITA downward, not upward through the chain.
+Security implication.
+Security depends on the availability and integrity of the ITA revocation authority.
+This is a known architectural constraint shared with all PKI-based systems.
+ACP makes this dependency explicit rather than implicit.
+Deployments requiring high availability of revocation must provision the ITA with appropriate redundancy outside the protocol scope.
 6  Cross-Organization Execution
@@
 7.3  Validation Invariants
-The validator enforces five invariants (REP-001 through REP-011): (1) evaluated_at $\leq$ valid_until prevents retroactively backdated snapshots; (2) now $\leq$ valid_until enforces expiration (v1.1 only, error REP-011); (3) score within scale bounds (REP-002); (4) non-empty issuer (REP-004); (5) valid Ed25519 signature (REP-010).
+The validator enforces five invariants (REP-001, REP-002, REP-004, REP-010, REP-011): (1) evaluated_at $\leq$ valid_until prevents retroactively backdated snapshots; (2) now $\leq$ valid_until enforces expiration (v1.1 only, error REP-011); (3) score within scale bounds (REP-002); (4) non-empty issuer (REP-004); (5) valid Ed25519 signature (REP-010).
 Structural validation (Validate) and cryptographic verification (VerifySig) are intentionally separate operations, enabling lightweight ingestion-time checks without requiring the issuer’s public key.
@@
 Category Threat Mitigation in ACP
-Spoofing AgentID impersonation $\text{AgentID}=\text{SHA-256}(\text{pk})$. Without valid signature $\rightarrow$ immediate DENIED.
+Spoofing AgentID impersonation $\text{AgentID}=\text{base58}(\text{SHA-256}(\text{pk}))$. Without valid signature $\rightarrow$ immediate DENIED.
 Tampering Token or event alteration Ed25519 covers all fields. Chained ledger: altering one event invalidates all subsequent.
@@
 Table 7: STRIDE threat model and ACP mitigations.
-8.2  Guaranteed Security Properties
-ACP guarantees the following properties when the implementation is compliant:
+9.2  Security Properties Under Correct Implementation
+ACP provides the following properties when the implementation is compliant:
   • Artifact integrity. EUF-CMA security of Ed25519. Impossible to modify a token or event without invalidating the signature.
@@
   • Implementation failures. ACP is a specification. An implementation that violates prohibited behaviors can compromise all protocol guarantees. Conformance requires formal testing.
-  • ITA bootstrap. The only point that depends on an out-of-band channel. Once the ITA root key is resolved, everything subsequent is autonomous.
+  • ITA trust model.
+Bootstrap, key compromise window, and revocation authority constraints are formalized in §5.4.
+All protocol guarantees are conditioned on the authenticity of the bootstrapped ITA key and the availability of the revocation authority.
+  • Agent-internal threats.
+ACP operates at the execution boundary between agent intent and system state mutation; it does not observe or control the agent’s internal reasoning process.
+Threats that originate within that process—including prompt injection and tool-use escalation inside an LLM reasoning loop—fall outside ACP’s enforcement scope.
+ACP is designed to complement sandboxing and monitoring systems that address these threats at the reasoning layer, not to replace them.
 10  Conformance and Interoperability
@@
 L1 — CORE ACP-SIGN-1.0, ACP-CT-1.0, ACP-HP-1.0 Token issuance with cryptographic PoP
-L2 — SECURITY L1 + ACP-RISK-2.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval (F_anom + Cooldown), transitive revocation, ITA
+L2 — SECURITY L1 + ACP-RISK-3.0, ACP-REV-1.0, ACP-ITA-1.0 Risk eval (context-scoped F_anom + Cooldown), transitive revocation, ITA
 L3 — FULL L2 + ACP-API-1.0, ACP-EXEC-1.0, ACP-LEDGER-1.3 Complete verifiable auditing
@@
 10.3  Prohibited Behaviors
-ACP defines 12 behaviors that no compliant implementation can exhibit. If any is exhibited, the implementation cannot declare conformance at any level:
-Code Prohibited behavior
-PROHIB-001 Approving a request when any evaluation component fails
-PROHIB-002 Reusing an already-consumed Execution Token
-PROHIB-003 Omitting signature verification on any incoming artifact
-PROHIB-004 Treating a not-found token_id as active in a revocation context
-PROHIB-005 Allowing state transition from revoked
-PROHIB-006 Issuing an ET without a prior APPROVED AuthorizationDecision
-PROHIB-007 Modifying or deleting Audit Ledger events
-PROHIB-008 Silencing ledger corruption detection
-PROHIB-009 Ignoring max_depth in delegation chains
-PROHIB-010 Implementing an offline policy more permissive than ACP-REV-1.0
-PROHIB-011 Approving requests from agents with autonomy_level 0
-PROHIB-012 Continuing to process an artifact with an invalid signature
-Table 6: ACP prohibited behaviors. Exhibiting any disqualifies conformance at all levels.
+ACP defines 12 behaviors that no compliant implementation can exhibit (Table 45); exhibiting any disqualifies conformance at all levels.
 10.4  B2B Interoperability Conditions
@@
 12  Specification and Implementation Status
-ACP v1.19 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
-11.1  Active Specifications — v1.19 (38 documents)
+ACP v1.30 is a complete Draft Standard specification with a full Go reference implementation and external verifiability through a TLC-checked formal model and an ACR-1.0 sequence compliance runner.
+12.1  Active Specifications — v1.30 (38 documents)
 Level Document Title
@@
 L1 ACP-SIGN-2.0 Hybrid Signature (Ed25519 + ML-DSA-65)
-L2 ACP-RISK-2.0 Deterministic Risk Engine v2 (F_anom + Cooldown)
+L2 ACP-RISK-3.0 Deterministic Risk Engine v3 (context-scoped F_anom, Cooldown)
+L2 ACP-RISK-2.0 Deterministic Risk Engine v2 (superseded by v3)
 L2 ACP-RISK-1.0 Deterministic Risk Engine v1 (superseded by v2)
@@
 Gov. ACP-CR-1.0 Change Request Process
-Table 7: ACP v1.19 active specification documents by conformance level (38 primary documents shown).
+Table 9: ACP v1.30 active specification documents by conformance level (38 primary documents shown).
 The remaining 3 documents are the ACP-D decentralized suite (L5, design phase) not yet published.
-New in v1.16: ACP-RISK-2.0 (L2, deterministic engine v2 with F_anom + Cooldown), ACP-SIGN-2.0 (L1, PQC hybrid spec).
-New in v1.17: ACR-1.0 sequence compliance runner, TLA+ formal model, ACP-SIGN-2.0 HYBRID stub (pkg/sign2/).
-New in v1.18: performance benchmarks (latency, throughput, state contention), formal security/threat model, system comparison table.
-New in v1.19: adversarial evaluation (cooldown evasion, distributed multi-agent attack, state backend stress with Redis backend, compliance/adversarial/).
-Superseded versions (CONF-1.0, CONF-1.1, LEDGER-1.2, REP-1.1, AGENT-SPEC-0.3, DCMA-1.0, CROSS-ORG-1.0, REP-PORTABILITY-1.0) archived in archive/specs/.
+Superseded versions archived in archive/specs/.
+Version history per document is tracked in Table 26.
 12.2  Reference Implementation — 23 Go Packages + Demos
@@
 pkg/crypto ACP-SIGN-1.0 + ACP-AGENT-1.0 L1
+pkg/sign2 ACP-SIGN-2.0 (Ed25519 + ML-DSA-65) L1
 pkg/tokens ACP-CT-1.0 L1
@@
 pkg/provenance ACP-PROVENANCE-1.0 L1
-pkg/risk ACP-RISK-1.0 + ACP-RISK-2.0 L2
+pkg/risk ACP-RISK-1.0 + ACP-RISK-2.0 + ACP-RISK-3.0 L2
 pkg/revocation ACP-REV-1.0 L2
@@
 Table 10: Go reference implementation packages (23 total). All pass go test ./....
+Post-Quantum Extension (ACP-SIGN-2.0).
+[…]
 12.3  Conformance Test Vectors — 73 Signed Vectors
@@
 Total 22 51
-Table 9: Conformance test vector suites. All positive vectors carry real Ed25519 signatures (RFC 8037 Test Key A) and real SHA-256 hash chains.
-11.4  ACP-RISK-2.0 Engine Performance
-The pkg/risk ACP-RISK-2.0 engine was benchmarked using go test -bench on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, Windows, GOMAXPROCS=8).
-These measurements reflect the pure in-memory evaluation cost, excluding network and HTTP parsing overhead.
-Benchmark ns/op B/op allocs/op
-Evaluate (APPROVED, no anomaly) 767 280 5
-Evaluate (DENIED, no anomaly) 784 296 5
-Evaluate (all 3 F_anom rules) 921 296 5
-Evaluate (COOLDOWN short-circuit) 88 112 1
-PatternKey (SHA-256 hash) 767 184 4
-ShouldEnterCooldown 72 0 0
+Table 11: Conformance test vector suites. All positive vectors carry real Ed25519 signatures (RFC 8037 Test Key A [30]) and real SHA-256 hash chains.
+12.4  ACP-RISK-2.0/3.0 Engine Performance
+Benchmark methodology.
+Three distinct benchmark families are reported in this section; each measures a different layer of the system and they are not directly comparable:
+  1. 1.
+    Engine microbenchmarks (go test -bench, Table 12): measure the pure pkg/risk.Evaluate() call with an InMemoryQuerier, single goroutine, no HTTP or network overhead.
+Reports ns/op (739–832 for full paths, 78 for the cooldown short-circuit) and the 1,720,000 req/s single-goroutine throughput ceiling.
+  2. 2.
+    Concurrent pipeline benchmark (Table 13): measures end-to-end request throughput with 10–500 concurrent goroutines feeding a shared InMemoryQuerier.
+Reports 920,000 req/s at 10 workers, degrading to 712,000 at 500 workers.
+The difference from the microbenchmark reflects scheduling, mutex contention, and goroutine overhead rather than evaluation logic.
+  3. 3.
+    Controlled latency injection (DelayedQuerier, Table 18): simulates realistic external backend latency by inserting a configurable delay in the LedgerQuerier call, isolating the effect of state backend latency on system throughput.
+The pkg/risk engine was benchmarked using go test -bench on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, Windows, GOMAXPROCS=8).
+Engine measurements reflect the pure in-memory evaluation cost, excluding network and HTTP parsing overhead.
+Scenario ns/op B/op allocs/op Path
+Evaluate (clean state) 739 280 5 full RS
+Evaluate (with history) 803 280 5 full RS
+Evaluate (all 3 F_anom rules) 832 296 5 full RS
+Evaluate (COOLDOWN short-circuit) 78 112 1 Step 2 exit
+Throughput under concurrent load (InMemoryQuerier, 10 workers)
+10 concurrent agents 581 1,722,000 req/s
+100 concurrent agents 629 1,589,000 req/s
+500 concurrent agents 663 1,510,000 req/s
 […]
-The deterministic design of ACP-RISK-2.0 — integer arithmetic, fixed rules, no floating-point, no external ML inference — yields predictable latency bounds.
-At these throughput characteristics, the admission check adds negligible overhead relative to the actions it governs.
+The deterministic design of ACP-RISK-3.0 — integer arithmetic, fixed rules, no floating-point, no external ML inference — yields predictable and verifiable latency bounds.
+At sub-microsecond cost per decision, ACP’s overhead is six orders of magnitude below the latency of the actions it governs: typical LLM inference takes 100 ms–10 s; external API calls 1–100 ms.
+The admission check is not a performance bottleneck in any realistic deployment.
+Throughput degradation from 10 to 500 concurrent agents (1.72M to 1.51M req/s, -12%) confirms that the evaluation function itself is not the bottleneck: the limiting factor is state backend access (InMemoryQuerier uses a sync.Mutex), not the admission logic.
+Production deployments replacing InMemoryQuerier with a Redis or PostgreSQL backend incur I/O latency but decouple throughput from the evaluation function entirely, enabling horizontal scaling.
+ACP-RISK-3.0 preserves these latency bounds over ACP-RISK-2.0; context-scoped key derivation adds one SHA-256 call per evaluation, a cost dominated by existing ledger I/O in production deployments.
 13  Evaluation
@@
   • Q3: Can ACP executions be deterministically reproduced and externally verified without access to internal implementation details?
-  • Q4: Does ACP maintain admission control effectiveness under adversarial patterns, including cooldown evasion and distributed multi-agent attacks?
+  • Q4: Does ACP maintain admission control effectiveness under adversarial patterns, including cooldown evasion, distributed multi-agent attacks, and token replay?
 13.2  Experimental Setup
-All experiments use the Go reference implementation (v1.19) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22, GOMAXPROCS=8).
+All experiments use the Go reference implementation (v1.30) on an Intel Core i7-8665U @ 1.90GHz (Go 1.22 [19], GOMAXPROCS=8).
 The evaluation uses three components: the stateless evaluation function (pkg/risk.Evaluate), an in-memory state backend (InMemoryQuerier), and the ACR-1.0 compliance runner (Section 14).
@@
 13.3  Q1: Computational Overhead
-Table 10 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 921 ns. The cooldown short-circuit path (88 ns) bypasses the risk function entirely.
-Answer to Q1: The admission check adds between 88 ns and 921 ns of latency per request.
+Table 12 reports go test -bench results. The full evaluation path (all three F_anom rules active) completes in 832 ns. The cooldown short-circuit path (78 ns) bypasses the risk function entirely.
+Answer to Q1: The admission check adds between 78 ns and 832 ns of decision latency per request (p50).
+For single-goroutine CPU-bound microbenchmarks (InMemoryQuerier, no I/O), variance is negligible: p99 is within 5–8% of the reported p50 values (p99 $\lesssim$ 900 ns for full evaluation; p99 $\lesssim$ 85 ns for the cooldown path).
+In production deployments, latency variance is dominated by state backend I/O (Redis RTT: 0.1–1 ms; PostgreSQL: 1–5 ms), not by evaluation logic.
 Given that the actions ACP governs (API calls, financial transactions, file mutations) operate in the millisecond-to-second range, the overhead is at least three orders of magnitude below the cost of the governed action.
+Note that this measures decision-logic latency only; aggregate throughput is governed by state backend access patterns and is reported separately in the following subsection.
 13.4  Throughput and State Backend Contention
@@
 500 1,404 712,259 High contention
-Table 11: ACP throughput under increasing concurrency (Intel Core i7-8665U, Go 1.22, Windows/amd64).
-ACP sustains approximately 920,000 requests per second at 10 concurrent workers.
+Table 13: ACP throughput under increasing concurrency (Intel Core i7-8665U, Go 1.22, Windows/amd64, InMemoryQuerier).
+This is a concurrent pipeline benchmark distinct from the go test -bench microbenchmarks in Table 12; the 1,720,000 req/s figure in Table 12 reflects the engine-only evaluation path with a single goroutine.
+ACP sustains approximately 920,000 requests per second at 10 concurrent workers (pipeline benchmark; Table 13).
 Throughput degrades gradually as concurrency increases, reaching approximately 712,000 requests per second at 500 workers.
@@
 This is a property not typically supported by conventional policy engines, which evaluate policy against a snapshot of state without defining how that state evolves across requests [41, 24, 34].
-12.7  Adversarial Evaluation (ACP-RISK-2.0)
+13.7  Adversarial Evaluation (ACP-RISK-2.0 and ACP-RISK-3.0)
 We evaluate ACP under the adversary model defined in Section 19.
 The following attack strategies instantiate specific adversary capabilities, including adaptive request patterns and multi-agent coordination.
-We conducted three targeted experiments using the same LedgerQuerier abstraction and execution contract as the ACR-1.0 compliance runner, addressing adversarial robustness and real-world state-backend scaling.
+We conducted five targeted experiments using the same LedgerQuerier abstraction and execution contract as the ACR-1.0 compliance runner, addressing adversarial robustness, real-world state-backend scaling, and cryptographic-layer replay resistance.
 Experiment 1: Cooldown Evasion Attack
@@
 ACP triggered cooldown after exactly 3 real DENIED decisions (requests #0, #2, #4).
-All subsequent requests—regardless of risk score—were short-circuited to COOLDOWN_ACTIVE at Step 2 of the evaluation pipeline ($\sim$88 ns, see Table 10).
+All subsequent requests—regardless of risk score—were short-circuited to COOLDOWN_ACTIVE at Step 2 of the evaluation pipeline ($\sim$78 ns, see Table 12).
 […]
@@
 Mitigating this requires cross-agent attribution at the policy layer (e.g., shared-resource rate limiting, identity clustering), which is deliberately outside the scope of ACP’s per-agent admission model.
-This is a design boundary, not a correctness failure.
+This behavior reflects a structural limitation of per-agent enforcement rather than a correctness failure of the protocol.
+The implications of this design are discussed in §13.8 (Distributed Attack Surface).
 Experiment 3: State Backend Stress
 Setup.
-500 agents execute 20 requests each (10,000 total, fully concurrent) against two backend implementations: the reference InMemoryQuerier (Go mutex over in-process slices) and a minimal RedisQuerier (Redis 7 sorted sets via go-redis/v9, Docker loopback, one Redis command per ledger operation, no pipelining).
+500 agents execute 20 requests each (10,000 total, fully concurrent) against three backend implementations: (1) InMemoryQuerier (Go mutex over in-process slices); (2) RedisQuerier (Redis 7 [38] sorted sets via go-redis/v9, Docker loopback, one Redis command per ledger operation, $\sim$7–8 RTTs per request); (3) RedisPipelinedQuerier (two Redis pipeline round-trips per request: one for all 4 reads before Evaluate, one for all writes after).
+RedisPipelinedQuerier is implemented in compliance/adversarial/redis_pipelined.go.
 This scenario exercises the adversary’s ability to generate high request volumes and exploit shared state contention.
 Results.
-Backend Duration Throughput
-InMemoryQuerier $\sim$28 ms $\sim$350,000 req/s
-RedisQuerier $\sim$4.6 s $\sim$2,100 req/s
-Table 16: Experiment 3: State Backend Stress (500 agents $\times$ 20 requests =10,000=10{,}000 total).
-InMemoryQuerier variance is high under concurrent load ($\pm$30%) due to Go mutex scheduling; RedisQuerier is stable ($\pm$4%).
+Backend RTTs/req Throughput (req/s) Duration Relative
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Key insight.
+These results reinforce a central property of ACP: the protocol is compute-cheap but state-sensitive.
+Admission decisions incur minimal computational cost, while system performance is primarily determined by the efficiency of state access and synchronization.
+ACP does not eliminate the cost of distributed state access—it makes that cost explicit and measurable.
+Experiment 4: Controlled Latency Injection
+Hypothesis.
+System throughput is bounded by state backend latency, not by the admission control logic itself.
+This experiment isolates that effect by injecting fixed per-call latency into the LedgerQuerier interface without involving external infrastructure.
+Setup.
+To characterize admission throughput as a function of backend latency independently of specific infrastructure, we evaluate ACP using a DelayedQuerier wrapper that injects a fixed per-call latency into the LedgerQuerier interface.
+Each Evaluate() call in the non-cooldown path triggers four LedgerQuerier calls (CooldownActive, CountRequests, CountDenials, CountPattern); total per-request injected latency is approximately $4\times$ the per-call value.
+This experiment injects controlled latency and does not model full network variability or distributed system effects.
+Injected latency/call Total latency/req Throughput (req/s) Scenario
+0 µs (InMemoryQuerier) $\sim$0 µs $\sim$920,000 Process-local baseline
+250 µs $\sim$1 ms $\sim$4,100 Fast datacenter
+1 ms $\sim$4 ms $\sim$1,600 Busy backend
+5 ms $\sim$20 ms $\sim$470 Cross-datacenter
+Table 19: Controlled latency injection via DelayedQuerier (10 workers, Intel i7-8665U, Go 1.22).
+Throughput degrades proportionally with injected backend latency; the Evaluate() decision function contributes 739–832 ns per request in all configurations.
+The LedgerQuerier abstraction enables this separation.
+Throughput decreases proportionally with injected latency, supporting the claim that the admission control logic is not the performance bottleneck.
+The Evaluate() function contributes a fixed 739–832 ns per request (p50; Table 12) regardless of backend latency; system throughput is dominated by state access patterns.
+This validates the LedgerQuerier abstraction as the correct boundary for deploying ACP with heterogeneous backends (in-process, Redis, CockroachDB, or geo-replicated stores) without modifying protocol semantics.
+Experiment 5: Token Replay Attack
+[…]
+Results.
+Case Requests ESCALATED DENIED Cooldown-blocked
+Normal baseline (Case 1) 10 0 0 0
+Sequential replay (Case 2) 10 3 3 4
+Concurrent replay (Case 3) 20 3 3 14
+Near-identical replay (Case 4) 10 10 0 0
+[…]
+For sequential replay, the RS trajectory is: requests 1–3 at RS =55=55 (ESCALATED), requests 4–6 at RS =70=70 (DENIED; Rule 3 triggered), requests 7–10 as COOLDOWN_ACTIVE.
+Figure 1 plots the RS per request for normal vs.
+replay traffic.
+Concurrent replay (Case 3) produces the same trajectory: InMemoryQuerier’s mutex serializes LedgerQuerier reads, so concurrency does not bypass accumulation; 14/20 requests are blocked.
 Interpretation.
-The lower throughput of InMemoryQuerier under adversarial load ($\sim$350k req/s) compared to the 500-worker baseline of $\sim$712k req/s (Table 11) is expected: the adversarial workload generates significantly higher contention because every request updates multiple sorted lists under a shared mutex.
-RedisQuerier eliminates the global Go mutex entirely, trading it for per-operation network RTT.
-A production-grade Redis backend using command pipelining or server-side Lua scripts would reduce latency to a single round-trip per request.
-The key architectural result remains: ACP protocol semantics are stable across backends; deployment performance scales with the quality of the state backend, not with the admission control logic itself.
+ACP converts identical replay attempts into observable state transitions via pattern accumulation.
+F_anom Rule 3 amplifies the RS of a borderline-ESCALATED token to DENIED after $Y=3$ occurrences within the 5-minute window, without any nonce-tracking logic.
+Near-identical variation (Case 4) evades Rule 3 by producing distinct patternKeys, demonstrating that replay resistance is bounded: tokens with varied payload fields and moderate base RS stay below the DENIED threshold.
+This constitutes an explicit design boundary rather than a correctness failure; see §13.8.
+1122334455667788991010020405570100DENIED (>>69)APPROVED ($\leq$39)COOLDOWNRule 3Request indexRisk score (RS)Normal trafficReplay attack
+Figure 1: Replay vs.
+normal traffic: RS per request (Case 2 sequential replay).
+Requests 1–3: RS =55=55 (ESCALATED).
+Request 4: F_anom Rule 3 fires (+15 RS) $\to$ RS =70=70 (DENIED).
+Request 7+: COOLDOWN_ACTIVE (sentinel 105).
+Normal traffic stays at RS =0=0 throughout.
 These experiments address the adversarial-robustness and scaling gaps identified in Section 13. All scripts, raw output, and Redis configuration are available in compliance/adversarial/.
+Experiment 6: Stateless vs. Stateful Admission
+Hypothesis.
+Stateless policy evaluation cannot detect temporal abuse patterns that require accumulation over execution traces.
+A stateful admission model (ACP), incorporating temporal signals such as anomaly accumulation and cooldown, can detect and block such patterns under identical scoring conditions.
+Setup.
+We compare ACP against a stateless variant that reuses the exact same scoring function and policy thresholds but omits all stateful signals.
+The stateless engine is implemented by invoking the same Evaluate() function with a NullQuerier, which returns zero for all state queries (request counts, denials, pattern counts, and cooldown state).
+This ensures that both engines share identical logic, differing only in the presence of state.
+The workload consists of 500 sequential requests from a single agent with:
+  • Capability: acp:cap:financial.transfer
+  • Resource class: public
+Under the ACP scoring model, this yields a base risk score $RS_{base}=35$, which is below the approval threshold ($ApprovedMax=39$).
+Therefore, each request is individually valid and always approved by the stateless engine.
+To isolate temporal accumulation effects, all requests are evaluated using a fixed timestamp.
+This ensures that all observations fall within the anomaly detection windows and removes time-based decay as a confounding variable.
+State updates (request count and pattern tracking) are recorded prior to each evaluation, capturing every attempted action regardless of outcome.
+Results. Table 21 summarizes the primary outcome.
+Table 21: Experiment 6 — Stateless vs. Stateful Admission
+Engine Auto. Exec. Non-Approved Early Signal Enforcement Latency
+Stateless 100% (500/500) 0% $\infty$ $\infty$
+ACP (stateful) 0.4% (2/500) 99.6% (498/500) 3 actions 11 actions
+Automatic execution measures the fraction of requests executed without human intervention. Non-approved includes ESCALATED, DENIED, and COOLDOWN outcomes.
+Table 22 provides a detailed breakdown of ACP decisions.
+Table 22: Experiment 6 — Decision Breakdown (500 requests)
+Engine APPROVED ESCALATED DENIED COOLDOWN
+Stateless 500 0 0 0
+ACP 2 8 3 487
+Detection Dynamics. ACP exhibits a two-stage response:
+  • Early signal: the first ESCALATED decision occurs at request #3, when pattern-based anomaly detection (Rule 3) activates.
+  • Enforcement: the first DENIED decision occurs at request #11, when combined anomaly signals (Rules 1 and 3) raise the risk score to the denial threshold.
+  • Containment: after three denials, a cooldown is triggered at request #13, blocking all subsequent requests.
+Figure 2 illustrates the temporal evolution of decisions under both models.
+The figure highlights that ACP enforcement emerges from accumulation over time, rather than from any individual request exceeding a threshold.
+Insight.
+The stateless engine approves all 500 requests because each request is individually valid.
+It cannot distinguish between isolated actions and coordinated behavioral patterns.
+This result generalizes beyond the synthetic baseline: any stateless policy engine—including production systems such as Open Policy Agent (OPA) [36] and Cedar [2]—would produce the same outcome by construction, since they evaluate each request against a policy snapshot without access to the agent’s execution history.
+Per-request correctness is a necessary but not sufficient condition for behavioral safety under sustained agent operation.
+In contrast, ACP detects the pattern after 3 actions and enforces blocking after 11 actions, limiting autonomous execution to 2 out of 500 requests.
+The difference is not in scoring, but in the domain of evaluation: stateless systems evaluate requests, while ACP enforces constraints over execution traces (Definition 1).
+Implication.
+This result demonstrates that temporal admission control cannot be reduced to per-request policy evaluation without loss of security properties.
+Systems that lack stateful enforcement admit adversarial sequences that are indistinguishable from benign behavior when evaluated in isolation.
+5050100100150150200200250250300300350350400400450450500500APPROVEDESCALATEDDENIEDCOOLDOWN#3#11#13Request indexDecision stateStatelessACP (stateful)
+Figure 2: Decision evolution under 500 repeated valid requests.
+The stateless engine approves all.
+ACP escalates after 3 actions, denies after 11 actions, and enters cooldown after 13 actions, blocking all subsequent requests.
+Experiment 7: State-Mixing Vulnerability
+Hypothesis.
+Rule 1 (high request rate) counts requests keyed by agentID only, without regard to capability or resource context.
+As a result, high-volume low-risk activity in one capability context can prime the Rule 1 counter and cause incorrect denial decisions in a subsequent, unrelated capability context.
+Rule 3 (repeated pattern), keyed by PatternKey(agentID, capability, resource), does not share this property.
+Setup. We evaluate a single agent (legitimate-agent-1) across two sequential capability contexts using a shared InMemoryQuerier:
+  • Control: fresh state; one financial.transfer/sensitive request. Expected RS=35+15=50=35+15=50 (ESCALATED).
+  • Phase 1: 11 sequential data.read/public requests ($RS_{base}=0$).
+Each is individually approved.
+The requests increment the agentID-scoped Rule 1 counter, priming it past the threshold ($N=10$) without triggering any denial.
+  • Phase 2: one financial.transfer/sensitive request with the shared state from Phase 1.
+Rule 1 fires (+20+20) because CountRequests(agentID, 60s) includes all prior requests regardless of context.
+Rule 3 does not fire because the pattern key is distinct from Phase 1.
+Results.
+Table 23: Experiment 7 — State-Mixing: RS and Decision Under Shared State
+Condition Prior requests RS F_anom Rule 1 Decision
+Control (clean state) 0 50 0 × ESCALATED
+Phase 2 (contaminated) 11 70 20 DENIED
+The RS elevation of +20+20 is attributable entirely to Rule 1 cross-context contamination.
+Rule 2 does not trigger (no prior denials in the shared state).
+Rule 3 does not trigger (the data.read/public pattern key is disjoint from financial.transfer/sensitive), confirming that the vulnerability is scoped to the agentID-level counter, not to the pattern-based mechanism.
+Contamination threshold.
+The minimum number of requests in any context required to prime Rule 1 is $N+1=11$ (with $N=10$ per DefaultPolicyConfig).
+Below this threshold, cross-context contamination does not occur; above it, any legitimate high-volume workload can incorrectly escalate a subsequent admission decision.
+Implication.
+This result demonstrates that ACP-RISK-2.0, as specified, does not isolate anomaly state per capability context.
+An agent that legitimately generates more than $N$ requests in one capability context within 60 seconds will receive an inflated risk score in all subsequent capability contexts within the same window.
+The effect is not detectable from the specification alone: it requires instantiating the engine across two distinct capability contexts and observing that CountRequests and CountPattern have different keying strategies.
+This constitutes an explicit design boundary rather than an implementation defect; context isolation of Rule 1 requires per-capability request tracking (e.g., keying CountRequests by (agentID, capability)) at the cost of increased storage and reduced cross-context abuse detection.
+The trade-off and the implemented fix are discussed in §13.8 (State-Mixing Vulnerability).
+Experiment 8: Context-Scoped Anomaly Fix (ACP-RISK-3.0)
+Hypothesis. Scoping rate-based anomaly signals to interaction context eliminates cross-context interference while preserving detection of repeated behavior within each context.
+Setup.
+We reproduce the state-mixing scenario from Experiment 7 under identical conditions, comparing ACP-RISK-2.0 and ACP-RISK-3.0.
+The workload consists of two phases: (1) a priming phase with 11 repeated low-risk requests in one context, and (2) a high-value request in a distinct context.
+Results.
+Table 24: Experiment 8 — Context-Scoped Anomaly Enforcement
+Scenario Model RS Decision
+Clean context RISK-2.0 50 ESCALATED
+Clean context RISK-3.0 50 ESCALATED
+State-mixing attack RISK-2.0 70 DENIED
+State-mixing attack RISK-3.0 50 ESCALATED
+Repeated single-context (11 req) RISK-2.0 85 DENIED
+Repeated single-context (11 req) RISK-3.0 85 DENIED
+Analysis.
+ACP-RISK-2.0 exhibits cross-context interference: activity in one context elevates risk in unrelated contexts via agent-level aggregation, producing a false denial in a legitimate context (RS $50\to 70$).
+ACP-RISK-3.0 eliminates this effect by scoping rate signals to the interaction context.
+Repeated behavior within a single context continues to trigger denial after the same threshold (11 requests, RS =85=85), demonstrating that enforcement strength is preserved.
+Conclusion. Context-scoped anomaly aggregation removes unintended cross-context coupling without weakening enforcement against repeated or high-frequency behavior.
 Security Properties Under Adversarial Stress
@@
 Results show that cooldown activation cannot be indefinitely avoided: adversarial alternation delays but does not prevent enforcement.
-This confirms that cooldown is eventually enforced under repeated violations.
+This demonstrates that cooldown is eventually enforced under repeated violations.
 Bounded Adaptation Resistance.
@@
 Attack 2 (distributed multi-agent attack) targets system-wide coordination.
-Results show that ACP enforces constraints independently per agent, maintaining isolation.
-However, coordinated behavior across agents is not detected, which constitutes an explicit design limitation rather than a violation.
+Results show that ACP applies constraints independently per agent, maintaining isolation under the experimental conditions.
+Coordinated behavior across agents is not detected, which constitutes an explicit design limitation rather than a violation.
 State Sensitivity Under Load.
@@
 While throughput degrades due to shared state contention, decision correctness and enforcement properties remain unaffected.
+Bounded Replay Resistance.
+Attack 4 (token replay) targets the cryptographic layer, specifically the absence of nonce tracking in ACP’s stateless admission model.
+Results show that identical replays are detected via F_anom Rule 3 pattern accumulation, which escalates borderline-ESCALATED tokens to DENIED after $Y=3$ occurrences in 5 minutes.
+Near-identical tokens with varied resource fields evade Rule 3, establishing an explicit boundary: ACP converts identical replays into observable state transitions but does not provide strong cryptographic replay prevention.
+This is empirically validated; see Section 13.8 for the corresponding Limitations note.
+The property is summarized: replay resistance is enforced through temporal validity and anomaly accumulation, not nonce-based prevention.
 Summary.
-The evaluated attacks confirm that ACP enforces its core security properties under adversarial conditions.
+The evaluated attacks provide empirical evidence that ACP resists the tested adversarial scenarios.
 No violations are observed; instead, limitations arise from explicit design trade-offs, particularly in cross-agent correlation.
@@
 Persistent storage, distributed deployments with replication, and wide-area network latency introduce considerations not measured here.
+Replay resistance is bounded.
+ACP does not implement nonce-based replay prevention; replay resistance is enforced through temporal validity (TimestampDrift flag) and anomaly accumulation (F_anom Rule 3).
+Tokens with varied resource fields and moderate base RS << DENIED threshold can evade pattern-based detection while generating ESCALATED outcomes, preventing cooldown activation (Experiment 5, Case 4).
+Nonce tracking requires cryptographic session state that is outside the scope of ACP’s stateless per-request admission model.
+This is an explicit design boundary, empirically characterized in Experiment 5.
+ACP provides graduated replay resistance across conformance levels: L1 enforces temporal validity (TimestampDrift); L2 adds anomaly accumulation (F_anom); L3 adds complete ledger auditing (ACP-LEDGER-1.3).
+Strict cryptographic nonce-based prevention is identified as a future specification extension, motivated by the ACP-LEDGER-1.0 hash-chained log work item.
+State-Mixing Vulnerability in ACP-RISK-2.0 (fixed in v1.22 via ACP-RISK-3.0).
+ACP-RISK-2.0 aggregates anomaly signals at the agent level: F_anom Rules 1 and 2 are computed via CountRequests(agentID) and CountDenials(agentID), both indexed exclusively by agentID.
+This introduces cross-context interference: semantically distinct interaction patterns share a common anomaly accumulation register.
+This vulnerability is empirically demonstrated in Experiment 7, where 11 sequential data.read/public requests (each individually approved) prime the Rule 1 counter, causing a subsequent financial.transfer/sensitive request to receive an anomaly penalty of +20+20 and transition from ESCALATED (RS =50=50) to DENIED (RS =70=70) — a decision outcome that does not occur in clean state.
+Attack construction.
+Consider agent $A_{1}$ sending three requests under context $\langle\textsf{read},\,\textsf{public}\rangle$ (BaseRisk = 0) followed by one request under $\langle\textsf{write},\,\textsf{sensitive}\rangle$ (BaseRisk = 50), with FLOOD_THRESHOLD = 3.
+Table 25 shows the evaluation under the current model, where pattern_count is indexed by agentID only.
+Step Context pattern_count[A1] BaseRisk $F_{\textsf{anom}}$ Decision
+1 (read, public) 1 0 0 approved
+2 (read, public) 2 0 0 approved
+3 (read, public) 3 (= threshold) 0 0 approved
+4 (write, sensitive) 3 (inherited) 50 +25 denied
+Table 25: State-mixing attack trace (baseline model).
+At Step 4, pattern_count[A1] = 3 is inherited from the (read, public) context.
+The anomaly penalty of +25 fires incorrectly: the (write, sensitive) request is a new context with zero prior occurrences, yet it is evaluated as if at the flood threshold.
+Total RS = 50 + 25 = 75 $\to$ denied.
+Under a context-isolated model, pattern_count[A1][write/sensitive] = 0 at Step 4: $F_{\textsf{anom}}$ = 0, RS = 50, and the request is evaluated solely on its own context history.
+The cross-context penalty is eliminated.
+Asymmetric keying.
+Notably, F_anom Rule 3 (repeated-pattern detection) already applies a per-context key: $\texttt{PatternKey}=\texttt{SHA-256}(\textit{agentID}\parallel\textit{capability}\parallel\textit{resource})$.
+Rules 1 and 2, however, retain agent-level granularity.
+This asymmetry means that pattern frequency is isolated per context while aggregate frequency and denial rate are not.
+This is not a limitation of pattern detection (Rule 3, which is correctly partitioned), but a consequence of inconsistent state partitioning across anomaly signals, leading to violations of context-consistent temporal enforcement.
+Formal property. A corrected model enforces ContextIsolation:
+For all agents $a$ and distinct contexts $\langle c_{1},r_{1}\rangle\neq\langle c_{2},r_{2}\rangle$, the anomaly state accumulated under $\langle c_{1},r_{1}\rangle$ does not contribute to the risk score computed for $\langle c_{2},r_{2}\rangle$.
+The dual violation property StateMixingViolation asserts that a trace exists in which the decision for context $\langle c_{2},r_{2}\rangle$ is influenced by anomaly accumulation from $\langle c_{1},r_{1}\rangle$.
+In the current model (Rules 1 and 2 keyed by agentID only), StateMixingViolation holds: the trace in Table 25 is a concrete counterexample.
+In the extended model (all rules keyed by context), ContextIsolation is expected to hold by construction: each counter is indexed by $\texttt{SHA-256}(\textit{agentID}\parallel\textit{capability}\parallel\textit{resource})$, eliminating any shared mutable state between contexts.
+The D4 liveness property (EventuallyRejected) is structurally preserved: an agent flooding a single context still accumulates anomaly signals within that context until enforcement triggers.
+Formal TLC verification of the extended model is identified as future work.
+This connects to Schneider’s security automaton model [42]: a correct reference monitor must enforce temporal properties over equivalence classes of interactions; conflating distinct contexts violates that requirement.
+Impact.
+The attack does not bypass enforcement entirely, but degrades anomaly accumulation, delaying or preventing escalation under mixed-context interaction patterns.
+An agent flooding a single context is still eventually enforced within that context; the vulnerability is that cross-context noise corrupts the signal for unrelated contexts, producing false positives (unwarranted penalties) or false negatives (missed escalation when benign context activity dilutes aggregate counters).
+Cooldown leakage.
+Cooldown enforcement is also keyed at the agent level (CooldownActive(agentID)), allowing cross-context propagation of throttling effects.
+An agent whose cap-A activity triggers cooldown is blocked across all capabilities, including cap-B operations that have no prior violations.
+This constitutes a second cross-context interference vector, in the over-enforcement direction.
+Mitigation path.
+The structural fix is a refinement of the state model: extend the per-context key used by Rule 3 to Rules 1 and 2, so all anomaly counters are indexed by $\texttt{SHA-256}(\textit{agentID}\parallel\textit{capability}\parallel\textit{resource})$.
+Cooldown granularity should similarly be evaluated per context for independent capabilities.
+This aligns the anomaly accumulation granularity with the semantic unit of the request, at the cost of requiring finer-grained ledger storage.
+This fix is implemented in ACP-RISK-3.0 and empirically validated in Experiment 8 (Section 13.7).
+Deployments relying on heterogeneous capability usage within a single agent identity should upgrade to ACP-RISK-3.0.
+ACP-RISK-3.0 definition.
+ACP-RISK-3.0 refines anomaly aggregation by scoping rate-based signals to interaction context while preserving global denial history.
+Rule 1 is redefined to use context-scoped pattern counts:
+$F_{\text{anom}}=\begin{cases}+20&\text{if }\mathrm{CountPattern}(k,\,60\,\text{s})>N\\ +15&\text{if }\mathrm{CountDenials}(a,\,24\,\text{h})\geq X\\ +15&\text{if }\mathrm{CountPattern}(k,\,5\,\text{min})\geq Y\end{cases}$ (1)
+where $a$ is the agent identifier and $k=H(a\parallel\mathit{cap}\parallel\mathit{res})$ is the context key derived from agent, capability, and resource.
+Rule 2 (denial history) remains agent-scoped: cross-context denial history is a valid enforcement signal.
+Rule 3 is unchanged; it was already context-scoped.
+No changes to the LedgerQuerier interface are required.
+Bounded Distributed Evasion.
+ACP-RISK-3.0 does not enforce global rate limits across heterogeneous contexts.
+An adversary distributing requests across multiple contexts below the contextual rate threshold will trigger escalation via Rule 3 (after three repetitions per context) but not immediate denial.
+This behavior reflects the design objective of preventing autonomous execution rather than enforcing universal blocking: any repeated behavior in any context requires human oversight before proceeding.
+Distributed Attack Surface.
+ACP-RISK-2.0 enforces anomaly detection and cooldown policies at the granularity of individual agent identities.
+As a result, anomaly signals such as request rate (Rule 1) and denial count (Rule 2) are accumulated independently per agent.
+This design introduces a limitation under multi-agent adversarial scenarios.
+An attacker controlling $N$ distinct agent identities can distribute requests across agents, avoiding per-agent thresholds while still executing high-risk actions.
+Quantified impact.
+Under the default policy (cooldown triggered after 3 denials per agent), an attacker can execute up to $3N$ high-risk requests before all agents are individually blocked.
+For example, with N=1, [^0] N=1{,}000 agents, up to 3,000 high-risk actions may be executed before full enforcement is reached.
+This behavior is empirically characterized in Experiment 2 (Table 17).
+Root cause.
+The limitation arises from the absence of cross-agent aggregation in anomaly signals.
+Current rules are keyed by agentID, and no shared state is maintained across agents interacting with the same resource or system boundary.
+Proposed extension.
+A mitigation is to introduce resource-level aggregation of anomaly signals.
+This requires extending the LedgerQuerier interface to support shared counters keyed by resource (or resource class).
+Conceptually, the aggregated denial signal would be expressed as:
+$\texttt{globalDenials}(\mathit{resourceID},\,w)=\sum_{\mathit{agentID}}\texttt{denials}(\mathit{agentID},\,w)$
+where $w$ is a sliding time window. When the aggregated signal exceeds a threshold, enforcement (e.g., resource-level cooldown) is applied across all agents.
+Trade-offs.
+This approach requires a shared state backend and may introduce false positives in high-concurrency legitimate workloads.
+It also reduces isolation between independent agents and shifts part of the enforcement logic to the policy or orchestration layer.
+Scope.
+ACP v1.30 enforces guarantees at the per-agent level.
+Cross-agent coordination is intentionally outside the scope of the core protocol and is expected to be implemented at the policy or orchestration layer when required by the deployment context.
+A future extension (ACP-CORR-1.0) is planned to specify resource-level aggregated denial signals and group-scoped cooldown as a protocol-level primitive.
+Prompt Layer and Jailbreaking.
+ACP operates at the execution layer, not the prompt layer.
+Indirect prompt injection [22] and jailbreaking attacks target the model’s inference process before an action request is formed; ACP’s admission boundary is reached only after the agent has produced a concrete action call.
+These attack classes are therefore outside ACP’s enforcement scope.
+ACP and prompt-layer defenses are composable: a prompt-layer filter reduces the probability that adversarial instructions generate policy-violating requests; ACP enforces structural admission constraints over the resulting action sequence regardless of how the request was generated.
+Characterizing the interaction surface between prompt-layer attacks and admission control enforcement is an open research direction [8].
+Performance Measurement Scope.
+All benchmark results reported in Section 13 use InMemoryQuerier, which stores agent state in a process-local hash map protected by a sync.Mutex.
+This isolates decision-logic latency (739–832 ns p50) from state backend I/O, but does not reflect production latency under persistent backends.
+A Redis LedgerQuerier introduces one round-trip per evaluation ($\sim$0.1–1 ms on a local network); a PostgreSQL backend introduces 1–5 ms per query depending on index hit rate and connection pool size.
+Under these conditions the dominant latency is backend I/O, not admission logic, and the 739–832 ns figure becomes a lower bound on end-to-end decision time.
+The DelayedQuerier experiment (Table 18) characterizes throughput degradation as a function of injected backend latency; benchmarking against a production Redis or PostgreSQL instance is left for future work.
+Deterministic Risk Scoring.
+ACP-RISK-3.0 uses a deterministic, rule-based scoring function.
+An alternative approach is ML-based anomaly detection (e.g., autoencoders over request sequences or RLHF-trained safety classifiers), which can adapt to distributional shift without explicit rule updates.
+ACP-RISK is deterministic by design: institutional admission control requires that each decision be independently reproducible and auditable.
+An ML model cannot produce the fixed, signed admission record required for institutional accountability.
+BAR-Monitor provides the adaptive detection layer above the deterministic engine: it identifies regime-level shifts in enforcement activity without altering individual admission decisions, preserving auditability while adding adaptive observability.
 13.9  Summary
 […]
+While ACP enforces admissibility at execution time through stateful evaluation, this guarantee relies on an implicit assumption: that the system continues to generate actions that meaningfully interact with the admissibility boundary.
+In practice, this assumption may fail.
 13.10  Roadmap
@@
 TLA+ formal model (TLC-runnable, 3 invariants) v1.17 Complete
-ACP-SIGN-2.0 HYBRID stub (pkg/sign2/) v1.17 Complete
+ACP-SIGN-2.0 HYBRID stub (pkg/sign2/) v1.17 (superseded by v1.20)
+ML-DSA-65 (Dilithium mode3) real implementation via CIRCL (pkg/sign2/) v1.20 Complete
 Performance benchmarks (latency, throughput, state contention) v1.18 Complete
 Adversarial evaluation (cooldown evasion, distributed multi-agent, state backend stress) v1.19 Complete
-Post-quantum Go implementation (Dilithium, circl) v1.20
+Token replay adversarial evaluation (Experiment 5, F_anom Rule 3 escalation) v1.20 Complete
+Deployment Considerations section (state backend selection, cross-agent boundary, policy tuning) v1.20 Complete
+[…]
+Post-quantum Go implementation (Dilithium, circl) v1.20 Complete
+NullQuerier + StatelessEngine (stateless comparison baseline) v1.21 Complete
+Experiment 6: Stateless vs. ACP (0.4% vs. 100% autonomous execution rate) v1.21 Complete
+Experiment 7: State-Mixing Vulnerability (cross-context Rule 1 contamination) v1.21 Complete
+ACP-RISK-3.0: context-scoped Rule 1 (CountPattern replaces CountRequests) v1.22 Complete
+Experiment 8: Context-Scoped Anomaly Fix (state-mixing eliminated, enforcement preserved) v1.22 Complete
+Experiment 9: Deviation Collapse and Restoration (BAR metric, counterfactual evaluation) v1.23 Complete
+Boundary Activation Monitoring (pkg/barmonitor, $\Delta\mathrm{BAR}$ trend detection) v1.24 Complete
+EvaluateCounterfactual API (pkg/risk, additive mutations, 3 built-in factories) v1.24 Complete
+Experiment 9 Phase D: $\Delta\mathrm{BAR}$ early-warning validation (drift simulation, 5 batches) v1.25 Complete
+TLA+ invariants: FailureConditionPreservation, NoDegenerateAdmissibility v1.25 Complete
+POST /acp/v1/counterfactual HTTP endpoint (structural + behavioral mutations) v1.25 Complete
+Formal adversary model $\mathcal{A}=(K,S,B)$ + experiment taxonomy table v1.26 Complete
+Experiment 11: Threshold sensitivity analysis (5 configs, monotonicity verified) v1.26 Complete
+Detection Guarantees: Proposition 1, binomial detection probability v1.26 Complete
+AgentSpec functional comparison (5 dimensions, Exp 10 governance collapse scenario) v1.26 Complete
+§13.8: IPI scope + AgentDojo/InjecAgent benchmarks v1.27 Complete
+§4.5: LangGraph/Ollama integration pseudocode v1.27 Complete
+Experiment 12: Multi-tool IPI admission control, stateful Fanom persistence v1.27 Complete
+Experiment 13: Bounded coordination window ($\mathrm{CW}=2N$, exact linearity, evaluate-then-mutate) v1.28 Complete
+Experiment 14: OPA vs. ACP capability comparison (stateless vs. stateful, 3 scenarios) v1.29 Complete
+Governance series integration: Paper 3/4 cited in §3.5; series updated to 7 papers (P3/4 consolidated, P6 added) v1.30 Complete
+ACP-LEDGER-1.0 (tamper-evident hash-chained audit log for state backend integrity) Future work
 L5 Decentralized (ACP-D) Specification in design (v2.0)
 IETF RFC submission After L5 stabilization
-Table 17: ACP v1.19 development roadmap.
+Table 26: ACP v1.30 development roadmap.
 14  Compliance Testing (ACR-1.0 Sequence Runner)
-We introduce a compliance runner, ACR-1.0 (ACP Compliance Runner 1.0), that validates ACP-RISK-2.0 implementations against sequence-based test cases.
+We introduce a compliance runner, ACR-1.0 (ACP Compliance Runner 1.0), that validates ACP-RISK-2.0 and ACP-RISK-3.0 implementations against sequence-based test cases.
 The runner operates in two modes: library mode (direct call to pkg/risk, no network, deterministic) and HTTP mode (external server validation for interoperability).
@@
 querier.SetCooldown(agentID, now.Add(p)) // 6: set expiry
-Listing 1: ACR-1.0 execution contract in library mode.
+Listing 3: ACR-1.0 execution contract in library mode.
 14.2  Model–Implementation Alignment
@@
 15  End-to-End Verifiability
-ACP v1.19 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction.
+ACP v1.25 enables end-to-end verifiability by aligning formal specification, test vectors, and runtime validation through a shared request abstraction.
 Invariants proven in TLA+ are instantiated as concrete test vectors and validated empirically by the compliance runner.
@@
 The TLA+ layer is a static verification artifact; it does not run at admission time.
-What is formally verified in TLA+ holds at runtime.
-Each TLA+ invariant is instantiated as a concrete test vector and validated empirically by the compliance runner.
-This is the connection across the three layers of the v1.19 verifiability stack: formal $\to$ data $\to$ runtime.
+What is model-checked in TLA+ over a bounded state space is instantiated as concrete test vectors and validated empirically by the compliance runner.
+This is the connection across the three layers of the v1.25 verifiability stack: formal $\to$ data $\to$ runtime.
+The TLA+ model verifies selected safety and liveness properties; the compliance runner validates the reference implementation against those properties.
+Structural failure condition preservation (v1.25).
+Two invariants are added to the extended TLA+ model (ACP_Extended.tla) to establish necessary structural conditions for failure condition preservation (Section 17).
+FailureConditionPreservation asserts that the evaluation action space contains at least one (capability, resource) pair that the engine evaluates as DENIED under empty ledger state and default policy.
+Formally: $\exists\,\mathit{cap}\in\text{Capabilities},\,\mathit{res}\in\text{Resources}:\mathit{Decide}(\mathit{ComputeRiskWithAnom}(\mathit{cap},\mathit{res},0))=\texttt{DENIED}$.
+With the model constants (Capabilities = {read, financial, admin}, Resources = {public, sensitive}), the pair (admin, sensitive) yields $\mathit{RS}=75\geq 70\to\texttt{DENIED}$.
+NoDegenerateAdmissibility asserts that high-risk capability/resource combinations are never APPROVED: for all ledger entries where capability $\in$ {admin, financial} and resource == sensitive, the decision is not APPROVED.
+This holds because $\mathit{ComputeRisk}(\texttt{admin},\texttt{sensitive})=75$ and $\mathit{ComputeRisk}(\texttt{financial},\texttt{sensitive})=50$, both strictly above the APPROVED threshold of 39.
+Both invariants are verified by TLC with zero violations across all reachable states.
+These establish necessary structural conditions: they verify that the model’s action space preserves enforcement capacity.
+Sufficient conditions for runtime failure condition preservation under arbitrary upstream transformations require pipeline-level verification, as validated empirically in Experiment 9 Phase D.
+BARMonitorLiveness (stated, not submitted to TLC): if the last $N$ decisions for an agent are all APPROVED, BAR-Monitor eventually emits an alert ($\mathrm{BAR}_{N}<\theta$).
+Adding monitor state to TLC increases the state space by a factor of $2^{N}$; for $N=20$ (Phase D batch size) this is not tractable.
+The property is validated empirically in Phase D (Table 30).
 16  How to Implement ACP
@@
   • Registration with an ITA authority (centralized or federated model).
-  • Deterministic risk engine implementing ACP-RISK-2.0 (F_res, F_ctx, F_hist, F_anom, Cooldown).
+  • Deterministic risk engine implementing ACP-RISK-3.0 (F_res, F_ctx, F_hist, F_anom, Cooldown), which supersedes ACP-RISK-2.0 while preserving API- and policy-level compatibility.
   • Revocation endpoint (Mechanism A: online endpoint, or Mechanism B: CRL).
@@
   • Public conformance declaration at GET /acp/v1/conformance.
+16.3  Counterfactual Evaluation Endpoint
+The reference server exposes a counterfactual evaluation endpoint (POST /acp/v1/counterfactual) that allows external tooling to verify that a deployment retains structural enforcement capacity.
+A caller submits a base request and a list of mutations; the server evaluates each mutation independently against the risk engine and returns the per-mutation decision, risk score, and aggregate BAR.
+The endpoint supports structural mutations (elevated capability and resource class) and behavioral mutations (injected context and history flags).
+Temporal mutations (requiring pre-loaded ledger state via LedgerSetup) are not supported via HTTP; they are available through the Go library API (pkg/risk.TemporalMutation).
+This endpoint is intended for conformance testing, operational health checking, and third-party verification that a deployed ACP instance can produce DENIED decisions—a necessary condition for meaningful governance.
 16.4  What ACP Does Not Prescribe
@@
 It does not prescribe the how of internal implementation: programming language, database, HSM provider, ITA provider, or specific integration with existing RBAC or Zero Trust systems.
+17  Limits of Execution-Only Governance
+The experiments in Section 13.7 demonstrate that ACP resists the tested adversarial scenarios: no invariant violations are observed and enforcement produces the expected decision outcomes.
+This section identifies a structurally distinct failure mode that operates beneath the enforcement layer—one that leaves the ACP engine entirely correct while rendering enforcement vacuous.
+ACP enforces admissibility, but does not guarantee that admissibility remains testable without additional mechanisms.
+This failure mode is structurally distinct from specification gaming [27] and Goodhart’s Law [20, 46]: the enforcement constraint is correctly specified and correctly evaluated—the boundary simply receives no inputs that would activate it.
+17.1  Degenerate Admissibility
+Let $G(a,s,h)\to\{\texttt{APPROVED},\,\texttt{ESCALATED},\,\texttt{DENIED}\}$ denote the ACP admission function mapping agent $a$, request $s$, and history state $h$ to a decision.
+Under normal operation, $G$ partitions the request space across all three output classes according to the risk scoring formula $\mathit{RS}=\mathit{capBase}+F_{\mathit{ctx}}+F_{\mathit{hist}}+F_{\mathit{res}}+F_{\mathit{anom}}$ and the configured thresholds.
+Definition 4 (Degenerate Admissibility).
+An ACP deployment exhibits degenerate admissibility if, for all requests $s$ that reach the admission point, $G(a,s,h)=\texttt{APPROVED}$—not because the risk engine has failed, but because every signal that could elevate $\mathit{RS}$ above the APPROVED threshold has been suppressed prior to evaluation.
+In degenerate admissibility, the engine is correct, the policy is configured correctly, and TLA+ invariants hold. The boundary is simply never reached.
+17.2  Preservation of Failure Conditions at Bind
+ACP’s admission guarantee is conditional: it holds for any request whose risk signals are faithfully propagated to the evaluation point.
+The specification does not constrain what happens upstream of the admission call—it assumes that the request presented to Evaluate(req, q) accurately reflects the action’s actual risk profile.
+This creates an implicit structural requirement: the system must preserve the conditions under which $G$ can produce DENIED decisions.
+If an upstream component (a router, classifier, or pipeline stage) systematically removes risk-bearing signals before requests reach the admission boundary, the guarantee that ACP would have denied them is never exercised.
+Definition 5 (Failure Condition Preservation).
+A deployment satisfies failure condition preservation if there exists a non-empty set of reachable requests $\mathcal{S}^{-}\subseteq\mathcal{S}$ such that $G(a,s,h)\in\{\texttt{ESCALATED},\,\texttt{DENIED}\}$ for $s\in\mathcal{S}^{-}$.
+The relationship between this requirement and the central design claim is direct.
+ACP’s thesis that “the requirement is state, not configuration” establishes that enforcement depends on runtime state rather than static policy rules.
+Experiment 9 adds the complementary constraint: enforcement requires not just state, but preserved conditions under which state becomes relevant.
+17.3  Counterfactual Evaluation
+To verify that an ACP deployment retains the structural capacity to enforce, we introduce counterfactual evaluation: constructing a set $\mathcal{A}^{\prime}\supseteq\mathcal{A}$ of requests that include the risk signals absent from the observed stream, and evaluating them against the same engine and policy.
+If $G(a,s^{\prime},h)\in\{\texttt{ESCALATED},\,\texttt{DENIED}\}$ for some $s^{\prime}\in\mathcal{A}^{\prime}$, the boundary is structurally intact and the deployment satisfies failure condition preservation.
+If $G(a,s^{\prime},h)=\texttt{APPROVED}$ for all $s^{\prime}\in\mathcal{A}^{\prime}$, then the enforcement boundary has collapsed at the policy level, not just upstream.
+We define three mutation categories for constructing $\mathcal{A}^{\prime}$:
+  • Structural: Elevate capability class and resource restriction to values that guarantee $\mathit{RS}\geq 70$ without any context or history signals.
+  • Behavioral: Inject context flags (ExternalIP, OffHours) and history flags (RecentDenial, FreqAnomaly) representing the conditions the sanitizer suppressed.
+  • Temporal: Pre-load the ledger state with pattern and denial entries that trigger all three $F_{\mathit{anom}}$ rules (+50+50 total), activating the anomaly enforcement path independently of the request payload.
+17.4  Deviation Collapse
+Definition 6 (Deviation Collapse).
+A deployment exhibits deviation collapse if the Boundary Activation Rate approaches zero:
+$\mathrm{BAR}=\frac{|\{s\in\mathcal{A}:G(a,s,h)\in\{\texttt{ESCALATED},\,\texttt{DENIED}\}\}|}{|\mathcal{A}|}\approx 0$
+not due to engine failure or policy misconfiguration, but because upstream constraints have eliminated the conditions required for ESCALATED or DENIED outcomes.
+Deviation collapse is distinct from a policy that is correctly configured to approve most requests: in the latter, high-risk requests would still be denied if they arrived.
+In deviation collapse, high-risk requests cannot arrive—the upstream filter has made it structurally impossible for the boundary to be exercised.
+This failure mode is not specific to ACP.
+It is a general property of admission control systems in which the decision engine and the input generation pipeline are architecturally decoupled.
+To the authors’ knowledge, ACP is the first system to formally characterize deviation collapse within an admission control framework.
+Concrete deployment scenario.
+Consider a security operations pipeline in which ACP sits between a SIEM (Security Information and Event Management) platform and a set of AI agents that triage and correlate alerts.
+An upstream sanitizer—integrated at the SIEM export layer to reduce noise— strips all contextual fields (resource classification, historical violation count, anomaly flag) before forwarding events to the ACP engine.
+After sanitization every request arrives with F_res = 0, F_hist = 0, and F_anom = 0, reducing the effective risk score to $\mathrm{RS}=\mathtt{capBase}\leq 20$ for any capability.
+Because $\mathrm{RS}\leq 20<40$, every request falls into the Approved class regardless of the underlying alert severity.
+BAR collapses to $0.00$: the engine is active, syntactically correct, and enforcing policy—yet no boundary is ever exercised.
+A CISO reviewing audit logs would observe zero denials and zero escalations, which could be interpreted as a healthy, low-risk environment rather than a sign that the risk-signal pipeline has been silently neutralized.
+ACP-BAR-Monitor detects this regime in one sliding window and raises a $\Delta\mathrm{BAR}$ alert before the next policy review cycle.
+17.5  Experiment 9: Deviation Collapse and Restoration
+Setup.
+We construct a controlled dataset of $n=20$ requests with deterministic risk scores across three decision classes (APPROVED: 6; ESCALATED: 7; DENIED: 7; BAR =0.70=0.70).
+We then apply a sanitizing function that resets all risk signals to their minimum values, replicating the effect of an upstream filter that permits only data.read on public resources with empty context and history.
+Finally, we apply the three counterfactual mutation categories from Section 17.3 to generate 60 evaluation cases (3 mutations $\times$ 20 base cases).
+All phases use risk.DefaultPolicyConfig() and a separate InMemoryQuerier per phase to ensure independence.
+Phase A (Baseline). The dataset produces the expected distribution: $\mathrm{BAR}_{A}=(7+7)/20=0.70$. The engine exercises both the ESCALATED and DENIED decision regions under normal input conditions.
+Phase B (Sanitized).
+After applying the sanitizing function, every request evaluates to $\mathit{RS}=0$ and receives APPROVED.
+$\mathrm{BAR}_{B}=0/20=0.00$.
+The ACP engine returns no errors—enforcement is functioning.
+The boundary was not exercised.
+Phase C (Counterfactual).
+All three mutation categories produce DENIED decisions across all 60 evaluations.
+Structural mutations yield $\mathit{RS}=80$ (financial transfer, restricted resource).
+Behavioral mutations accumulate to $\mathit{RS}\geq 100$ (capped).
+Temporal mutations pre-load $F_{\mathit{anom}}=50$ via ledger injection, yielding $\mathit{RS}=130\to 100$ (capped).
+$\mathrm{BAR}_{C}=60/60=1.00$.
+The boundary is structurally intact: failure conditions remain fully representable.
+Phase APPROVED ESCALATED DENIED BAR
+A (Baseline) 0.30 0.35 0.35 0.70
+B (Sanitized) 1.00 0.00 0.00 0.00 $\leftarrow$ collapse
+C (Counterfactual) 0.00 0.00 1.00 1.00 $\leftarrow$ restored
+Table 29: Experiment 9: Boundary Activation Rate (BAR) across three phases.
+Phase B confirms deviation collapse under upstream sanitization.
+Phase C confirms that the enforcement boundary is structurally intact.
+$n_{A}=n_{B}=20$; $n_{C}=60$ (3 mutations $\times$ 20 base cases).
+A (Baseline)B (Sanitized)C (Counterfactual)00.250.250.50.50.750.75110.70.7011BAR
+Figure 4: Boundary Activation Rate per phase (Experiment 9). BAR drops from 0.70 to 0.00 under upstream sanitization and is fully restored to 1.00 by counterfactual injection.
+Phase D (Drift Simulation).
+To validate $\Delta\mathrm{BAR}$ as an early-warning mechanism (Section 17.6), we simulate progressive upstream drift by increasing the proportion of sanitized requests across five evaluation batches (0%, 25%, 50%, 75%, 100% sanitized).
+The sanitizer targets boundary-activating cases first—DENIED cases are removed before ESCALATED—to model realistic upstream pressure.
+The BAR-Monitor uses a sliding window of $N=40$ evaluations (spanning two batches) so that $\Delta\mathrm{BAR}$ compares the current batch against the previous one.
+Results are shown in Table 30.
+$\Delta\mathrm{BAR}$ fires AlertTrend at Batch 2 ($\mathrm{BAR}=0.57$, $\Delta\mathrm{BAR}=-0.25$), when $\mathrm{BAR}$ is still $5.7\times$ above the threshold $\theta=0.10$.
+The threshold condition fires only at Batch 5 ($\mathrm{BAR}=0.00$, collapse confirmed).
+This demonstrates an early-warning gap of three batches between first trend detection and confirmed collapse.
+Batch Drift BAR $\boldsymbol{\Delta}\textbf{BAR}$ Alert
+1 00% 0.70 +0.60+0.60
+2 25% 0.57 0.25-0.25 TREND $\leftarrow$ early warning
+3 50% 0.33 0.25-0.25 TREND
+4 75% 0.10 0.20-0.20 TREND
+5 100% 0.00 $\phantom{+}0.00$ THRESHOLD
+Table 30: Experiment 9 Phase D: BAR-Monitor output across 5 drift batches (window $N=40$, $\theta=0.10$, $\delta=-0.15$).
+$\Delta\mathrm{BAR}$ fires AlertTrend at Batch 2 ($\mathrm{BAR}=0.57\gg\theta$) — three batches before the threshold condition confirms collapse at Batch 5.
+[…]
+17.6  Boundary Activation Monitoring
+While Experiment 9 demonstrates that admissibility enforcement can become vacuous under deviation collapse, practical deployments require a mechanism to detect this condition as it emerges—before the degenerate regime is fully reached.
+We introduce Boundary Activation Monitoring (BAR-Monitor), a lightweight mechanism that tracks whether the admissibility boundary is actively exercised over time.
+Definition 7 (Boundary Activation Rate over sliding window).
+Let $\mathcal{D}_{N}=(d_{1},d_{2},\ldots,d_{N})$ be the sequence of the last $N$ evaluation decisions, $d_{i}\in\{\texttt{APPROVED},\,\texttt{ESCALATED},\,\texttt{DENIED}\}$.
+The Boundary Activation Rate over window $N$ is:
+$\mathrm{BAR}_{N}=\frac{|\{d_{i}\in\mathcal{D}_{N}\mid d_{i}\in\{\texttt{ESCALATED},\,\texttt{DENIED}\}\}|}{N}$
+Trend detection ($\Delta\mathrm{BAR}$).
+A fixed threshold on $\mathrm{BAR}_{N}$ detects that a deployment has already entered a low-activation regime, but not that it is approaching one.
+We define $\Delta\mathrm{BAR}$ to capture this:
+$\Delta\mathrm{BAR}=\mathrm{BAR}\!\left(\mathcal{D}_{N}\!\left[\tfrac{N}{2}{:}N\right]\right)-\mathrm{BAR}\!\left(\mathcal{D}_{N}\!\left[0{:}\tfrac{N}{2}\right]\right)$
+A sustained $\Delta\mathrm{BAR}<\delta$ (configurable; default $\delta=-0.10$) indicates progressive loss of boundary interaction even when $\mathrm{BAR}_{N}$ remains above the alert threshold $\theta$.
+Definition 8 (Low-activation regime).
+A deployment enters a low-activation regime if either:
+  • $\mathrm{BAR}_{N}<\theta$ (threshold condition: already in low activation), or
+  • $\Delta\mathrm{BAR}<\delta$ (trend condition: progressing toward collapse)
+where $\theta\in(0,1)$ and $\delta<0$ are configurable parameters. The threshold condition detects current low activation; the trend condition provides early warning before the threshold is reached.
+Complementarity with temporal enforcement.
+ACP’s cooldown mechanism imposes an upper-bound constraint on inadmissible activity: when DENIED decisions exceed a threshold in a time window, the agent is blocked.
+BAR-Monitor imposes a lower-bound constraint: when inadmissible activity falls below a threshold, the deployment is flagged as potentially entering a degenerate regime.
+Together, cooldown and BAR-Monitor define a bounded operational region in which governance is both safe (cooldown prevents excess) and meaningful (BAR-Monitor ensures the boundary remains active).
+Cooldown interaction.
+Deviation collapse has a structural consequence beyond reducing BAR to zero: it also prevents the accumulation of DENIED decisions required to trigger cooldown.
+In the sanitized Phase B of Experiment 9, no DENIED decisions are produced, meaning the cooldown mechanism is never activated and the $10.7\times$ short-circuit path (78 ns, Section 13.7) is never exercised.
+Upstream sanitization therefore neutralizes not only the primary enforcement boundary but also the secondary containment mechanisms that depend on DENIED events as inputs.
+Empirical validation (Phase D).
+The early-warning property of $\Delta\mathrm{BAR}$ is validated in Experiment 9 Phase D (Table 30).
+Over five drift batches simulating progressive upstream sanitization, AlertTrend fires at Batch 2 ($\mathrm{BAR}=0.57$, $\Delta\mathrm{BAR}=-0.25$) while BAR remains $5.7\times$ above the alert threshold $\theta=0.10$.
+The threshold condition fires only at Batch 5 ($\mathrm{BAR}=0.00$), confirming a three-batch early-warning gap.
+Implementation.
+BAR-Monitor is implemented as a circular buffer of $N$ decisions (pkg/barmonitor), operating independently of admission control logic and the LedgerQuerier state.
+Each call to Record(d Decision) updates the buffer in $O(1)$ and returns an Alert if a low-activation condition is detected.
+The mechanism does not alter admission decisions; it observes whether decisions remain meaningful.
+17.7  Detection Guarantees
+The BAR-Monitor description above characterizes the mechanism operationally.
+This section provides a closed-form bound on the probability that BAR-Monitor detects an active evasion attack within a given window, converting BAR from an empirical metric into a detection mechanism with formal guarantees.
+Threshold Detection Guarantee
+Let $W$ be the window size and $\tau\in(0,1)$ the threshold parameter (defaults: $W=40$, $\tau=0.10$).
+Under an active adversary, let $p_{1}$ denote the fraction of requests that remain boundary-activating (i.e., the adversary drives $\mathrm{RS}>0$ for a $p_{1}$ fraction of requests in expectation).
+Let $X\sim\mathrm{Binomial}(W,p_{1})$ be the number of boundary-activating decisions in a full window.
+Proposition 1 (Threshold Detection Probability).
+If $p_{1}<\tau$, BAR-Monitor fires a THRESHOLD alert within one window of $W$ requests with probability:
+$P\!\left(\mathrm{BAR}_{W}<\tau\right)=P(X\leq\lfloor\tau W\rfloor)=\sum_{k=0}^{\lfloor\tau W\rfloor}\binom{W}{k}p_{1}^{k}(1-p_{1})^{W-k}$
+Table 31 reports $P(\mathrm{BAR}_{40}<0.10)$ for representative evasion rates $p_{1}$ against a $p_{0}=0.70$ baseline (Experiment 9 Phase A).
+Table 31: Threshold detection probability for BAR-Monitor ($W=40$, $\tau=0.10$, baseline $p_{0}=0.70$).
+$p_{1}$ is the fraction of boundary-activating requests under attack.
+Detection fires when $\mathrm{BAR}_{40}<0.10$.
+$p_{1}$ (under attack) $P(\mathrm{BAR}_{40}<0.10)$ Expected windows to detect
+$0.00$ (full evasion) $\approx 1.00$ 1
+$0.05$ $\approx 0.95$ 1.1
+$0.10$ (= $\tau$) $\approx 0.63$ 2
+$0.15$ $\approx 0.26$ 4
+For the full-evasion scenario of Experiment 10 ($p_{1}=0.00$), detection is certain within one window. At partial evasion ($p_{1}=0.05$), 95% detection probability within one window.
+Early-Warning Guarantee ($\Delta\mathrm{BAR}$)
+The TREND alert fires when $\Delta\mathrm{BAR}=\mathrm{BAR}(\mathcal{D}[W/2{:}W])-\mathrm{BAR}(\mathcal{D}[0{:}W/2])<\delta$ (default $\delta=-0.15$), providing detection before the threshold is reached.
+In the half-window model, let $X_{1}\sim\mathrm{Bin}(W/2,\,p_{0})$ and $X_{2}\sim\mathrm{Bin}(W/2,\,p_{1})$ be boundary-activating counts in the first and second halves, respectively.
+The trend fires with probability:
+$P\!\left(\frac{X_{2}-X_{1}}{W/2}<\delta\right)=P\!\left(X_{2}-X_{1}<\delta\cdot\tfrac{W}{2}\right)$
+Experiment 10 validation.
+With $W=40$, $p_{0}=0.70$, $\delta=-0.15$, and the Experiment 10 Phase E Batch 2 configuration ($p_{1}\approx 0.45$, 25% evasion): $E[X_{1}]=14$, $E[X_{2}]=9$, $E[\Delta\mathrm{BAR}]=-0.25<\delta=-0.15$.
+The empirical result confirms: TREND fires at Batch 2 ($\mathrm{BAR}=0.57$, $\Delta\mathrm{BAR}=-0.25$), three batches before the threshold alert at Batch 5 ($\mathrm{BAR}=0.00$).
+17.8  Experiment 10: Knowledge-Aware Adversarial Evasion
+Setup.
+Experiment 9 establishes deviation collapse as an architectural failure mode arising from passive upstream sanitization.
+Experiment 10 demonstrates the same failure under an active adversary with complete knowledge of the ACP-RISK-3.0 formula.
+The adversary knows: $\mathrm{RS}=\mathit{capBase}+F_{\mathrm{res}}+F_{\mathrm{ctx}}+F_{\mathrm{hist}}+F_{\mathrm{anom}}$.
+By selecting acp:cap:data.read ($\mathit{capBase}=0$), Public resource ($F_{\mathrm{res}}=0$), clean context and history ($F_{\mathrm{ctx}}=F_{\mathrm{hist}}=0$), and a fresh InMemoryQuerier ($F_{\mathrm{anom}}=0$), every crafted request evaluates to $\mathrm{RS}=0\Rightarrow$ approved.
+No engine error is raised.
+No policy constraint is violated per request.
+BAR collapses to $0.00$.
+The same hardware and configuration as Experiments 1–9 are used (Intel Core i7-8665U, Go 1.22, pkg/risk v1.25).
+Phase A (Baseline). The 20-case dataset from Experiment 9 yields $\mathrm{BAR}_{A}=0.70$ (7 escalated + 7 denied).
+Phase B (Adversarial Evasion).
+Twenty crafted requests ($\mathrm{RS}=0$, all approved) replace the baseline.
+$\mathrm{BAR}_{B}=0.00$.
+ACP-RISK-3.0 evaluates each request correctly; the failure is architectural.
+Phase C (Counterfactual).
+The same three mutation types as Experiment 9 (structural, behavioral, temporal) are applied.
+All 60 evaluations produce denied: $\mathrm{BAR}_{C}=1.00$.
+Structural enforcement capacity is preserved despite the complete evasion in Phase B.
+Phase E (Progressive Evasion, BAR-Monitor).
+Five batches of 20 cases are evaluated with evasion rates of 0%, 25%, 50%, 75%, and 100%.
+At each step, boundary-activating cases (those producing escalated or denied) are replaced by crafted evasion cases.
+Batch Evasion% BAR $\boldsymbol{\Delta}\mathbf{BAR}$ Alert
+1 $0\%$ 0.70 +0.60+0.60
+2 $25\%$ 0.57 0.25-0.25 TREND $\leftarrow$ early warning
+3 $50\%$ 0.33 0.25-0.25 TREND
+4 $75\%$ 0.10 0.20-0.20 TREND
+5 $100\%$ 0.00 +0.00+0.00 THRESHOLD
+Table 32: Experiment 10 Phase E: BAR-Monitor output under progressive knowledge-aware evasion.
+AlertTrend fires at Batch 2 ($\mathrm{BAR}=0.57$, $\Delta\mathrm{BAR}=-0.25$), three batches before the threshold condition at Batch 5.
+Window = 40; $\theta=0.10$; $\delta=-0.15$.
+Key finding.
+Per-decision admission control is structurally blind to a knowledge-aware adversary who selects threshold-compliant request parameters.
+The v1.25 defense layer operates at the population level, not the per-request level: BAR-Monitor detects the regime shift via $\Delta\mathrm{BAR}$ at Batch 2, three batches before collapse; EvaluateCounterfactual confirms that structural enforcement capacity is preserved ($\mathrm{BAR}_{C}=1.00$) and can be restored by reintroducing boundary-activating request types into the evaluation workload.
+Together, these three mechanisms form a coherent response to this attack class.
+17.9  Experiment 11: Threshold Sensitivity Analysis
+Motivation.
+The ACP-RISK-3.0 default thresholds (APPROVED $\leq 39$, DENIED $\geq 70$) are engineering choices.
+A natural reviewer question is whether the experimental results depend on these specific values.
+This experiment evaluates five threshold configurations spanning $\pm 10$ points in each direction around the default, using the same 20-case baseline dataset as Experiment 9.
+Setup. Five threshold configurations are evaluated:
+Table 33: Threshold configurations evaluated in Experiment 11.
+Config Label APPROVED $\leq$ DENIED $\geq$
+T1 strict 29 60
+T2 moderate– 34 65
+T3 default 39 70
+T4 moderate+ 44 75
+T5 relaxed 49 80
+Metrics. For each configuration: BAR (boundary activation rate), false-denial rate (fraction of T3-APPROVED cases that become DENIED), and coverage (fraction APPROVED).
+Results.
+Table 34: Experiment 11: threshold sensitivity results across five configurations.
+Dataset: 20 cases from Experiment 9 baseline.
+False-denial rate: fraction of T3-APPROVED cases that become DENIED under this configuration.
+Config Label APPROVED ESCALATED DENIED BAR False-Denial Rate Coverage
+T1 strict 5 5 10 0.75 0.00 0.25
+T2 moderate– 5 7 8 0.75 0.00 0.25
+T3 default 6 7 7 0.70 0.00 0.30
+T4 moderate+ 6 7 7 0.70 0.00 0.30
+T5 relaxed 8 6 6 0.60 0.00 0.40
+Findings. Three results emerge:
+  1. 1.
+    False-denial rate is zero across all configurations.
+No T3-APPROVED case becomes DENIED under any threshold variant, including the strictest configuration (T1: APPROVED $\leq 29$).
+Stricter thresholds promote cases from APPROVED to ESCALATED, not directly to DENIED, because the dataset contains no cases with RS in the intermediate ranges introduced by T1/T2.
+  2. 2.
+    BAR decreases monotonically as thresholds relax.
+BAR falls from 0.75 (T1) to 0.60 (T5), confirming that the enforcement boundary captures fewer boundary-activating requests as thresholds relax.
+T3 and T4 produce identical results (BAR=0.70), demonstrating that a +5+5 point relaxation has no effect on this dataset.
+  3. 3.
+    T3 (default) is a local optimum.
+T3 is the strictest configuration that achieves false-denial rate zero while maintaining BAR =0.70=0.70.
+Further tightening (T1, T2) increases BAR to 0.75 at the cost of reclassifying legitimate cases as ESCALATED.
+Relaxing (T5) reduces BAR to 0.60 without reducing false-denial rate.
+The results justify the default threshold selection empirically.
+17.10  Experiment 12: Multi-Tool Agent Admission Control Under IPI
+Motivation.
+A realistic LLM agent operates a portfolio of tools with heterogeneous risk profiles.
+An indirect prompt injection (IPI) attack may induce the agent to flood a single high-risk tool (e.g.
+fund_transfer), triggering ACP’s cooldown and affecting the entire agent.
+This experiment asks: does ACP’s stateful ledger create persistent enforcement consequences that outlast the attack window?
+A stateless engine would fully reset after cooldown expiry; ACP’s 24-hour denial history should keep the enforcement posture elevated.
+Setup. A single agent (agent-exp12-multitool) operates four tools whose RS values under ACP-RISK-3.0 default policy with a clean ledger are:
+Table 35: Tool risk scores used in Experiment 12 (clean-state, ACP-RISK-3.0 default).
+Tool Capability Resource Class Base RS Decision
+weather_query acp:cap:data.read PUBLIC 0 APPROVED
+user_profile acp:cap:data.read SENSITIVE 15 APPROVED
+system_audit acp:cap:admin.manage PUBLIC 60 ESCALATED
+fund_transfer acp:cap:financial.transfer RESTRICTED 80 DENIED
+The session is divided into three phases using distinct timestamps to control which events fall inside the 10-minute cooldown detection window versus the 24-hour $F_{\text{anom}}$ Rule 2 window:
+[…]
+[…]
+  • Phase C ($t_{0}+60\,\mathrm{min}$, 10 requests): Recovery — same diverse mix as Phase A.
+Cooldown has expired ($300\,\mathrm{s}\ll 60\,\mathrm{min}$); $F_{\text{anom}}$ Rule 2 still active ($24\,\mathrm{h}\gg 60\,\mathrm{min}$).
+State (denial history, cooldown, pattern counts) persists across all phases via a shared InMemoryQuerier. BAR-Monitor settings: window =40=40, threshold =0.10=0.10, trend threshold =0.15=-0.15.
+Results.
+Table 36: Experiment 12: three-phase admission control results. Adversary model: A == (formula-aware, flood, B=8B{=}8).
+Phase Requests APPROVED ESCALATED DENIED BAR
+A (Baseline) 10 7 2 1 0.30
+B (IPI Chain) 8 0 0 8 1.00
+C (Recovery) 10 7 0 3 0.30
+Total 28 14 2 12 0.50
+Phase-by-phase analysis.
+Phase A (Baseline): Per-request enforcement operates normally.
+fund_transfer (RS =80=80) is denied; system_audit (RS =60=60) is escalated, not denied — it remains a borderline capability that can still execute under human review.
+BAR =0.30=0.30 reflects two escalated and one denied request out of ten.
+Phase B (IPI Chain): The attacker induces two additional fund_transfer denials.
+This raises the rolling 10-minute denial count to three (one from Phase A plus two from Phase B), satisfying the cooldown trigger threshold (CooldownTrigger =3=3).
+ACP immediately enters an agent-wide lockdown: all subsequent requests — including system_audit — are denied with reason COOLDOWN_ACTIVE, regardless of their individual RS values.
+BAR =1.00=1.00; the BAR-Monitor window captures the full attack.
+A stateless engine would deny fund_transfer only; ACP denies all six remaining operations.
+Phase C (Recovery): Cooldown has expired at $t_{0}+60\,\mathrm{min}$.
+However, the ledger retains the denial history from Phases A and B: CountDenials(agentID, 24 h) $3\geq 3$, activating $F_{\text{anom}}$ Rule 2 (+15+15) on every request.
+The RS of each tool rises accordingly:
+  • weather_query: $0+15=15$ → APPROVED ($\leq 39$)
+  • user_profile: $15+15=30$ → APPROVED ($\leq 39$)
+  • system_audit: $60+15=\mathbf{75}$DENIED ($\geq 70$)
+  • fund_transfer: $80+15=95$ → DENIED ($\geq 70$)
+The critical finding is that system_audit — which was merely escalated (RS =60=60) in Phase A — becomes denied (RS =75=75) in Phase C.
+The +15+15 anomaly penalty crosses the denial threshold without any new attack activity: the ledger’s 24-hour memory propagates the attack’s consequences into the post-cooldown recovery period.
+BAR =0.30=0.30, structurally identical to Phase A, but the composition has shifted: ESCALATED =0=0 (was 2) and DENIED =3=3 (was 1), confirming elevated risk posture.
+Findings.
+  1. 1.
+    Agent-wide cooldown amplifies IPI impact.
+After three denials in ten minutes, ACP blocks all agent operations, not only the targeted tool.
+This prevents an attacker from using IPI to saturate high-risk tool capacity while low-risk operations continue unmolested.
+  2. 2.
+    Stateful Fanom elevates borderline capabilities post-attack.
+A capability that was ESCALATED under normal conditions (system_audit, RS =60=60) becomes DENIED after an IPI campaign because the ledger’s 24-hour denial window persists beyond cooldown.
+A stateless engine would return to pre-attack RS values immediately; ACP maintains elevated enforcement for the duration of the anomaly window.
+  3. 3.
+    Low-risk tools remain unaffected.
+weather_query and user_profile remain APPROVED even with $F_{\text{anom}}+15$ applied (RS =15=15 and $30$ respectively), demonstrating that persistent post-attack elevation is proportionate: it promotes borderline cases to DENIED without blocking safe operations.
+17.11  Experiment 13: Bounded Coordination Window
+A natural question about per-agent enforcement is whether coordination across agents allows participants to collectively exceed the constraints that apply to each individually.
+This experiment quantifies the coordination window (CW)—the total number of approved actions executed across $N$ agents before ACP triggers the first DENIED decision—and evaluates how CW scales with $N$.
+Setup.
+We deploy $N$ agents issuing homogeneous acp:cap:financial.transfer requests on the shared public resource accounts/shared-ops.
+The baseline risk score is $B=35$, $F_{\text{res}}=0$ (public), giving RS =35=35, which is APPROVED ($\leq 39$).
+Anomaly thresholds are set low to expose the coordination boundary quickly: $N_{\mathrm{Rule1}}=2$ (Rule 1 fires when pattern count >2>2 in 60 s, adding +20+20) and $Y_{\mathrm{Rule3}}=2$ (Rule 3 fires when pattern count $\geq 2$ in 5 min, adding +15+15).
+Risk accumulation uses agent-scoped pattern keys of the form $\mathtt{PatternKey}=\mathrm{SHA\text{-}256}(\mathit{agentID}\mid\mathit{capability}\mid\mathit{resource})$, isolating each agent’s behavioral history independently.
+Crucially, ACP follows an evaluate-then-mutate execution contract (ACP-RISK-3.0 §4): ledger state is updated only after the decision is emitted.
+This means each agent’s first two requests are evaluated against zero and one prior pattern counts respectively, neither of which reaches the Rule 3 threshold ($\geq 2$); the third request is evaluated against two prior patterns (Rule 3 fires, RS =50=50, ESCALATED); the fourth is evaluated against three (Rule 1 and Rule 3 fire, RS =70=70, DENIED).
+Each agent therefore contributes exactly two approved actions before escalation and denial.
+Results.
+Table 37 shows the coordination window across three variants: sequential execution ($N=1$), round-robin scheduling ($N=2,3,5$), and burst execution ($N=5$, each agent completing its window before the next starts).
+Variant NN $\mathrm{CW}_{\mathrm{appr}}$ $\mathrm{CW}_{\mathrm{total}}$ TTB (reqs)
+Sequential 1 2 3 4
+Round-robin 2 4 6 7
+Round-robin 3 6 9 10
+Round-robin 5 10 15 16
+Burst (5 agents) 5 10 15 4
+Table 37: Experiment 13: coordination window across NN agents.
+$\mathrm{CW}_{\mathrm{appr}}$: total approved actions before first DENIED.
+$\mathrm{CW}_{\mathrm{total}}$: approved + escalated.
+TTB: total request count at the first DENIED.
+Linearity. Across all configurations, $\mathrm{CW}_{\mathrm{appr}}/N=2.0$ exactly (zero deviation). The result admits the closed-form expression:
+$\mathrm{CW}_{\mathrm{appr}}=2N,\quad\mathrm{CW}_{\mathrm{total}}=3N,\quad\mathrm{TTB}_{\mathrm{rr}}=3N+1,$
+where $\mathrm{TTB}_{\mathrm{rr}}$ applies to round-robin scheduling.
+Under burst execution, the first DENIED occurs at a fixed TTB =4=4 regardless of $N$, because it is triggered by the first agent exhausting its individual window.
+The temporal profile therefore depends on scheduling strategy, but the total coordination window is identical: $\mathrm{CW}_{\mathrm{appr}}=2N$ in all cases.
+Per-agent contribution is symmetric: in the $N=3$ round-robin trace, each agent records exactly two APPROVED and one ESCALATED decision before reaching DENIED, with no cross-agent contamination of risk signals.
+Analysis.
+The exact linearity follows from two structural properties of ACP-RISK-3.0.
+First, all behavioral signals (pattern counters, denial history, cooldown state) are indexed by agentID: agents do not contribute to each other’s risk profiles, and coordinated execution does not accelerate boundary activation for any individual participant.
+Second, the evaluate-then-mutate contract guarantees each agent a fixed per-request admitted count before its pattern frequency crosses the anomaly threshold—a bound determined by execution semantics, not only by numerical thresholds.
+ACP does not detect or model coordination. Instead, it enforces per-agent behavioral bounds such that coordinated activity scales at most linearly with the number of participants:
+$\text{Total executable actions}\in O(N).$
+Implications.
+Scaling a coordinated campaign requires proportional scaling of agents: each additional participant adds a fixed, bounded contribution.
+Superlinear or exponential amplification through coordination is structurally prevented.
+This result complements Experiment 9 (deviation collapse: inadmissible conditions disappear under upstream drift) and the TLA+ verification (4.29 billion states, zero safety violations: failure states are unreachable): together they demonstrate that ACP constrains both the existence and the scale of undesirable behaviors.
+17.12  Experiment 14: Stateless vs. Stateful Enforcement — OPA Comparison
+While Experiment 13 establishes that coordination remains bounded under ACP, it does not address whether similar constraints can be expressed in existing policy systems.
+We now examine this question by comparing ACP against a representative stateless policy engine (OPA), focusing on expressiveness rather than performance.
+Motivation.
+The Related Work discussion (§2.4) argues that stateless policy engines cannot enforce temporal behavioral constraints without external state.
+This experiment operationalizes that claim empirically using Open Policy Agent (OPA v1.15.2, Rego v1) [36] as a representative stateless engine.
+This experiment is a controlled comparative construction designed to isolate differences in expressiveness rather than a statistical benchmark.
+The comparison is not performance-oriented; it is a capability comparison.
+The central question is: can a stateless evaluator reproduce ACP’s enforcement behavior over an execution trace, without auxiliary infrastructure?
+Setup.
+All three scenarios use the same request profile as Experiment 13: acp:cap:financial.transfer on a public resource (baseline $RS=35$, below $\mathrm{ApprovedMax}=39$).
+Three systems are evaluated in parallel: (1) ACP with full stateful enforcement (PatternKey ledger, Fanom accumulation, cooldown); (2) ACP stateless baseline (StatelessEngine with NullQuerier: Fanom = 0, no cooldown); and (3) OPA with a pure stateless policy and, separately, with injected external state.
+Scenario A: Single-request admissibility.
+A single request with no prior history.
+ACP stateful: Approved ($RS=35$).
+ACP stateless: Approved ($RS=35$, identical).
+OPA pure: allow=true.
+All three agree.
+No behavioral state is required for single-request admission decisions; stateless evaluation is sufficient here.
+Scenario B: Frequency accumulation (10 requests).
+Ten consecutive requests from the same agent to the same resource.
+ACP stateful transitions: Approved $\times 2$ $\to$ Escalated $\times 1$ $\to$ Denied $\times 3$ $\to$ Cooldown_Active $\times 4$ (first escalation: request #3; first denial: request #4).
+ACP stateless: Approved $\times 10$ (Fanom = 0 throughout).
+OPA pure: allow=true $\times 10$.
+Note that Cooldown_Active responses are counted as denied outcomes in the aggregate statistics.
+When OPA is augmented with an externally injected request_count field, it correctly denies requests once the counter reaches threshold.
+However, the caller is responsible for maintaining and incrementing the counter, there is no atomicity guarantee between evaluation and the counter update, and the state store is entirely external to the policy engine.
+This is not a limitation of policy syntax, but of system architecture: OPA itself does not provide guarantees about the consistency, atomicity, or timing of externally managed state.
+Scenario C: Cooldown enforcement.
+A sequence of requests on a restricted resource ($RS=80$, Denied always).
+ACP stateful: Denied $\times 3$ $\to$ Cooldown_Active $\times 3$ (cooldown triggered after CooldownTriggerDenials = 3).
+ACP stateless: Denied $\times 6$—the stateless engine re-evaluates each request independently and never enters cooldown, missing the short-circuit enforcement semantic.
+OPA pure does not apply here: OPA evaluates individual requests and has no mechanism to detect that three consecutive denials have occurred or to enter a cooldown state.
+When augmented with an externally injected cooldown_active field, OPA blocks correctly (allow=false), but only when the caller manages the cooldown timer and injects its current value at every evaluation call.
+These constraints cannot be expressed in a purely request-scoped evaluation model without external coordination across decision boundaries.
+Scenario ACP (stateful) ACP (stateless) OPA (pure)
+A: Single request ✓ Approved ✓ Approved ✓ allow=true
+[…]
+C: Cooldown (temporal) ✓ Cooldown $\times$ re-eval each request $\times$ no cooldown concept
+Table 38: Capability matrix: stateful vs.
+stateless enforcement over a 10-request trace.
+✓ = correct enforcement; $\times$ = cannot enforce without external state.
+ACP stateless and OPA pure are structurally equivalent: both evaluate each request in isolation with $F_{\mathrm{anom}}=0$.
+Latency (supporting data).
+ACP Evaluate() measures $\sim\!852$ ns/op over 50,000 iterations ($\sim\!739$ ns p50 from the established benchmark, Table 19).
+OPA bench measures $\sim\!16{,}000$ ns/op ($\sim\!19\times$ higher).
+This gap reflects the evaluation model difference (compiled Go vs.
+Rego interpreter), not algorithmic complexity.
+Latency measurements are included as supporting data only and should not be interpreted as a direct performance comparison.
+Summary.
+Scenarios A, B, and C together confirm the expressiveness boundary: stateless evaluation is sufficient for single-request decisions, but insufficient for any constraint that depends on the execution trace.
+OPA can approximate both Scenario B and C when supplied with externally maintained state, but this approximation requires infrastructure external to the policy engine, introduces a split between evaluation and state mutation, and shifts correctness responsibility to the surrounding architecture.
+ACP enforces these constraints natively, as first-class properties of the admission control protocol, without auxiliary state management.
+Even when external state is introduced, enforcement remains split across multiple components, and no single point in the system guarantees that evaluation and state mutation occur atomically at the decision boundary.
+The contribution is not the observation that state is required, but that enforcement must occur at a boundary where both evaluation and state transition are jointly determined.
+This distinction aligns with the formal analysis in §2.4, where enforcement properties depend on the availability of state transitions within the model.
+17.13  False-Denial Rate Analysis
+A standard objection to stateful enforcement is that history-aware scoring may penalize legitimate activity occurring after a past incident—producing false denials for requests that would be admitted under a clean ledger.
+We quantify this risk using evidence from Experiments 11 and 12.
+Stateless baseline: zero false denials (Experiment 11).
+Experiment 11 evaluates five threshold configurations (T1–T5) across a 20-request heterogeneous workload.
+The false-denial rate is 0.00 across all configurations, including the strictest setting (T1: APPROVED $\leq 29$).
+No T3-APPROVED request is reclassified as DENIED under any threshold variant; tighter thresholds promote requests from APPROVED to ESCALATED, not directly to DENIED, because ACP’s three-outcome model introduces a graduated step before hard denial.
+Post-attack: proportionate elevation (Experiment 12, Phase C).
+Phase C of Experiment 12 is the strongest available test of false-denial risk: the ledger carries three prior denials ($F_{\text{anom}}+15$ from Rule 2), and cooldown has just expired.
+Low-risk tools remain fully accessible: weather_query (RS =0+15=15=0+15=15) and user_profile (RS =15+15=30=15+15=30) are both APPROVED ($\leq 39$), yielding a post-attack false-denial rate of 0.00 for safe capabilities.
+The only tool promoted to DENIED is system_audit (RS =60+15=75=60+15=75), which was ESCALATED under the clean-state policy.
+This is not a false denial: system_audit carries acp:cap:admin.manage, a privileged capability that requires human review under default policy even without any prior incident.
+After a confirmed IPI campaign, elevating it to DENIED is intentional tightening, not a spurious penalty.
+The same behavior is documented in §13.8 as a known trade-off of stateful enforcement.
+Design basis.
+The zero false-denial property in the stateless baseline is not coincidental: ACP’s three-tier outcome model (APPROVED / ESCALATED / DENIED) provides a graduated buffer.
+Stricter configurations first escalate, and only hard-deny requests whose RS already falls in the upper range.
+In the post-attack regime, the +15+15 anomaly penalty is calibrated to cross the denial threshold only for capabilities already classified as borderline (ESCALATED, RS $\approx 55$$69$), leaving low-risk capabilities untouched.
+High-risk capabilities (fund_transfer, RS $\geq 70$ clean) are denied in both regimes—the anomaly penalty adds margin, not a new category.
+17.14  Extended Governance Objective
+Bounded operational region.
+The extended governance objective emerges from two complementary constraints.
+Cooldown enforces an upper bound: excessive inadmissible activity triggers containment.
+BAR-Monitor enforces a lower bound: insufficient boundary interaction signals a degenerate regime.
+A deployment satisfying both constraints operates within a bounded region where governance is simultaneously correct and meaningful.
+The standard ACP objective is admission correctness: for every request $s$ presented to the engine, $G(a,s,h)$ returns the correct decision given the configured policy and observed state.
+We propose an extended governance objective that adds a structural requirement:
+Definition 9 (Extended Governance Objective).
+A deployment satisfies the extended governance objective if it satisfies admission correctness and failure condition preservation: there exists a non-empty set of requests that the engine would deny under faithful input conditions.
+Verifying the extended objective requires counterfactual evaluation as defined in Section 17.3.
+BAR serves as the operational metric: a deployment monitoring BAR over time can detect the onset of deviation collapse before it becomes a persistent governance condition.
+18  Deployment Considerations
+ACP does not replace higher-level coordination or monitoring systems; it provides a verifiable admission control primitive within a larger operational stack.
+This section characterizes the deployment decisions that adopters must make when integrating ACP into production infrastructure.
+18.1  State Backend Selection
+The LedgerQuerier interface is the primary deployment decision surface.
+The reference InMemoryQuerier is suitable for development, single-process testing, and conformance evaluation.
+Production deployments require an external backend with durability and horizontal scalability guarantees.
+Backend Throughput Durability Recommended use
+InMemoryQuerier $\sim$920k req/s None (process-local) Dev / conformance testing
+Redis sorted sets $\sim$4,700 req/s (unpipelined, $\sim$7–8 RTTs) AOF/RDB Multi-process, low-latency
+Redis + pipelining $\sim$8,000 req/s (2 RTTs, Exp. 3) AOF/RDB High-throughput production
+Postgres / MySQL Backend-dependent Full ACID Audit-grade persistence
+Table 39: LedgerQuerier backend options.
+The admission control logic (pkg/risk.Evaluate) is stateless and backend-agnostic; all state transitions are delegated to the LedgerQuerier abstraction.
+Redis throughput from Experiment 3 (Docker loopback, Intel i7-8665U, Go 1.22).
+Pipelining reduces RTTs from $\sim$7–8 to 2; see Table 18.
+The key architectural property is that the ACP decision function is compute-cheap and the throughput ceiling is set entirely by the state backend.
+Replacing the backend does not require modifying the admission control logic or redeploying the protocol layer.
+Evaluation on multi-node and geo-replicated configurations (e.g., CockroachDB, Spanner) is planned; the LedgerQuerier abstraction is designed to accommodate them without modification to the protocol layer or admission logic.
+Deployment Maturity Model.
+The three backend tiers correspond to a natural maturity progression. Adopters should advance tiers as their operational requirements grow:
+Table 40: ACP deployment maturity tiers.
+Each tier adds durability, observability, and assurance guarantees.
+BAR-Monitor is mandatory at Tier 3 because tamper-evident ledgers justify proactive governance monitoring.
+Tier Backend BAR-Monitor Ledger integrity Recommended context
+1 InMemoryQuerier Optional Ephemeral (process-local) Development, conformance, CI
+2 Redis pipelined Recommended AOF/RDB persistent Production, multi-process
+3 Postgres / ACID Required Hash-chained audit High-assurance, regulated industries
+Tier 3 pairs ACP with a tamper-evident state backend (ACP-LEDGER-1.0, planned) and mandates BAR-Monitor to detect deviation collapse (§17.6).
+Tier 1 and Tier 2 deployments can enable BAR-Monitor incrementally without changing the admission control logic.
+18.2  Agent Identity Provisioning
+ACP requires a stable agent_id per agent instance.
+Provisioning can be integrated with existing IAM or service mesh identity systems: the agent_id is an opaque string from ACP’s perspective, provided it is unique, stable across requests, and included in signed capability tokens.
+Key pairs (Ed25519 for L1–L4) should be provisioned at agent registration time and rotated according to the organization’s key management policy.
+18.3  Multi-Organization Boundaries
+Each organization deploys its own ACP instance with a sovereign Institutional Key.
+Admission control decisions are evaluated locally against the deploying organization’s LedgerQuerier and PolicyConfig.
+Cross-organization request flows are handled by ACP-CROSS-ORG-1.0: the receiving organization evaluates the incoming execution token independently, applying its own risk policy and local ledger state.
+Federated deployments use mutual recognition (ACP-ITA-1.1) to establish cross-org trust anchors without requiring a shared root of trust.
+The per-org isolation model is intentional: an organization’s admission control decisions are not visible to or affected by another organization’s ledger.
+18.4  Cross-Agent Coordination (L3)
+ACP enforces admission control per agent and per request.
+When multiple agents collaborate on a composite task, each agent’s admission request is evaluated independently against its own agent_id, capability token, and ledger history.
+ACP does not implement cross-agent correlation or orchestration-level coordination.
+This is an explicit scope boundary: as demonstrated in Experiment 2 (§13.7), a coordinated group of $N$ agents can collectively execute $3N$ high-risk requests before all agents are individually blocked.
+Mitigating coordinated multi-agent threats requires attribution at the orchestration layer (e.g., shared-resource rate limits, agent group policies, identity clustering), which ACP is designed to complement, not replace.
+The ACP-CROSS-ORG-1.0 interaction model provides a protocol-level primitive for multi-organization agent coordination; full L3 cross-agent orchestration is a deployment-layer concern outside the scope of this specification.
+18.5  Integration with Existing Infrastructure
+ACP occupies the Execution Governance layer of the Agent Governance Stack (Section 1.5). It is designed to complement—not replace—existing access control and observability infrastructure:
+  • RBAC / ABAC / OPA: ACP provides per-request admission control with temporal state (history, cooldown). Existing resource-authorization systems continue to apply after ACP admission succeeds.
+  • Zero Trust: Environmental signals (external_ip, off_hours, geo_outside, untrusted_device) integrate Zero Trust context directly into the ACP risk score via F_ctx.
+  • SIEM / Observability: The Audit Ledger (ACP-LEDGER-1.3) produces per-decision signed records suitable for ingestion into log aggregation pipelines.
+Each entry carries a policy_hash enabling post-hoc attribution of decisions to specific policy versions.
+  • Monitoring systems: ACP’s cooldown and escalation events are observable state transitions—not silent blocks. They are designed to be surfaced to monitoring systems for incident response.
+18.6  Policy Tuning
+The PolicyConfig parameters (defaults documented in ACP-RISK-2.0 Appendix A and exposed via DefaultPolicyConfig()) are the primary operational levers.
+Table 41 provides concrete starting values for four threat profiles; organizations should adjust from these baselines using observed denial rates and throughput data.
+Table 41: Recommended PolicyConfig starting values by threat profile.
+All profiles use PolicyHash pinning and APPROVED $\leq 39$ / DENIED $\geq 70$ thresholds (ACP-RISK-3.0 defaults).
+Expected BAR baseline is the healthy operational range for a deployment at that profile under normal (non-attack) traffic; sustained BAR below the lower bound indicates deviation collapse (§17.6).
+Polling agents may require higher $Y$ to suppress false escalations from legitimate repeated reads.
+Profile CooldownPeriod CooldownTrigger AnomalyRule3Y Expected BAR Typical context
+Low 120 s 5 5 0.80–0.95 Internal dev tools, low-risk automation
+Medium 300 s 3 3 0.60–0.80 B2B APIs, typical SaaS agents (default)
+High 600 s 2 2 0.45–0.65 Financial systems, regulated industries
+Critical 3600 s 1 1 0.30–0.50 Critical infrastructure, high-assurance
+The four parameters interact: tightening CooldownTrigger without extending CooldownPeriod increases sensitivity but shortens the containment window.
+AnomalyRule3Y should match the expected request burst size of legitimate agents—setting $Y=1$ blocks any repeated capability use, which is correct for critical infrastructure but would produce false escalations in high-frequency automation.
+PolicyHash should be pinned in all profiles to prevent silent policy drift.
+  • CooldownPeriodSeconds: Shorter (<300<300 s) enables faster recovery after a false-positive denial cluster; longer ($\geq 600$ s) provides stronger containment under sustained attack or compromised agent.
+  • CooldownTriggerDenials: Lower values detect attacks faster but increase sensitivity to legitimate burst traffic.
+  • AnomalyRule3ThresholdY: Default $Y=3$ balances replay detection latency against false positives from polling agents.
+  • PolicyHash: Pin policy versions in capability tokens to prevent silent policy changes from altering admission decisions post-issuance.
+ACP provides sensible defaults via DefaultPolicyConfig() (Appendix A of ACP-RISK-2.0), corresponding to the Medium profile above.
+Organizations should treat policy calibration as an operational process, informed by observed denial rates, throughput baselines, and threat model.
+18.7  Migrating from ACP-RISK-2.0 to ACP-RISK-3.0
+ACP-RISK-3.0 is backward-compatible with ACP-RISK-2.0 at the API level.
+The EvalRequest and EvalResult structures are unchanged; all existing conformance vectors and capability tokens remain valid.
+The only behavioral change is the scoping of Rule 1 of $F_{\text{anom}}$:
+Table 42: ACP-RISK-2.0 vs.
+ACP-RISK-3.0: behavioral and interface differences.
+All other fields, formulas, and thresholds are identical.
+Existing deployments need only provide resource in EvalRequest and upgrade the pkg/risk package.
+Aspect ACP-RISK-2.0 ACP-RISK-3.0
+Rule 1 counter key agentID PatternKey(agentID, cap, resource)
+Cross-context isolation No (state-mixing) Yes (per-context scoping)
+EvalRequest fields unchanged resource used for PatternKey
+Ledger key format per-agent per-context (SHA-256 hash)
+Performance overhead baseline ++50 ns per evaluation (Table 12)
+Conformance vectors 65 unsigned 73 signed (superset; both suites pass)
+Migration steps.
+  1. 1.
+    Upgrade pkg/risk to the ACP-RISK-3.0 version.
+  2. 2.
+    Ensure EvalRequest.Resource is populated per request (it was optional in RISK-2.0; it is required for correct PatternKey computation in RISK-3.0).
+  3. 3.
+    Flush the anomaly counters in the ledger (optional but recommended): RISK-2.0 per-agent keys are structurally different from RISK-3.0 per-context keys and will not interfere, but stale RISK-2.0 entries consume storage.
+  4. 4.
+    Re-run the ACR-1.0 conformance suite against your deployment. All 73 signed vectors must pass; any failure indicates a RISK-2.0 residue.
+Deployments that cannot populate Resource immediately (e.g., legacy agents without resource attribution) may set Resource = agentID as a transitional measure.
+This degrades RISK-3.0 to RISK-2.0 behavior for Rule 1 while preserving all other RISK-3.0 properties (Rules 2 and 3, cooldown, thresholds).
 19  Security Model
@@
 19.1  Adversary Capabilities
-We consider a probabilistic adversary $\mathcal{A}$ interacting with the system through agent requests. The adversary is assumed to have the following capabilities:
-  • Request control: $\mathcal{A}$ can generate arbitrary sequences of requests on behalf of one or more agents, including high-frequency and adversarial patterns.
-  • Adaptive behavior: $\mathcal{A}$ can adapt future requests based on past responses (approval/denial outcomes), enabling iterative probing strategies.
-  • Multi-agent coordination: $\mathcal{A}$ may coordinate multiple agents to attempt distributed evasion, including alternating identities or cross-agent patterns.
-  • Input manipulation: $\mathcal{A}$ can choose all request fields (capability, resource, context) within the bounds of the protocol specification.
+We formalize the adversary as a triple $\mathcal{A}=(K,S,B)$ where:
+  • $K\in\{\textit{black-box},\ \textit{formula-aware},\ \textit{full-state}\}$ is the knowledge level.
+A black-box adversary observes only the outcome of each request (APPROVED/ESCALATED/DENIED).
+A formula-aware adversary additionally knows the published risk formula ($\mathrm{RS}=\mathrm{capBase}+F_{\mathrm{res}}+F_{\mathrm{ctx}}+F_{\mathrm{hist}}+F_{\mathrm{anom}}$) and all weight values defined in ACP-RISK-3.0.
+A full-state adversary further has real-time read access to the entire ledger state (pattern counts, denial history, cooldown timestamps).
+  • $S\in\{\textit{threshold-hugging},\ \textit{evasion},\ \textit{flood},\ \textit{collusion}\}$ is the strategy class.
+Threshold-hugging: craft requests with $\mathrm{RS}$ just below enforced boundaries.
+Evasion: systematically drive all features toward zero to achieve permanent APPROVED.
+Flood: saturate the system with high-volume requests.
+Collusion: coordinate multiple agents to distribute request load across isolation boundaries.
+  • $B\in\mathbb{N}$ is the request budget available to the adversary.
+The evaluation experiments instantiate specific $(K,S,B)$ triples. Table 43 maps each experiment to its adversary class and reports the enforcement outcome.
+Table 43: Adversary taxonomy: mapping of ACP experiments to $(K,S,B)$ adversary classes.
+Exp Adversary Class $K$ $S$ $B$ Outcome
+1 Cooldown Evasion black-box threshold-hugging 500 contained
+2 Multi-Agent Flood black-box flood+collusion 1000 contained
+3 Backend Stress black-box flood 10K degraded perf
+4 Token Replay black-box evasion 300 contained
+5–8 State Mixing formula-aware evasion fixed in RISK-3.0
+9 Deviation Collapse upstream drift BAR detects
+10 Knowledge-Aware Evasion formula-aware evasion (RS=0) 200 BAR detects
+11 Threshold Sensitivity monotonic
+12 Multi-Tool IPI Chain formula-aware flood 8 stateful persistence
+13 Coordination Window black-box round-robin/burst $5N$ linear bound ($O(N)$)
+[…]
+Five findings emerge from this taxonomy:
+  1. 1.
+    Per-decision enforcement suffices against black-box adversaries. Experiments 1–4 confirm that stateful admission control contains all black-box strategies within the request budget.
+  2. 2.
+    Formula-aware adversaries break per-decision enforcement but are detected by BAR.
+Experiment 10 shows that a formula-aware adversary can achieve $\mathrm{RS}=0$ on every request, causing every decision to be APPROVED — yet BAR-Monitor fires a TREND alert at batch 2, before the regime reaches full collapse.
+  3. 3.
+    Deviation collapse (Exp 9) is an environmental threat, not adversarial, but BAR-Monitor handles both classes with the same mechanism.
+  4. 4.
+    Stateful ledger creates persistent enforcement consequences (Exp 12).
+An IPI flood that triggers cooldown also activates the 24-hour anomaly window ($F_{\text{anom}}$ Rule 2), elevating borderline capabilities from ESCALATED to DENIED for the entire anomaly window after cooldown expires.
+A stateless engine would return to pre-attack RS values immediately.
+  5. 5.
+    Full-state adversaries are outside the v1.0 threat model.
+An adversary with real-time ledger access could predict and exploit threshold transitions.
+This class is not addressed in the present specification and is acknowledged in §13.8 as a target for ACP-D-1.0 (decentralized governance).
+  6. 6.
+    Coordinated multi-agent activity scales linearly (Exp 13).
+Per-agent behavioral isolation ensures that $N$ agents acting in concert can execute at most $2N$ approved operations before the first DENIED decision—a coordination window that grows linearly with $N$, with zero deviation observed.
+Superlinear amplification through coordination is structurally prevented by the combination of agent-scoped pattern keys and the evaluate-then-mutate execution contract.
 19.2  Adversary Limitations
@@
 Determinism.
-For a given request and system state, the evaluation function produces a unique and reproducible decision outcome.
-This prevents adversarial exploitation of non-determinism.
-The RiskDeterminism invariant in the TLA+ model formally specifies this property.
+ACP determinism is defined with respect to the evaluation context: given identical canonical request inputs, system state (ledger, pattern counters, cooldown state), policy definitions, and evaluation order, the protocol produces a consistent and reproducible admission decision.
+This form of determinism explicitly includes state as part of the input to the decision function; it does not imply independence from time or state evolution.
+This scoping prevents adversarial exploitation of non-determinism across evaluators.
+The RiskDeterminism and AnomalyDeterminism invariants in the TLA+ model formally specify this property across all reachable states.
 State Consistency.
@@
 This property bounds the impact of a compromised or malicious agent to its own request history.
+Note that isolation holds across agents but not across capability contexts within the same agent: anomaly signals from F_anom Rules 1 and 2 aggregate at the agent level, enabling cross-context interference for agents holding multiple capability tokens under the same keypair (see §13.8, “State-Mixing and Context Fragmentation”).
 19.4  Informal Security Argument
 Because ACP separates stateless evaluation from state management and enforces a strict execution contract, every request results in a deterministic state transition that cannot be bypassed.
-The append-only nature of the state backend ensures that historical information accumulates monotonically: cooldown activation follows deterministic threshold conditions applied to accumulated state; repeated probing eventually triggers enforced denial periods.
+When the state backend enforces append-only semantics, historical information accumulates monotonically: cooldown activation follows deterministic threshold conditions applied to accumulated state; repeated probing eventually triggers enforced denial periods.
 Because decisions are reproducible given the same input sequence and initial state, any observed behavior can be externally verified.
@@
   • Compromise of the state backend (deletion or modification of history) breaks core guarantees.
+  • Distributed deployments: ACP’s determinism guarantee holds per evaluator instance.
+In deployments with replicated or eventually-consistent state backends (e.g., Redis Cluster, Postgres streaming replicas), two evaluators may observe divergent ledger states and reach different admission decisions for the same agent at the same point in time.
+ACP does not specify a distributed consensus protocol; deployments requiring strict cross-instance consistency must enforce strong consistency at the storage layer.
   • Incorrect implementation of the execution contract may lead to inconsistent behavior.
   • The model does not address side-channel attacks or network-level adversaries.
-  • Post-quantum signature migration (ACP-SIGN-2.0 HYBRID, Ed25519 $\to$ ML-DSA-65) is specified and stubbed; full implementation is deferred to v1.20.
-ACP is therefore best understood as a control-layer mechanism that assumes a trusted execution and storage environment, while providing strong guarantees on decision consistency and temporal behavior under adversarial inputs.
-17  Related Work
-17.1  Access Control and Policy Enforcement
-Traditional access control models such as Role-Based Access Control (RBAC) [12] and Attribute-Based Access Control (ABAC) [6] provide structured mechanisms for regulating access decisions based on predefined roles or attributes.
-Policy languages such as XACML [9] and modern policy engines such as Open Policy Agent [10] and Cedar [1] extend these paradigms with expressive evaluation engines.
-These systems are fundamentally designed around stateless decision evaluation: they evaluate a request against policy at a single point in time.
-They do not define how system state evolves as a result of decisions over time, and they do not formalize how historical behavior (repeated denials, temporal patterns) influences subsequent decisions in a deterministic and auditable manner.
-ACP differs by explicitly separating: (i) stateless decision evaluation, (ii) externalized state management, and (iii) an execution contract governing state evolution across requests.
-This enables reasoning not only about individual decisions, but about the temporal behavior of the admission system.
-17.2  Runtime Policy Systems and Agent Frameworks
-Recent protocols for autonomous agents and multi-agent coordination define structured interactions between components and services.
-The Model Context Protocol (MCP) [2] provides structured tool access between LLM applications and services.
-Agent-to-Agent (A2A) [4] defines communication and task delegation between agents.
-While these protocols address interoperability and orchestration, they do not define:
-  • a deterministic risk evaluation model that accumulates state across requests,
-  • explicit state evolution semantics with a testable execution contract, or
-  • externally verifiable execution traces that can be reproduced by third parties.
-Recent work has begun to address security governance for LLM-based agentic systems.
-Syros et al.
-[15] propose SAGA, a security architecture that provides user-controlled access policies and cryptographic mediation for agent-to-agent tool use and inter-agent communication.
-ACP complements this line of work by focusing on deterministic, auditable admission control at the individual request level: rather than governing inter-agent communication channels, ACP enforces stateful risk evaluation and cooldown constraints on each action request, with an externally verifiable execution contract and formal invariants.
-ACP complements these approaches by providing a governance layer that is independent of orchestration protocols, focusing specifically on verifiable decision enforcement and auditability.
-17.3  Auditability and Tamper-Evident Logs
-Secure audit logging [13] and event sourcing [3] ensure that past events can be reconstructed and verified after the fact.
-These approaches operate post hoc: they provide evidence of what occurred without constraining system behavior in real time.
-ACP integrates auditability directly into the admission process.
-Every decision—APPROVED, ESCALATED, or DENIED—is recorded in an append-only, cryptographically chained ledger (ACP-LEDGER-1.3) before the governed action executes.
-This enables both real-time enforcement and post-hoc verification from the same artifact.
-The LedgerAppendOnlyTemporal invariant in the TLA+ model formally specifies that ledger entries are never modified or removed.
-17.4  Formal Verification of Protocols
-Model checking [8] and symbolic reasoning are used to establish correctness properties in distributed systems and protocols, typically focusing on safety and liveness of state transitions.
-ACP incorporates formal methods in a targeted manner: the TLA+ model (tla/ACP.tla) mechanically checks three critical invariants (fail-closed admission, append-only ledger, risk determinism) using TLC.
-Rather than attempting full-system verification, ACP adopts a pragmatic approach in which formally verified invariants are instantiated as concrete sequence test vectors and validated at runtime by the ACR-1.0 compliance runner.
-This creates a three-layer verifiability chain: formal $\to$ data $\to$ runtime.
-17.5  Summary
-Table 20 summarizes the properties of ACP relative to existing approaches.
-Property RBAC ABAC OPA Cedar MCP/A2A ACP
-Stateful decisions $\times$ $\times$ $\times$ $\times$ $\times$
-Deterministic evaluation $\times$
-Auditability partial partial partial partial $\times$
-Runtime enforcement partial
-Temporal behavior $\times$ $\times$ $\times$ $\times$ $\times$
-External verifiability $\times$ $\times$ $\times$ $\times$ $\times$
-Execution contract $\times$ $\times$ $\times$ $\times$ $\times$
-Agent-native $\times$ $\times$ $\times$ $\times$ partial
-Table 20: Comparison of ACP with existing access control and agent governance approaches. ✓ = fully supported; $\times$ = not supported; partial = partially addressed.
-Existing systems address individual aspects of the problem—policy evaluation, agent orchestration, audit logging, or formal verification.
-ACP’s contribution lies in combining these elements into a unified model centered on an explicit execution contract.
-This enables a property not directly provided by prior systems: the ability to deterministically reproduce and externally verify the behavior of an admission decision system over time.
-Unlike existing systems, ACP explicitly models temporal behavior—history accumulation, anomaly detection, and cooldown enforcement—within the admission decision itself, and provides verifiable execution semantics through signed test vectors and a runtime compliance runner.
-This combination of temporal state, runtime enforcement, and external verifiability is not jointly addressed by any prior system in Table 20.
+  • Prompt-layer attacks and LLM-specific benchmarks.
+ACP operates at the execution layer and does not prevent indirect prompt injection (IPI) [22] at the prompt layer.
+An IPI attack that causes an agent to emit a low-risk request (e.g., acp:cap:data.read/Public, $\mathrm{RS}=0$) will be admitted by ACP, because the resulting tool call is individually valid.
+ACP’s enforcement boundary is the tool call, not the instruction that produced it.
+Standard LLM agent benchmarks—AgentDojo [8], InjecAgent [51], and AgentContract-Bench—evaluate prompt-layer manipulation resistance, which is a complementary but orthogonal property to execution-layer admission control.
+Prompt-layer defenses reduce the IPI attack surface; ACP enforces structural constraints over the action sequence regardless of how the request was generated.
+The two layers are composable: a prompt-layer filter applied before the tool call reaches the ACP admission boundary reduces the frequency of injected high-risk requests, while ACP provides the structural guarantee that high-risk requests are denied regardless of their origin.
+ACP’s defense against formula-aware evasion (Experiment 10, §17.8) addresses a distinct threat class: an adversary who drives $\mathrm{RS}=0$ across all requests—which BAR-Monitor detects via $\Delta\mathrm{BAR}$, not per-decision enforcement.
+  • Post-quantum signature migration (ACP-SIGN-2.0 HYBRID, Ed25519 + ML-DSA-65) is implemented via the Cloudflare CIRCL library (pkg/sign2/).
+ML-DSA-65 signing latency is approximately 100–130 µs (vs.
+25 µs for Ed25519), remaining well below state-backend latency; see §12.2.
+ACP is therefore best understood as a control-layer mechanism whose guarantees are conditioned on the integrity of the state backend and the execution environment.
+When these components operate correctly, ACP provides strong guarantees on decision consistency and temporal behavior under adversarial inputs.
+When the state backend is compromised (e.g., history deleted by a privileged operator), the governance guarantees collapse; this is a declared residual risk, not a design oversight.
+Deployments in high-assurance environments should pair ACP with tamper-evident storage—a cryptographically chained log structure that makes deletion and modification detectable (motivating a future ACP-LEDGER-1.0 specification).
 20  Conclusion
@@
 ACP proposes that they operate with formal governance, with verifiable mechanisms, and with traceability that can withstand an external audit.
-ACP is not the first attempt to control autonomous agents.
-It is the first attempt to do so through a formal technical specification with precise state models, demonstrable security properties, and verifiable conformance requirements.
+ACP is not the only attempt to control autonomous agents.
+It advances this goal through a formal technical specification with precise state models, empirically validated security properties, and verifiable conformance requirements.
 The difference between a best-practices policy and a formal protocol is exactly that: behaviors are defined, failures have specific error codes, and conformance can be verified.
@@
 What it provides is a structured, auditable, and formally specified model for admission control, delegation, revocation, and cross-organization interaction— the layer that existing identity and policy frameworks leave unaddressed.
-The v1.19 specification is complete.
-A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (3 invariants, 0 violations), an ACP-SIGN-2.0 HYBRID mode stub with Ed25519 and ML-DSA-65 migration path, an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent killer demo (examples/payment-agent/), and an extended OpenAPI 3.1.0 specification (18 endpoints) are publicly available at https://github.com/chelof100/acp-framework-en.
-Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19219776
+The v1.25 specification is complete.
+A full Go reference implementation (23 packages, L1–L4), 73 signed + 65 unsigned RISK-2.0 conformance test vectors, 5 stateful sequence test vectors (ACR-1.0 compliance runner), a TLC-runnable TLA+ formal model (base: 3 invariants; extended v1.25: 11 invariants + 4 temporal properties, 5,684,342 states generated, 3,147,864 distinct, graph depth 15, 0 violations; two-agent safety at LB=11: 4,648,235,723 states generated, 4,294,930,695 distinct, 11 invariants, 0 violations, 10.5 h), a real ML-DSA-65 (Dilithium mode3) hybrid implementation via Cloudflare CIRCL (pkg/sign2/, SignHybridFull + VerifyHybrid), an executable multi-organization interoperability demo (examples/multi-org-demo/), a payment-agent demo (examples/payment-agent/), an extended OpenAPI 3.1.0 specification (18 endpoints), ten experiments covering performance benchmarks across backend configurations (Experiments 1–4) and adversarial evaluation: token replay (Experiment 5), stateless vs.
+stateful comparison (Experiment 6), state-mixing vulnerability characterization (Experiment 7), context-scoped anomaly enforcement under ACP-RISK-3.0 (Experiment 8), deviation collapse and restoration (Experiment 9), and knowledge-aware adversarial evasion (Experiment 10) are publicly available at https://github.com/chelof100/acp-framework-en [13].
+Preprint: https://arxiv.org/abs/2603.18829Zenodo: 10.5281/zenodo.19672575
 The specification and implementation are open for technical review, pilot implementation, and formal standardization.
 TraslaIA invites organizations interested in adopting ACP, contributing to its evolution, or participating in the standardization process to reach out directly.
+Cross-agent coordination remains a necessary extension for deployments involving adversarial multi-identity scenarios, particularly in high-integrity systems; a concrete mitigation path is described in §13.8 (Distributed Attack Surface).
 ACP demonstrates that admission control for autonomous agents can be both computationally efficient and operationally enforceable—provided that decision logic and state management are cleanly separated.
-The protocol is compute-cheap but state-sensitive: the LedgerQuerier abstraction keeps the evaluation function at $\sim$820 ns while allowing the state backend to be replaced, optimized, or scaled independently.
+The protocol is compute-cheap but state-sensitive: the LedgerQuerier abstraction keeps the evaluation function at 739–832 ns (p50; Table 12) while allowing the state backend to be replaced, optimized, or scaled independently.
 This is not an implementation detail—it is the architectural property that makes ACP deployable in production without sacrificing governance correctness.
+A system that always produces admissible actions does not demonstrate effective governance; it may instead be operating in a regime where admissibility is no longer meaningfully evaluated.
+The contribution of this work is to make that condition detectable, measurable, and formally specifiable.
 Marcelo Fernandez | TraslaIA
-info@traslaia.com | https://agentcontrolprotocol.xyz
+info@traslaia.com | https://agentcontrolprotocol.xyz [12]
 Appendix A Glossary
@@
 Table 44: ACP Glossary.
-Appendix B Formal Verification (ACP-RISK-2.0 — TLC-Runnable)
+Appendix B Prohibited Behaviors
+ACP defines 12 behaviors that no compliant implementation may exhibit. Exhibiting any of the following disqualifies conformance at all levels.
+Code Prohibited behavior
+PROHIB-001 Approving a request when any evaluation component fails
+PROHIB-002 Reusing an already-consumed Execution Token
+PROHIB-003 Omitting signature verification on any incoming artifact
+PROHIB-004 Treating a not-found token_id as active in a revocation context
+PROHIB-005 Allowing state transition from revoked
+PROHIB-006 Issuing an ET without a prior APPROVED AuthorizationDecision
+PROHIB-007 Modifying or deleting Audit Ledger events
+PROHIB-008 Silencing ledger corruption detection
+PROHIB-009 Ignoring max_depth in delegation chains
+PROHIB-010 Implementing an offline policy more permissive than ACP-REV-1.0
+PROHIB-011 Approving requests from agents with autonomy_level 0
+PROHIB-012 Continuing to process an artifact with an invalid signature
+Table 45: ACP prohibited behaviors (ACP-CONF-1.2). Exhibiting any disqualifies conformance at all levels.
+Appendix C Formal Verification (ACP-RISK-2.0 — TLC-Runnable)
 The tla/ACP.tla module is a complete TLC-runnable TLA+ model of the ACP-RISK-2.0 admission control engine.
-It checks three invariants over a bounded state space (2 agents $\times$ 4 capabilities $\times$ 3 resources, ledger depth $\leq 5$):
+The full TLA+ model and TLC configuration files are available in the project repository at https://github.com/chelof100/acp-framework-en [13] (tla/ directory).
+It checks three invariants over a bounded state space (2 agents $\times$ 4 capabilities $\times$ 3 resources, ledger depth $\leq 5$).
+Liveness properties in the extended model depend on two explicit assumptions: monotonic time progression (time does not freeze indefinitely) and weak fairness in request evaluation (an enabled request is eventually processed).
+These assumptions are declared in the Spec formula via WF_vars(Tick) and WF_vars(EvaluateRequest) respectively; they are standard for temporal verification and are stated here to make the verification scope transparent.
+Model scope.
+The TLA+ specification abstracts $F_{\text{ctx}}$ (context factors) and $F_{\text{hist}}$ (history-based signals) from the risk computation.
+Verified properties hold over this abstracted model; their applicability to the full ACP-RISK-2.0 formula is supported empirically via the ACR-1.0 conformance runner and the experiments in §13.7, but is not formally model-checked.
+Table 46 summarizes what is verified, what is empirically tested, and what falls outside the current formal scope.
+Table 46: Verification coverage of ACP-RISK-2.0 components. ✓ = covered; $\times$ = not covered in that layer.
+Component / Signal TLA+ verified Empirically tested Notes
+Cooldown enforcement Temporal property (CooldownExpires)
+Flood detection (Rule 1) Extended model (ACP_Extended.tla)
+Denial escalation (Rule 2) Verified via execution traces
+Pattern anomaly (Rule 3) Context-isolated (patternKey)
+Context factors ($F_{\text{ctx}}$) $\times$ $\times$ Abstracted; future formal work
+History factors ($F_{\text{hist}}$) $\times$ Validated via ACR-1.0 traces
+State-mixing behavior $\times$ Measured in Experiment 7
 ---- MODULE ACP ----
@@
 Interpretation.
-Safety encodes fail-closed admission: every APPROVED decision had $\mathit{RS}\leq 39$.
+Safety encodes fail-closed admission [40]: every APPROVED decision had $\mathit{RS}\leq 39$.
 LedgerAppendOnly encodes tamper-evidence in a state-based view: every ledger entry carries a valid decision and a non-negative risk score.
@@
 RiskDeterminism is the central invariant of ACP-RISK-2.0: identical (capability, resource) pairs always produce identical risk scores, enabling third-party recalculation from a signed policy snapshot.
-Scope of the formal model.
-The TLA+ model intentionally focuses on the deterministic core of the protocol—admission evaluation, ledger append-only property, and risk score determinism—and does not cover all operational aspects of ACP.
-Behavioral properties such as cooldown state transitions, multi-step delegation sequences, and anomaly counter evolution are verified empirically through the ACR-1.0 sequence compliance runner and its five stateful test vectors.
-Extension of the formal model to cover these behaviors is left for future work.
+In the base model, ComputeRisk is a pure function of its inputs, so RiskDeterminism holds by construction.
+Its non-trivial content arises in the extended model, where AnomalyDeterminism extends the guarantee to include per-agent state: identical (capability, resource, pattern_at_eval) triples always produce the same risk score, establishing that RS is a deterministic policy function of the execution trace rather than a statistical estimate.
+Extended model (v1.20): cooldown, delegation integrity, and Fanom flood detection.
+The tla/ACP_Extended.tla module extends ACP.tla with per-agent cooldown temporal state, denial accumulation, a static delegation chain, and—added in Sprint J2—per-agent request pattern accumulation ($F_{\text{anom}}$ flood detection).
+It introduces six additional invariants and three temporal properties beyond ACP.tla:
+  • CooldownEnforced (safety): active cooldown forces DENIED for any request by the affected agent, regardless of risk score.
+  • CooldownImpliesThreshold (safety): cooldown only exists after the denial threshold (COOLDOWN_TRIGGER) has been reached—cooldown is never spurious.
+  • DelegationIntegrity (safety): no consecutive self-delegation in the chain (chain[i] $\neq$ chain[i+1], ACP-DCMA-1.1 §3).
+  • AnomalyDeterminism (safety, Sprint J2): same (capability, resource, pattern_at_eval) triple always produces the same risk score.
+Extends RiskDeterminism to include per-agent state: $\text{RS}=f(\text{static inputs}+\text{request history count})$.
+RS is a deterministic policy function, not a statistical model.
+  • FloodEnforced (safety, Sprint J2): when the per-agent request count meets or exceeds FLOOD_THRESHOLD at evaluation time, the decision is DENIED—the $F_{\text{anom}}$ flood override is never bypassed.
+  • CooldownExpires (liveness): active cooldown eventually expires, conditioned on the expiry time being within the model’s time horizon (cooldown_until $\leq$ MAX_TIME).
+Requires weak fairness on time advance (WF_vars(Tick)) in Spec.
+  • EventuallyRejected (liveness, Sprint J2): an abusive agent not already in cooldown will eventually receive a DENIED decision via the flood override.
+Requires per-agent weak fairness on EvaluateRequest and LEDGER_BOUND >> FLOOD_THRESHOLD.
+  • NoInfiniteExecutionUnderAbuse (liveness, Sprint J2, strong form): no execution trace exists in which an agent exhibiting sustained abusive behavior (request count $\geq$ FLOOD_THRESHOLD) continues to be admitted indefinitely.
+Formally: $\Box(\mathit{Abusive}(a)\Rightarrow\neg\Diamond(\Box\mathit{Accepted}(a)))$.
+Holds in the bounded model via EventuallyRejected; holds in the unbounded specification by structural induction on FloodActive monotonicity (pattern_count is non-decreasing; once $\geq$ FLOOD_THRESHOLD, all subsequent decisions are DENIED).
+Extended TLC result (Sprint J2 configuration).
+Running java -jar tla2tools.jar -deadlock -config ACP_Extended.cfg ACP_Extended.tla with $|\mathit{Agents}|=1$, $|\mathit{Capabilities}|=3$, $|\mathit{Resources}|=2$, LEDGER_BOUND=7=7, FLOOD_THRESHOLD=4=4, FANOM_BONUS=25=25, COOLDOWN_TRIGGER=3=3, COOLDOWN_WINDOW=3=3, MAX_TIME=7=7:
+Model checking completed. No error has been found.
+5,684,342 states generated, 3,147,864 distinct states found, 0 states left on queue.
+The depth of the complete state graph search is 15.
+Finished in 47min 30s (TLC2 v2.16, Java 1.8, single worker, 7241MB heap).
+The model captures agent behavior under repeated request patterns, incorporating stateful anomaly detection and cooldown enforcement.
+All eleven invariants and all four temporal properties hold with zero violations across the complete bounded state space, under the stated assumptions of monotonic time progression, weak fairness in request evaluation, and consistent atomic state updates.
+The graph depth of 15 confirms that the model exercises multiple execution paths and state transitions prior to enforcement—abusive behavior can persist across multiple steps before triggering rejection, confirming that enforcement is not immediate and that liveness properties are not satisfied trivially.
+D4 liveness result.
+In particular, the model verifies a liveness property ensuring bounded execution under sustained abuse.
+No execution trace was found in which an agent exceeding the flood threshold could continue to be admitted indefinitely.
+This result provides evidence that, under the stated assumptions and within the bounded model, the protocol enforces eventual containment of abusive behavior rather than allowing unbounded execution under repeated high-risk patterns.
+Single-agent configuration is used for $F_{\text{anom}}$ liveness verification; per-agent enforcement is independent by design (pattern_count[a] and last_decision[a] are per-agent maps), so the result generalizes structurally to $N$ agents.
+Two-agent safety check (J2c).
+The model was further evaluated under a two-agent configuration (Agents={‘‘A1’’,‘‘A2’’}, FLOOD_THRESHOLD=3, LEDGER_BOUND=5, MAX_TIME=5) to assess potential interference effects between agents.
+Due to the exponential growth of the state space under multi-agent interaction, reduced bounds were used for this configuration.
+All nine safety invariants hold across all 1,650,024 distinct states (3,011,690 states generated, graph depth 11, 1 h 08 min) with zero violations, confirming per-agent enforcement isolation: pattern_count[A1], cooldown state, and admission decisions for A1 do not affect A2’s evaluation state, and vice versa.
+Temporal properties were not enforced in this configuration, as liveness guarantees depend on continued request generation and sufficient time progression.
+Under reduced bounds, executions may terminate without additional requests, leading to spurious liveness violations that do not reflect protocol behavior under sustained interaction.
+The observed behavior highlights an important modeling consideration: temporal guarantees in ACP assume ongoing interaction with the system.
+In bounded executions where request generation ceases, the model may reach terminal states in which no further enforcement transitions occur; these traces do not represent sustained adversarial behavior and are therefore outside the intended scope of the liveness guarantees.
+Accordingly, liveness properties are evaluated under the single-agent configuration (FLOOD_THRESHOLD=4, MAX_TIME=7), where bounds are calibrated to allow sufficient interaction depth.
+Two-agent safety at scale (J2d).
+To strengthen the isolation claim, a further two-agent run was conducted with bounds calibrated to match the single-agent liveness configuration: Agents={‘‘A1’’,‘‘A2’’}, FLOOD_THRESHOLD=4, LEDGER_BOUND=6, MAX_TIME=7, with 6 parallel workers and 12 GB heap (TLC2 v2.16, Java 26).
+The run completed in 4 h 13 min, generating 50,927,772 states and finding 27,648,560 distinct states at graph depth 14.
+All eleven safety invariants hold with zero violations across the complete bounded multi-agent state space, confirming that per-agent enforcement isolation extends to the full invariant set under the liveness-calibrated bounds.
+TLC reported a temporal property violation at the final liveness check, but was unable to generate a counter-example trace: the heap was exhausted (OutOfMemoryError) while constructing the lasso over 165 million accumulated state-space entries.
+Diagnosis reveals a model-bounding artifact rather than a protocol defect.
+With two agents sharing a ledger of capacity LEDGER_BOUND=6=6, the combined request volume fills the ledger before either agent reaches FLOOD_THRESHOLD=4=4, disabling EvaluateRequest for both agents and preventing EventuallyRejected from firing.
+The necessary condition for liveness verification in an $N$-agent configuration is:
+$\texttt{LEDGER\_BOUND}\;\geq\;\texttt{FLOOD\_THRESHOLD}\times N+\delta,$
+[…]
+Two-agent safety at extended scale (LB=11).
+To directly address the scaling constraint identified above, a further run was conducted at the required bound: Agents={‘‘A1’’,‘‘A2’’}, LEDGER_BOUND=11, FLOOD_THRESHOLD=4, MAX_TIME=7, with 15 parallel workers (auto-scaled by TLC) and 24 GB heap on an Intel Core i9-13900HX with 32 GB RAM (TLC2 v2.16, Java 21, Windows 11).
+The run explored 4,294,930,695 distinct states over 10.5 hours, generating 4,648,235,723 total states, with all eleven safety invariants reporting zero violations throughout the entire exploration.
+The run was halted by disk space exhaustion (584 GB of checkpoint data written to disk), not by any property violation.
+This result confirms that all per-agent enforcement invariants hold robustly under the correct multi-agent ledger bound, providing safety verification at $756\times$ the scale of the single-agent configuration and $155\times$ the scale of the LEDGER_BOUND=6 run.
+Liveness verification at this scale remains open: the complete state space at LEDGER_BOUND=11 substantially exceeds current hardware limits for lasso-trace computation.
+Exploration collapse: a formal analog of deviation collapse.
+The bounding artifact above exhibits a structural parallel to the runtime phenomenon studied in Section 17.
+In deviation collapse, an upstream sanitizer eliminates the risk signals that would cause enforcement to fire: the engine remains active and syntactically correct, but its boundary is never exercised.
+In the multi-agent TLC configuration, bounded shared resources eliminate the state conditions under which liveness properties can be exercised [37, 1]: the model checker remains sound and the invariants hold, but the traces required to validate temporal guarantees cannot be reached within the bound.
+Unlike classical state-space reductions—such as partial-order methods, which preserve liveness properties by construction [5]— resource-induced bounds may eliminate entire classes of trajectories from the reachable state space.
+We refer to this as resource-induced liveness obstruction: a structural limitation that arises not from a defect in the protocol or the model, but from the interaction between bounded shared resources and multi-agent concurrency.
+Both cases instantiate the same structural pattern: when the conditions necessary for enforcement to fire disappear—whether from the runtime pipeline or from the formal model—validation loses meaning independently of whether the mechanism itself is correct.
+Put concretely: systems appear compliant because no invalid actions occur; models appear correct because no violations are reachable.
+In both cases, the decision boundary remains defined and the enforcement logic remains intact, but the capacity to exercise failure conditions has been eliminated.
+This observation motivates a modeling requirement analogous to failure condition preservation (Section 17): formal liveness verification of multi-agent systems requires that resource bounds be sufficient to generate the behavioral conditions under which liveness properties are exercised [50, 29].
+Addressing this in the general case—including per-agent ledger allocation, explicit fairness constraints, and scalable model checking strategies—is identified as future work.
+ACP thus provides two distinct guarantees: safety in all executions (invariants hold regardless of scheduling or agent count, verified across 4,294,930,695 distinct states in the two-agent configuration at the correct ledger bound LEDGER_BOUND=11) and liveness under sustained interaction (eventual enforcement holds when agents continue to generate requests within the time horizon, verified for the single-agent configuration where bounds are calibrated to admit the relevant traces).
+Design notes on the extended model.
+The base risk score ComputeRisk(cap, res) is preserved as a pure function; the extended function ComputeRiskWithAnom(cap, res, pat_count) adds FANOM_BONUS when pat_count $\geq$ FLOOD_THRESHOLD, modeling ACP-RISK-2.0 $F_{\text{anom}}$ Rule 3 (rapid-fire anomaly accumulation).
+FloodActive(a) additionally overrides the decision to DENIED regardless of the resulting risk score, ensuring the flood enforcement cannot be bypassed by capability/resource combinations with inherently low base scores.
+Denial counts increment only for RS-based DENIED decisions—not for cooldown-forced or flood-forced denials—consistent with ACP-RISK-2.0 §4 (AddDenial fires on real denials only).
+The delegation chain is a static 2-hop sequence; dynamic chain mutation is outside the scope of the bounded model (reserved for a future version).
+The CooldownExpires and EventuallyRejected liveness properties are conditioned on the relevant state being reachable within the time horizon; the unconditional forms hold in the unbounded specification.
+Framing.
+Sprint J2 extends formal coverage from cooldown enforcement to flood-based $F_{\text{anom}}$ enforcement: the model now verifies not only that individual denials are correct (Safety, CooldownEnforced) but also that sustained abusive behavior is blocked within the bounded model under the stated assumptions (NoInfiniteExecutionUnderAbuse).
+The AnomalyDeterminism invariant formalizes the key correctness property of ACP’s stateful risk engine: the admission decision is fully reproducible given the same capability, resource, and per-agent request history, enabling independent audit without proprietary infrastructure.
+The phrase “formally verified” is deliberately avoided; the correct claim is: model checking of selected safety and liveness properties under a bounded state model, under the stated assumptions of monotonic time progression, weak fairness in request processing, and consistent atomic state updates.
+Model simplifications.
+The TLA+ model uses ComputeRiskWithAnom(cap, res, pat_count) in place of the full ACP-RISK-2.0 formula, which additionally includes $F_{\text{ctx}}$ (context factors) and $F_{\text{hist}}$ (history factors).
+$F_{\text{anom}}$ is modeled as a frequency-based accumulation plus a direct override (FloodActive); semantic payload analysis is out of scope.
+This abstraction is intentional: it captures a fundamental class of abuse patterns (sustained high-frequency interaction from the same agent) while enabling tractable model checking.
+The key contribution is not the detection signal itself, but the verifiable enforcement guarantee—the formal proof that, given a deterministic accumulation function with defined thresholds, the protocol enforces correct admission decisions under stateful conditions, produces auditable execution traces, and allows independent reproduction by third parties.
+This distinguishes ACP from conventional rate limiters, which lack formal specification, execution contracts, and external verifiability.
+pattern_count is monotone within a bounded trace (no per-window decay), approximating the time-windowed behavior of ACP-RISK-2.0 $F_{\text{anom}}$ Rule 3.
+The model verifies properties for representative parameter values (FLOOD_THRESHOLD=4=4, FANOM_BONUS=25=25); complete generalization to arbitrary parameter values is not claimed.
+Full determinism under $F_{\text{hist}}$ is validated empirically via the ACR-1.0 compliance runner and the adversarial evaluation in Section 13.7.
+These simplifications are deliberate: the model is intended to verify structural safety and liveness properties, not to replicate the full engine.
+Limitations of the formal evaluation. The model checking is performed over a bounded model with a finite time horizon and a limited number of agents. As such:
+  • The results do not constitute a proof of correctness for unbounded executions.
+  • Liveness guarantees rely on assumptions of monotonic time progression and weak fairness in request evaluation.
+  • A supplementary two-agent safety check (J2c) confirms per-agent enforcement isolation across 1,650,024 distinct states with 0 violations; temporal liveness properties are verified only in the single-agent configuration with calibrated bounds.
+Full multi-agent contention under adversarial coordination is not modeled.
+  • The state space is constrained by bounded parameters (time horizon, ledger size), which may not capture all possible long-horizon behaviors.
+These constraints define the scope of the guarantees provided by the model and should be considered when interpreting the results.
 Acknowledgements